### v1

# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

Abstract
To support extremely high data rates, reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is considered to be a promising technology for future sixth-generation networks. However, due to the typical employment of hybrid beamforming architecture in THz systems, as well as the passive nature of RIS which lacks the capability to process pilot signals, obtaining channel state information (CSI) is facing significant challenges. To accurately estimate the cascaded channel, we propose a novel low-complexity channel estimation scheme, which includes three steps. Specifically, we first estimate full CSI within a small subset of subcarriers (SCs). Then, we acquire angular information at base station and RIS based on the full CSI. Finally, we derive spatial directions and recover full-CSI for the remaining SCs. Theoretical analysis and simulation results demonstrate that the proposed scheme can achieve superior performance in terms of normalized mean-square-error and exhibit a lower computational complexity compared with the existing algorithms.
Index Terms:
Terahertz, beam split, reconfigurable intelligent surface, wideband channel estimation, direction estimation
I Introduction
Envisioned as a cornerstone technology to meet the requirements of future 6G network, terahertz (THz) communication holds immense potential for a myriad of applications within the field of Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1]. However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [2]. One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [3]. To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide LoS paths by intelligently manipulating the propagation environment [4]. RIS can dynamically adjust reflection properties of a large number of passive reflecting elements to optimize signal transmission and reception [5]. By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems. However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), of which acquisition can be challenging. Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
I-A Prior Work
Over the past few years, numerous research works have been dedicated to explore channel estimation in RIS-assisted high frequency band systems [6, 7, 8, 9, 10, 11]. Specifically, a sparse representation of RIS-assisted millimeter (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [6]. In [7], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel. To reduce computational complexity, [8] proposed a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation. The authors in [9] proposed a two-phase channel estimation strategy to reduce pilot overhead. [10] investigated cooperative-beam-training based channel estimation procedure. In [11], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
The above solutions have effectively addressed the channel estimation problem of mmWave and narrowband THz systems with utilization of the angular domain sparsity. However, THz communication typically adopts a hybrid beamforming architecture to reduce hardware costs[12]. This architecture introduces the beam split effect, where path components split into distinct spatial directions at different SC frequencies due to the wider bandwidth of THz signals and the enormous number of RIS elements [13]. This phenomenon leads to a considerable degradation in achievable data rates, counteracting the benefits of RIS-assisted THz systems.
To address this issue, beam split effect for RIS-assisted THz systems have been investigated in [14, 15, 16]. In particular, [14] introduced a beam-split-aware-orthogonal-matching-pursuit (BSA-OMP) approach to automatically mitigate beam split. In [15], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator. The authors in [16] proposed a robust regularized sensing-beam-split-OMP based scheme to estimate beam split affected RIS-assisted cascaded channel. The above works have achieved accurate channel estimation under beam split effect with the identical core idea of mitigating beam split effect by a well-designed dictionary covering the entire beamspace. However, these algorithms will need to be executed at each SC to ensure estimation accuracy, which may introduce additional unaffordable computations. To further reduce the complexity, literatures [17, 18, 19] have investigated methods by prioritizing the solution of physical directions. Specifically, the authors in [17] regarded channel estimation problem as a block-sparse recovery problem. [18] proposed a beam split pattern detection based wideband channel estimation scheme where the physcical DoAs are prioritized for determination. In [19], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals(AoAs) and angle-of-departure(AoDs) are determined and refined to reconstructed the wideband channel with beam split effect. However, in RIS-assisted THz systems, due to the additional angular information and the coupling of different physical directions introduced by RIS, the aforementioned solutions might be challenging to maintain their low-complexity advantages. Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
I-B Main Contributions
To address the issue discussed above, we presented some preliminary results to estimate the RIS-assisted cascaded channel [20]. In this paper, we significantly expand the previous work and propose a low-complexity channel estimation scheme for RIS-assisted wideband THz systems with beam split. Our main contributions are summarized as follows:
  • We focus on the channel estimation in a general RIS-assisted wideband THz communication scenario with consideration of the beam split effect. In this scenario, we propose a novel low-complexity channel estimation scheme with three steps. To recover the full CSI within a small subset of SCs, we propose a cyclic-beam-split generalized-approximate-message-passing (CBS-GAMP) approach in Step $1$. To reduce computational complexity, we propose a novel beam split angle decoupling (BSAD) scheme where we propose energy-maximum (EnM) and double-sensing-multiple-signal-classification (DS-MUSIC) algorithm to acquire angular information in Step $2$ and we utilize least-squares (LS) solution to obtain the rest angle-excluded information in Step $3$.
  • For the CBS-GAMP approach, we design simplified CBS dictionaries to provide comprehensive coverage of the entire beam space and construct the sparse representation of cascaded channel. We employ this algorithm with expectation-maximum (EM) extensions to acquire full CSI within a small subset of SCs, without the need of prior knowledge of channel distribution and signal-to-noise-ratio (SNR).
  • For the BSAD scheme, we extend the full CSI acquired by CBS-GAMP approach within a small subset of SCs to all the SCs. Specifically, we propose a novel EnM algorithm to estimate the directions-of-departure (DoDs) at BS. Additionally, we design an extended-log multiple-signal-classification pseudo-spectrum (EL-MUSIC-PS) and propose a novel DS-MUSIC algorithm to estimate the coupled directions at RIS. Subsequently, we incorporate the spatial directions derived from the estimated angular information to the unknown CSI and reconstruct the RIS-assisted cascaded channel by a simple LS solution.
The remainder of this paper is organized as follows. In Section II, we explain the RIS-assisted wideband THz system model with beam split. In Section III, we propose a novel CBS-GAMP based full CSI estimation approach. In Section IV, we propose a further simplified BSAD scheme. In Section V, we discuss the numerical simulations and then draw conclusions in Section VI.
Notations: Boldface lower-case and capital letters represent column vectors and matrices, respectively. $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively. $\mathbb{E}$ and $\mathrm{Var}$ represent statistical expectation and variance, respectively. $\mathbb{N}$, $\mathbb{C}$ and $\mathbb{R}$ denote the set of integers, complex numbers and real numbers, respectively. $\mathbf{I}_{N}$ is $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace. $\otimes$ denote Kronecker product, *** and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
II RIS-assisted wideband THz system model
In this section, we first present system model and downlink channel estimation protocol for the considered RIS-assisted wideband THz system with beam split effect. Then, we formulate the RIS-assisted THz channel model and further elaborate beam split effect.
II-A RIS-Assisted Wideband THz System Model
Figure 1: RIS-assisted THz communication system with beam split.
Figure 2: Downlink channel estimation protocol and frame structure.
We consider a RIS-assisted wideband THz MIMO system with hybrid analog/digital beamforming architecture over $M$ subcarriers (SCs), as depicted in Fig. 1. Assuming that the BS with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}(N_{\mathrm{RF}}<N_{\mathrm{T}})$ ratio-chains are deployed to serve $K$ single-antenna UEs. The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS. Besides, the OFDM system is considered with the central frequency $f_{c}$ and bandwidth $B$. The overall downlink channel estimation protocol is shown in Fig. 2. We assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [9]. For each coherence block, channel estimation is firstly performed during $T$ subframes, and the downlink data transmission is then carried out during the remaining $D-T$ subframes. Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots.
Assuming that the direct BS-UE links are obstructed by obstacles [17], the downlink channel $\mathbf{H}_{k}^{\mathrm{RIS}}\left[m\right]\in\mathbb{C}^{1\times N_{\mathrm{T}}}$ from BS to UE-$k$ at the $m$-th SC can be expressed as
$\begin{split}\mathbf{H}_{k}^{\mathrm{RIS}}\left[m\right]&=\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}\right)\mathbf{G}\left[m\right]\\ &=\boldsymbol{\theta}\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right],\end{split}$ (1)
where $\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right]\triangleq\mathbf{G}\left[m\right]\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ is the RIS-assisted cascaded channel and $\boldsymbol{\theta}=\left[e^{j\theta_{1}},\cdots,e^{j\theta_{N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS.
As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[12]. As mentioned above, we consider that $P$ frequency-dependent baseband beamformers $\mathbf{W}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}\triangleq\mathrm{Diag}\left\{s_{1},\cdots,s_{P}\right\}\in\mathbb{C}^{P\times P}$ per time slot. Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain augment in each user direction, wherein each entry is subject to the unit-modulus constraint, i.e., $|\left[\mathbf{W}^{\mathrm{RF}}\right]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ as $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$. Define $\mathbf{W}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}[m]=\left[\mathbf{w}_{1}[m],...,\mathbf{w}_{P}[m]\right]$. At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by UE-$k$ at the $m$-th SC is denoted as
$\begin{split}y_{k,t,p}[m]&=\mathbf{\theta}_{t}\mathbf{H}_{k}^{\mathrm{cas}}[m]\mathbf{w}_{t,p}[m]s_{t,p}[m]+n_{k,t,p}[m],\end{split}$ (2)
where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN). Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t,p}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}\times P}}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
$\tilde{\mathbf{y}}_{k,t}[m]=\boldsymbol{\theta}_{t}\mathbf{H}_{k}^{\mathrm{cas}}[m]\mathbf{R}_{t}[m]+\tilde{\mathbf{n}}_{k,t}[m].$ (3)
We adopt the widely used Saleh-Valenzuela channel model [21] for BS-RIS channel $\mathbf{G}[m]$ and RIS-UE channel $\mathbf{h}_{k}[m]$. The BS-RIS channel $\mathbf{G}[m]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ at the $m$-th SC is expressed as
$\mathbf{G}[m]=\sum_{l=1}^{L}{\alpha_{l}\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{l})\mathbf{a}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\varphi}_{l}\right)e^{-j2\pi\tau_{l}f_{m}}},$ (4)
where $L$ represents the number of propagation paths between BS and RIS, $\alpha_{l}$, $\tau_{l}$, $\tilde{\psi}_{l}$ and $\tilde{\varphi}_{l}$ represent the complex gain, the time delay, the physical angle-of-departure (AoD) at BS and angle-of-arrival (AoA) at RIS for the $l$-th path, respectively. Besides, $f_{m}$ denotes the frequency of the $m$-th SC, which is given by
$f_{m}=f_{c}+\frac{B}{M}\left(m-1-\frac{M-1}{2}\right),\forall m=1,2,\dots,M,$ (5)
where $c$ is the speed of light. Besides, $\mathbf{a}_{\mathrm{T},m}\in\mathbb{C}^{N_{\mathrm{T}}}$ and $\mathbf{a}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}}$ are the normalized array response vectors (ARVs) associated with BS and RIS, respectively, given by
$\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi})=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{\varphi}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
$\mathbf{a}_{\mathrm{R},m}(\tilde{\psi})=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{\psi}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th suncarrier. $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively. Besides, $\mathbf{n}_{t}=\left[0,1,\dots,N_{\mathrm{T}}-1\right]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively. By defining the angle vectors $\tilde{\boldsymbol{\psi}}=\left[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}\right]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as
$\mathbf{G}[m]=\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\boldsymbol{\varphi}}\right),$ (7)
where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=\left[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})\right]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively. $\mathbf{\Upsilon}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix. Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and UE-$k$ at the $m$-th SC can be formulated as
$\mathbf{h}_{k}[m]=\sum_{j=1}^{J_{k}}{\beta_{k,j}\mathbf{a}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\vartheta}_{k,j})e^{-j2\pi\mu_{k,j}f_{m}}},$ (8)
where $J_{k}$ represents the number of propagation paths between RIS and UE-$k$, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the AoD at RIS for the $j$-th path, respectively. By defining $\tilde{\boldsymbol{\vartheta}}_{k}=\left[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}\right]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form, given by
$\mathbf{h}_{k}[m]=\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}),$ (9)
where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-j2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-j2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE. $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the AoDs at RIS. Plugging (7) and (9) into (1), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be written as
$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Upsilon}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
II-B Beam Split Effect
In RIS-assisted wideband THz systems, beam split effect is essentially caused by the analog beamformer and RIS, which contain the frequency-independent phase shifters [22]. The antenna spacing $d_{\mathrm{BS}}$ and $d_{\mathrm{RIS}}$ are set according to the central frequency $f_{c}$, with half-wavelength spacing, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$ as a typical configuration, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [23]. Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
$\mathbf{a}_{\mathrm{T},m}(\varphi)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}\varphi\mathbf{n}_{t}},$ (11a)
$\mathbf{a}_{\mathrm{R},m}(\psi)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}\psi\mathbf{n}_{r}},$ (11b)
where $\eta_{m}=f_{m}/f_{c}$ represents the relative frequency. Without loss of generality, we define $\varphi=\sin\tilde{\varphi}\in[-1,1]$ and $\psi=\sin\tilde{\psi}\in[-1,1]$ as the physical direction for expression simplicity. In THz band, the physical directions $\varphi$ and $\psi$ split into totally separated spatial directions $\eta_{m}\varphi$ and $\eta_{m}\psi$, which is the so-called beam split effect. Due to the large bandwidth and enormous number of RIS elements, beam split effect occurs at both the BS and RIS, making channel estimation algorithms suffer from severe performance degradation [24].
III Full CSI Estimation by CBS-GAMP
In this section, we propose a novel CBS-GAMP based channel estimation approach where simplified CBS dictionaries are firstly presented to formulate angular domain channel. By employing CBS dictionaries, CBS-GAMP algorithm with EM extensions is implemented alternatively for each angular sub-channel to achieve accurate channel estimation.
III-A Angular Domain Sparse Representation
In this subsection, we transformed the THz channel to the angular domain sparse representation by utilizing the overcomplete CBS dictionaries. Specifically, the CBS dictionaries of the BS and RIS directions are formulated as
$\mathbf{C}_{\mathrm{T},m}=\left[\mathbf{a}_{\mathrm{T},m}\left(\bar{\vartheta}_{1}\right),\cdots,\mathbf{a}_{\mathrm{T},m}\left(\bar{\vartheta}_{Q_{\mathrm{T}}}\right)\right],$ (12a)
$\mathbf{C}_{\mathrm{R},m}=\left[\mathbf{a}_{\mathrm{R},m}\left(\bar{\vartheta}_{1}\right),\cdots,\mathbf{a}_{\mathrm{R},m}\left(\bar{\vartheta}_{Q_{\mathrm{R}}}\right)\right],$ (12b)
where $Q_{\mathrm{T}}\gg N_{\mathrm{T}}$ and $Q_{\mathrm{R}}\gg N_{\mathrm{R}}$ are the quantization level of the BS and RIS directions, respectively. Besides, $\bar{\vartheta}_{q}=\delta q-\frac{Q+1}{Q}$ with $Q\in\{Q_{\mathrm{T}},Q_{\mathrm{R}}\}$ denotes the spatial direction sample, $\delta=2/Q$ is the interval between two adjacent samples. Substituting the ARMs in (10) by the CBS dictionaries and extending the angle-excluded matrix/vector, the sparse angular domain representation of the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ is given by
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Upsilon}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Upsilon}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS and the angle-excluded vector of RIS, respectively. According to the operation of transposed Khatri-Rao product and the uniform sampling interval of the CBS dictionary, the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains a considerable amount of duplicated columns. For narrowband channel model, the coupled dictionary contains only $Q_{\mathrm{R}}$ distinct columns, which are exactly the first $Q_{\mathrm{R}}$ columns [6]. However, for wideband channel model with beam split effect, due to the frequency-dependent nature of ARVs, this simplification is no longer suitable. Note that sequence $\bar{\vartheta}_{q_{1}}=\delta q_{1}-\frac{Q+1}{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ corresponds one-to-one with $Q_{\mathrm{R}}$ ARVs in dictionary $\mathbf{C}_{\mathrm{R},m}$, the sequence $-\bar{\vartheta}_{q_{2}},\forall q_{2}=1,\dots,Q_{\mathrm{R}}$ can also establish such correspondence with $\mathbf{C}_{\mathrm{R},m}^{*}$. Therefore, it can be observed that the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ have the form of $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}}),\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. The $j$-th column of the coupled dictionary $\tilde{{\xi}}_{\mathrm{R},m,j}$ can be expressed as
$\tilde{\xi}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (14)
where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of distinct columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}\triangleq 2Q_{\mathrm{R}}-1$. By employing the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, one can easily construct the coupled CBS dictionary as
$\mathbf{\Xi}_{\mathrm{R},m}\triangleq[\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(Q_{\mathrm{R}}^{2}-Q_{\mathrm{R}}+1:Q_{\mathrm{R}}^{2}\right);\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(2:Q_{\mathrm{R}}\right)],$ (15)
where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary. Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (14), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
$\xi_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (16)
By substituting the simplified $\mathbf{\Xi}_{\mathrm{R},m}$ into (13), the RIS-assisted cascaded channel can be rewritten as
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\mathbf{\Xi}_{\mathrm{R},m}\mathbf{X}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (17)
where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Upsilon}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$. $\mathcal{P}_{i}$ is the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ which are identical to the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$. Therefore, there are maximum $\Sigma_{k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$. Employing the property of the Kronecker product, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is given by
$\mathbf{h}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\mathbf{x}_{k,m},$ (18)
where $\mathbf{x}_{k,m}=\mathbf{X}_{k,m}(:)\in\mathbb{C}^{Q_{\mathrm{B}}Q_{\mathrm{T}}}$ represents $\mathbf{X}_{k,m}$ in vector form, $\mathbf{\Pi}_{m}\triangleq\left(\mathbf{C}_{\mathrm{T},m}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},m}\right)\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ denotes the total dictionary. By substituting (18) into (3) and vectorizing, the sparse system formulation can be expressed as
$\mathbf{y}_{k,t}\left[m\right]=\left(\mathbf{R}_{t}^{\mathsf{T}}\left[m\right]\otimes\boldsymbol{\theta}_{t}\right)\mathbf{\Pi}_{m}\mathbf{x}_{k,m}+\mathbf{n}_{k,t}[m],$ (19)
where $\mathbf{y}_{k,t}\left[m\right]\in\mathbb{C}^{P}$ is the vector form of $\tilde{\mathbf{y}}_{k,t}[m]$. Stacking the measurement vectors collected at $T$ successive subframes $\mathbf{y}_{k}\left[m\right]=\left[\mathbf{y}_{k,1}^{\mathsf{T}}\left[m\right],\cdots,\mathbf{y}_{k,T}^{\mathsf{T}}\left[m\right]\right]^{\mathsf{T}}\in\mathbb{C}^{PT}$, the received measurement signal can be expressed as
$\mathbf{y}_{k}[m]=\mathbf{\Psi}_{m}\mathbf{x}_{k,m}+\mathbf{n}_{k}\left[m\right],$ (20)
where $\mathbf{\Psi}_{m}=\overline{\mathbf{R}}_{m}\mathbf{\Pi}_{m}$ and
$\overline{\mathbf{R}}_{m}\triangleq\left[\begin{array}[]{c}\mathbf{R}_{1}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{1}\\ \vdots\\ \mathbf{R}_{T}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{T}\\ \end{array}\right].$ (21)
It is obvious that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix. So far, by employing the CBS dictionaries, the cascaded channel estimation problem has been transformed into the sparse signal recovery problem. Our primary task is to reconstruct the total angle-excluded vectors $\mathbf{x}_{k,m}$ with knowledge of the measurement vectors $\mathbf{y}_{k}[m]$ and the observation matrices $\mathbf{\Psi}_{m}$.
III-B Accurate Sparse Recovery and Channel Reconstruction
The complex sparse recovery problem has been formulated in Subsection III-A. To map this problem to the real number domain for resolution, we then transform the complex form in (20) to the equivalent real form
[$\left[\begin{array}[]{c}\!\!\!\Re\{\mathbf{y}_{k}\left[m\right]\}\\ \!\!\!\Im\{\mathbf{y}_{k}\left[m\right]\}\\ \end{array}\right]\!\!\!=\!\!\!\left[\begin{matrix}\Re\{\mathbf{\Psi}_{m}\}&\!\!\!\!-\Im\{\mathbf{\Psi}_{m}\}\\ \Im\{\mathbf{\Psi}_{m}\}&\!\!\!\!\Re\{\mathbf{\Psi}_{m}\}\\ \end{matrix}\right]\!\!\left[\begin{array}[]{c}\!\!\!\!\Re\{\mathbf{x}_{k,m}\}\\ \!\!\!\!\Im\{\mathbf{x}_{k,m}\}\\ \end{array}\right]\!+\!\left[\begin{array}[]{c}\!\!\!\Re\{\mathbf{n}_{k}\left[m\right]\}\\ \!\!\!\Im\{\mathbf{n}_{k}\left[m\right]\}\\ \end{array}\right]\!\!,$ (22)
By systematically assigning symbols to each matrix in (22), the sparse recovery problem is formulated as
$\boldsymbol{y}=\mathbf{\Psi}\boldsymbol{x}+\boldsymbol{n}.$ (23)
For simplicity, we employ a simplified subscript notation by omitting the index for the matrix and vectors. By defining $T_{\mathrm{tot}}\triangleq PT$ and $Q_{\mathrm{tot}}\triangleq Q_{\mathrm{B}}Q_{\mathrm{T}}$, $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$ is the measurement vector, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ denotes the observation matrix and $\boldsymbol{x}\boldsymbol{\in}\mathbb{C}^{Q_{\mathrm{tot}}}$ with sparsity $\Sigma_{\mathrm{tot}}$ denotes the sparse representation vector to be estimated. Note that unlike conventional THz channel models where $\Sigma_{\mathrm{tot}}=\Sigma_{k},\forall k=1,\dots,K$, the sparsity of RIS-assisted THz channel satisfies $\Sigma_{\mathrm{tot}}\leq\Sigma_{k}$ and remains uncertain. This is because the introduction of simplified dictionary and RIS phase-shift matrix, resulting in coupling between AoAs and AoDs, which can be challenging to be separated. In this case, we adopt the Bayesian framework, i.e. GAMP approach, to recover the angle-excluded coefficients in $\boldsymbol{x}$, which do not require the sparsity level $\Sigma_{\mathrm{tot}}$ as the prior input. Bayesian approaches require only the prior distribution assumption of $\boldsymbol{x}$, where Gaussian distribution model is widely used in rich-scattering environments [25]. However, this assumption becomes inapplicable for THz band due to the existence of angular domain sparsity makes $\boldsymbol{x}$ a sparse vector with $Q_{\mathrm{tot}}\gg\Sigma_{\mathrm{tot}}$. Since the sparse vector $\mathbf{x}\in\mathbb{C}^{Q^{\rm tot}}$ is supported on a subspace with dimension smaller than $Q^{\rm tot}$, a mixture stochastic distribution can be employed to capture its sparsity [26, 27]:
$p_{X}(x)=\lambda f_{X}(x)+(1-\lambda)\delta(x),$ (24)
where $\delta(\cdot)$ is the Dirac delta, $f_{X}(x)$ is the active-coefficient probability density function (PDF), and $\lambda$ is the sparsity rate that parameterizes the signal sparsity. By the weak law of large numbers [28], $\left\|\boldsymbol{x}\right\|_{0}/Q_{\mathrm{tot}}\xrightarrow{\mathbb{P}}\lambda$, where the $\ell_{0}$-norm $\left\|\cdot\right\|_{0}$ denotes the number of nonzero entries of a vector.
Then, we adopt the Gaussian distribution as an active-coefficient PDF in (24) which is widely employed as a priori distribution [29, 25], i.e., $f_{X}(x)=\mathcal{N}(x,u,v)$, the mixture stochastic distribution in (24) can be rewritten as
$p_{X}(x)=\lambda\mathcal{N}(x,u,v)+(1-\lambda)\delta(x),$ (25)
where $u$ and $\nu$ are the mean and the variance, respectively, which correspond to the large-scale fading coefficient that accounts for the combined effects of shadowing and path loss. Inspired by (25), we treat the sparse vector as Bernoulli-Gaussian (BG) signal with unknown priori $\boldsymbol{q}\triangleq\left[\lambda,u,v\right]$ and variance $\varepsilon$ of measurement noise.
In the Bayesian estimation framework, we employ the widely adopted minimum mean squared error (MMSE) estimation for each $x_{j}$ separately by approximating the marginal posterior distributions $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y}\right)$ [30]
$\hat{x}_{j}^{\mathrm{MMSE}}=\mathbb{E}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y}\right\}.$ (26)
We then employ the GAMP based framework to find the true marginal posterior distributions $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y}\right)$, which is summarized in Algorithm 1.
Algorithm 1 GAMP for MMSE Estimation
0:  Observation matrix $\mathbf{\Psi}\in\mathbb{R}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$, measurement vector $\boldsymbol{y}\in\mathbb{R}^{T_{\mathrm{tot}}}$, noise variance $\varepsilon$, prior distribution $p_{X}(x;\boldsymbol{q})$, precision tolerance $\tau_{\mathrm{max}}$ and maximum number of iterations $N_{\mathrm{iter}}$.
1:  $n=1$.
2:  $\forall j:\hat{x}_{j}(n)=\int_{x}{xp_{X}(x;\boldsymbol{q})}$.
3:  $\forall j:\mu_{j}^{x}(n)=\int_{x}{\left|x-\hat{x}_{j}(n)\right|^{2}p_{X}(x;\boldsymbol{q})}$.
4:  $\forall i:\widehat{s_{i}}(n-1)=0$.
5:  while $n\leq N_{\mathrm{iter}}$ do
6:     $\forall i:\mu_{i}^{p}(n)=\sum_{j=1}^{Q_{\mathrm{tot}}}{\left|\mathbf{\Psi}_{i,j}\right|^{2}}\mu_{j}^{x}(n)$.
7:     $\forall i:\hat{p}_{i}(n)=\sum_{j=1}^{Q_{\mathrm{tot}}}{\mathbf{\Psi}_{i,j}}\hat{x}_{j}(n)-\mu_{i}^{p}(n)\hat{s}_{i}(n-1)$.
8:     $\forall i:\mu_{i}^{z}(n)=\mathrm{Var}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i}(n),\mu_{i}^{p}(n),\boldsymbol{q}\right\}$.
9:     $\forall i:\hat{z}_{i}(n)=\mathbb{E}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i}(n),\mu_{i}^{p}(n),\boldsymbol{q}\right\}$.
10:     $\forall i:\mu_{i}^{s}(n)=\frac{1}{\mu_{i}^{p}(n)}\left[1-\frac{\mu_{i}^{z}(n)}{\mu_{i}^{p}(n)}\right]$.
11:     $\forall i:\hat{s}_{i}(n)=\frac{1}{\mu_{i}^{p}(n)}\left[\hat{z}_{i}(n)-\hat{p}_{i}(n)\right]$.
12:     $\forall j:\mu_{j}^{r}(n)=\left(\sum_{i=1}^{T_{\mathrm{tot}}}{\left|\mathbf{\Psi}_{i,j}\right|^{2}}\mu_{i}^{s}(n)\right)^{-1}$.
13:     $\forall j:\hat{r}_{j}(n)=\hat{x}_{j}(n)+\mu_{j}^{r}(n)\sum_{i=1}^{T_{\mathrm{tot}}}{\mathbf{\Psi}_{i,j}^{*}}\hat{s}_{i}(n)$.
14:     $\forall j:\hat{x}_{j}(n+1)=\mathbb{E}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j}(n),\mu_{j}^{r}(n),\boldsymbol{q}\right\}$.
15:     $\forall j:\mu_{j}^{x}(n+1)=\mathrm{Var}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j}(n),\mu_{j}^{r}(n),\boldsymbol{q}\right\}$.
16:     $n\leftarrow n+1$.
17:     if $\frac{\sum_{j=1}^{Q_{\mathrm{tot}}}{\left|\hat{x}_{j}(n+1)-\hat{x}_{j}(n)\right|^{2}}}{\sum_{j=1}^{Q_{\mathrm{tot}}}{\left|\hat{x}_{j}(n)\right|^{2}}}<\tau_{\mathrm{max}}$ then
18:        break.
19:     end if
20:  end while
20:  MMSE estimation $\left\{\hat{x}_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$.
The relationship between the $i$-th measured output $y_{i}$ and the corresponding noiseless output $z_{i}\triangleq\boldsymbol{\psi}_{i}^{\mathsf{T}}\boldsymbol{x}$ in GAMP algorithm can be modeled by the conditional PDF $p_{Y|Z}\left(y_{i}|z_{i};\boldsymbol{q}\right)$, where $\boldsymbol{\psi}_{i}^{\mathsf{T}}$ is the $i$-th row of $\mathbf{\Psi}$. The message from the factor node $p\left(z_{i}|\boldsymbol{\psi}_{i}^{\mathsf{T}}\boldsymbol{x}\right)$ to the variable node $z_{i}$ is calculated by integrating all the variable nodes $\left\{x_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ in $\boldsymbol{x}$. As $Q_{\mathrm{tot}}$ is quite large and $x_{j}$ is assumed independent and identically distributed (i.i.d.), according to the central limit theorem (CLT), the conditional distribution $z_{i}$ given $x_{j}$ can be approximated by a Gaussian distribution $\mathcal{N}\left(z_{i};\hat{p}_{i},\mu_{i}^{p}\right)$ with mean $\hat{p}_{i}$ and variance $\mu_{i}^{p}$. Therefore, the true marginal posterior $p_{Z|\boldsymbol{Y}}\left(z_{i}|\boldsymbol{y};\boldsymbol{q}\right)$ can be expressed as
$p_{Z|\boldsymbol{Y}}\left(z_{i}|\boldsymbol{y};\hat{p}_{i},\mu_{i}^{p},\boldsymbol{q}\right)\triangleq\frac{p_{Y|Z}\left(y_{i}|z_{i};\boldsymbol{q}\right)\mathcal{N}\left(z_{i};\hat{p}_{i},\mu_{i}^{p}\right)}{\int_{z}{p_{Y|Z}}\left(y_{i}|z;\boldsymbol{q}\right)\mathcal{N}\left(z;\hat{p}_{i},\mu_{i}^{p}\right)},$ (27)
where the quantities $\hat{p}_{i}$ and $\mu_{i}^{p}$ are updated with iteration $n$ in lines $7$ and $6$ of Algorithm 1, respectively. Since the noise $\boldsymbol{n}$ is treated as AWGN, the conditional probability $p_{Y|Z}\left(y_{i}|z_{i};\boldsymbol{q}\right)$ follows Gaussian distribution $\mathcal{N}\left(y_{i};z_{i},\varepsilon\right)$. Therefore, the PDF in (27) has moments
$\hat{z}_{i}=\mathbb{E}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i},\mu_{i}^{p},\boldsymbol{q}\right\}=\frac{\mu_{i}^{p}y_{i}+\varepsilon\hat{p}_{i}}{\mu_{i}^{p}+\varepsilon},$ (28a)
$\mu_{i}^{z}=\mathrm{Var}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i},\mu_{i}^{p},\boldsymbol{q}\right\}=\frac{\mu_{i}^{p}\varepsilon}{\mu_{i}^{p}+\varepsilon},$ (28b)
where the quantities $\hat{z}_{i}$ and $\mu_{i}^{z}$ are updated with iteration $n$ in lines $9$ and $8$ of Algorithm 1, respectively. Employing CLT again, the conditional distribution $x_{j}$ given $z_{i}$ can be approximated by a Gaussian distribution $\mathcal{N}\left(x_{j};\hat{r}_{j},\mu_{j}^{r}\right)$ with mean $\hat{r}_{j}$ and variance $\mu_{j}^{r}$. Therefore, the true marginal posterior $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{q}\right)$ can be expressed as
$p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right)\triangleq\frac{p_{X}\left(x_{j};\boldsymbol{q}\right)\mathcal{N}\left(x_{j};\hat{r}_{j},\mu_{j}^{r}\right)}{\int_{x}{p_{X}}(x;\boldsymbol{q})\mathcal{N}\left(x;\hat{r}_{j},\mu_{j}^{r}\right)},$ (29)
where the quantities $\hat{r}_{j}$ and $\mu_{j}^{r}$ are updated with iteration $n$ in lines $13$ and $12$ of Algorithm 1, respectively. By substituting the BG distribution $p_{X}\left(x_{j};\boldsymbol{q}\right)$ into (29) and after simple algebraic operations, we define $\zeta_{j}$ as the integration at the denominator in (29), which is expressed as
$\begin{split}&\zeta_{j}\triangleq\int_{x}{p_{X}}(x;\boldsymbol{q})\mathcal{N}\left(x;\hat{r}_{j},\mu_{j}^{r}\right)\\ &=(1-\lambda)\mathcal{N}\left(0;\hat{r}_{j},\mu_{j}^{r}\right)+\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right),\end{split}$ (30)
Similarly, the numerator in (29) can be derived as
$\begin{split}&p_{X}\left(x_{j};\boldsymbol{q}\right)\mathcal{N}\left(x_{j};\hat{r}_{j},\mu_{i}^{r}\right)=(1-\lambda)\mathcal{N}\left(0;\hat{r}_{j},\mu_{j}^{r}\right)\delta\left(x_{j}\right)\\ &+\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right)\mathcal{N}\left(x_{j};\gamma_{j},\nu_{j}\right),\end{split}$ (31)
where $\gamma_{j}\triangleq(\hat{r}_{j}\nu+\mu\mu_{j}^{r})/(\mu_{j}^{r}+\nu)$ and $\nu_{j}\triangleq\mu_{j}^{r}\nu/(\mu_{j}^{r}+\nu)$ are the $\left(\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right)$-dependent quantities and the simplification in (30) and (31) holds by the Gaussian-PDF multiplication rule where
$\mathcal{N}\!(x;a,\!A)\mathcal{N}\!(x;b,\!B)\!=\!\mathcal{N}\!\!\left(\!x;\!\frac{aB\!\!+\!\!bA}{A\!+\!B}\!,\!\frac{AB}{A\!\!+\!\!B}\!\right)\!\mathcal{N}\!(0;\!a-b,\!A+B).$ (32)
By defining
$\pi_{j}\triangleq\frac{\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right)}{\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right)+(1-\lambda)\mathcal{N}\left(0;\hat{r}_{j},\mu_{j}^{r}\right)},$ (33)
we can further simplify (29) as
$p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right)=\left(1-\pi_{j}\right)\delta\left(x_{j}\right)+\pi_{j}\mathcal{N}\left(x_{j};\gamma_{j},\nu_{j}\right).$ (34)
So far, the posterior $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y}\right)$ is rewritten into the form of BG distribution with quantities $\left(\pi_{j},\gamma_{j},\nu_{j}\right)$. Wherein $\pi_{j}$ is the sparsity rate corresponing to the posterior support probability $\operatorname{Pr}\left\{x_{j}\neq 0|\boldsymbol{y};\boldsymbol{q}\right\}$, $\gamma_{j}$ and $\nu_{j}$ are the corresponding mean and variance, respectively. Given the PDF $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{q}\right)$ in (34), the moments can be computed in closed form
$\hat{x}_{j}=\mathbb{E}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j},\mu,\boldsymbol{q}\right\}=\pi_{j}\gamma_{j},$ (35a)
$\mu_{j}^{x}=\mathrm{Var}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right\}=\pi_{j}^{2}\nu_{j},$ (35b)
where the quantities $\hat{x}_{j}$ and $\mu_{j}^{x}$ are updated with iteration $n$ in lines $14$ and $15$ of Algorithm 1, respectively. Essentially, the quantities $\left(\pi_{j},\gamma_{j},\nu_{j}\right)$ all depend on the prior parameters $\boldsymbol{q}$ in (25), which are unknown beforehand. Therefore, in next subsection, we employ the EM based approach to learn $\boldsymbol{q}$ from the received measurements $\boldsymbol{y}$.
III-C EM-Based Prior Parameters Learning
In this subsection, we aim to determine the maximum-likelihood estimate $\boldsymbol{\alpha}\triangleq\left[\boldsymbol{q},\varepsilon\right]$ of the prior signal parameters and noise variance. The prior parameters $\boldsymbol{\alpha}$ can be learned and updated by EM approach. Specifically, with knowledge of the posterior probabilities acquired by GAMP algorithm in Section III-B, EM algorithm iteratively increases the log-likelihood function (LLF) to guarantee the convergence to a local maximum of the likelihood $p_{\boldsymbol{Y}}(\boldsymbol{y};\boldsymbol{\alpha})$ [31]. The LLF can be approximated by GAMP posteriors, which yields EM update as
$\boldsymbol{\alpha}^{n+1}=\mathop{\mathrm{arg}\max}\limits_{\boldsymbol{\alpha}}\mathbb{E}_{\boldsymbol{X|Y}}\left\{\ln p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\alpha})\right\},$ (36)
where $p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y})$ is the joint PDF parameterized by $\boldsymbol{\alpha}$, which is given by
$p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\alpha})=p_{\boldsymbol{Y|X}}\left(\boldsymbol{y}|\boldsymbol{x};\varepsilon\right)p_{\boldsymbol{X}}(\boldsymbol{x};\boldsymbol{q}).$ (37)
Substituting (37) into (36), the expectation can be divided into two parts
$\begin{split}&\mathbb{E}_{\boldsymbol{X|Y}}\left\{\ln p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\alpha})\right\}\\ &=\!\!\int_{\boldsymbol{x}}\!\!{\ln p_{\boldsymbol{Y\!|X}}\!\!\left(\boldsymbol{y}|\boldsymbol{x};\!\varepsilon\right)\!p_{\boldsymbol{X\!|Y}}\!\!\left(\boldsymbol{x}|\boldsymbol{y};\!\boldsymbol{\alpha}\right)}\!+\!\!\int_{\boldsymbol{x}}\!\!{\ln\!p_{\boldsymbol{X}}\!(\boldsymbol{x};\!\boldsymbol{q})p_{\boldsymbol{X\!|Y}}\!\!\left(\boldsymbol{x}|\boldsymbol{y};\!\boldsymbol{\alpha}\right)}.\end{split}$ (38)
As observed in (38), the expectation of LLF in (36) is decoupled in terms of the signal parameters $\boldsymbol{q}$ and noise variance $\varepsilon$, which can be updated separately. Following the well-established incremental updating schedule proposed in [32], we update only one parameter in $\boldsymbol{\alpha}$ at one time while keeping other parameters unchanged.
III-C1 EM update for $\lambda$
Since $\lambda$ is independent of noise variance $\varepsilon$, the update of $\lambda$ is only relevant to the latter part in (38), which is given by
$\lambda^{n+1}\!=\mathop{\mathrm{arg}\max}\limits_{\lambda\in\left(0,1\right)}\!\sum_{j=1}^{Q_{\mathrm{tot}}}\!{\int_{x_{j}}\!{\ln p_{X}(x_{j};\lambda,\mu^{n},\nu^{n})p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}},$ (39)
Clearly, $\lambda^{n+1}$ is the value of $\lambda$ where the expectation in (39) is at the point of zero derivative
$\sum_{j=1}^{Q_{\mathrm{tot}}}{\int_{x_{j}}{p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\lambda}\ln p_{X}(x_{j};\lambda,\mu^{n},\nu^{n})}}=0.$ (40)
Substituting (25) into (40), one can derive that
$\frac{d}{d\lambda}\ln p_{X}\left(x_{j};\lambda,\mu^{n},\nu^{n}\right)=\begin{cases}\frac{1}{\lambda-1}&x_{j}=0\\ \frac{1}{\lambda}&x_{j}\neq 0\end{cases},$ (41)
Given the calculation in (34) and (41), it is previous that the neighborhood around $x_{j}=0$ which we define as $\mathcal{U}_{e}=U(0,e)$ is nonconsecutive from the remainder $\bar{\mathcal{U}}_{e}=\mathbb{R}\backslash\mathcal{U}_{e}$ in (40). Therefore, we calculate the integration in (40) within the neighborhood $\mathcal{U}_{e}$ and the remainder $\bar{\mathcal{U}}_{e}$ separately
$\lambda\!=\!\underset{e\rightarrow 0}{\lim}\sum_{j=1}^{Q_{\mathrm{tot}}}\!{\frac{\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}{\int_{x_{j}\in\bar{\mathcal{U}}_{e}}\!{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\!+\!\int_{x_{j}\in\mathcal{U}_{e}}\!{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (42)
Substituting the above observation and (34) into (42), the update of $\lambda$ can be easily obtained
$\lambda^{n+1}=\underset{e\rightarrow 0}{\lim}\sum_{j=1}^{Q^{\mathrm{tot}}}{\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{q}^{n}\right)}=\frac{1}{Q^{\mathrm{tot}}}\sum_{j=1}^{Q^{\mathrm{tot}}}{\pi_{j}},$ (43)
where $\left\{\pi_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ are computed from GAMP outputs via (33).
III-C2 EM update for $\mu$
Similar to (39), the update for $\mu$ is given by
$\mu^{n+1}\!=\!\underset{\mu\in\mathbb{R}}{\mathrm{arg}\max}\sum_{j=1}^{Q_{\mathrm{tot}}}\!\!{\int_{x_{j}}{\ln p_{X}\left(x_{j};\lambda^{n},\mu,\nu^{n}\right)p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (44)
Similarly, $\mu^{n+1}$ corresponding to the maximum expectation in (44) is again at the point of zero derivative
$\sum_{j=1}^{Q_{\mathrm{tot}}}{\int_{x_{j}}{p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\mu}\ln p_{X}\left(x_{j};\lambda^{n},\mu,\nu^{n}\right)}=0}.$ (45)
Substituting (25) into (45), one can derive that
$\frac{d}{d\mu}\ln p_{X}\left(x_{j};\lambda^{n},\mu,\nu^{n}\right)=\begin{cases}0&x_{j}=0\\ \frac{x_{j}-\mu}{\nu^{n}}&x_{j}\neq 0\\ \end{cases},$ (46)
Excluding the neighborhood $\mathcal{U}_{e}$, (45) can be rewritten as
$\sum_{j=1}^{Q_{\mathrm{tot}}}{\lim_{e\rightarrow 0}\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{\left(x_{j}-\mu\right)}p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}=0.$ (47)
Then, the unique $\mu$ can be derived from (47) as
$\mu^{n+1}=\frac{1}{\lambda^{n+1}Q_{\mathrm{tot}}}\sum_{j=1}^{Q_{\mathrm{tot}}}{\pi_{j}}\gamma_{j},$ (48)
where $\left\{\pi_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ and $\left\{\gamma_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ are computed from GAMP outputs via (34).
III-C3 EM update for $\nu$
Similar to (39), the update for $\nu$ is given by
$\nu^{n+1}\!=\!\underset{\nu>0}{\mathrm{arg}\max}\sum_{j=1}^{Q_{\mathrm{tot}}}\!\!\!{\int_{x_{j}}{\ln p_{X}\left(x_{j};\lambda^{n},\mu^{n},\nu\right)p_{X|\boldsymbol{Y}}\!\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (49)
Similarly, $\nu^{n+1}$ corresponding to the maximum expectation in (49) is again at the point of zero derivative
$\sum_{j=1}^{Q_{\mathrm{tot}}}{\int_{x_{j}}{p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\nu}\ln p_{X}\left(x_{j};\lambda^{n},\mu^{n},\nu\right)}=0}.$ (50)
Substituting (25) into (50), one can derive that
$\nu^{n+1}=\begin{cases}0&x_{j}=0\\ \frac{1}{2}\left(\frac{\left|x_{j}-\mu^{n}\right|^{2}}{\left(\nu\right)^{2}}-\frac{1}{\nu}\right)&x_{j}\neq 0\\ \end{cases}.$ (51)
Excluding the neighborhood $\mathcal{U}_{e}$, (50) can be rewritten as
$\sum_{j=1}^{Q_{\mathrm{tot}}}{\lim_{e\rightarrow 0}\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{\left(\left|x_{j}-\mu^{n}\right|^{2}-\nu\right)}p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}=0.$ (52)
Then, the unique $\nu$ can be derived from (52) as
$\nu^{n+1}=\frac{1}{\lambda^{n+1}Q_{\mathrm{tot}}}\sum_{j=1}^{Q_{\mathrm{tot}}}{\pi_{j}\left(\left|\mu^{n}-\gamma_{j}\right|^{2}+\nu_{j}\right)},$ (53)
where $\left\{\pi_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$, $\left\{\gamma_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ and $\nu_{j}$ are computed from GAMP outputs via (33) and (34).
III-C4 EM update for $\varepsilon$
Given $\boldsymbol{q}^{n}$ as determined above, we finally update the noise variance $\varepsilon$. Under the AWGN assumption, the noise is i.i.d and independent from $\boldsymbol{x}$. Therefore, the update of $\varepsilon$ is only relevant to the former part in (38), which is given by
$\varepsilon^{n+1}=\underset{\varepsilon>0}{\mathrm{arg}\max}\sum_{i=1}^{T_{\mathrm{tot}}}{\int_{y_{i}}{\ln p_{Y|\boldsymbol{X}}\left(y_{i}|\boldsymbol{x};\varepsilon\right)p_{\boldsymbol{X}|\boldsymbol{Y}}\left(\boldsymbol{x}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (54)
Recalling that $z_{i}\triangleq\boldsymbol{\psi}_{i}^{\mathsf{T}}\boldsymbol{x}$ and we define $w_{i}\triangleq y_{i}-z_{i}$ as the AWGN variable, the posterior $p_{Y|\boldsymbol{X}}\left(y_{i}|\boldsymbol{x};\varepsilon\right)$ can be re-expressed as
$p_{Y|\boldsymbol{X}}\left(y_{i}|\boldsymbol{x};\varepsilon\right)=p_{W}\left(w_{i};\varepsilon\right)=\mathcal{N}\left(w_{i};0,\varepsilon\right).$ (55)
Similarly, $\varepsilon^{n+1}$ corresponding to the maximum expectation in (54) is again at the point of zero derivative
$\sum_{i=1}^{T_{\mathrm{tot}}}{\int_{w_{i}}{p\left(w_{i}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\varepsilon}\ln p_{W}\left(w_{i};\varepsilon\right)}}=0.$ (56)
Substituting (55) into (56), one can derive that
$\frac{d}{d\varepsilon}\ln p_{W}\left(w_{i};\varepsilon\right)=\frac{1}{2}\left(\frac{\left|w_{i}\right|^{2}}{(\varepsilon)^{2}}-\frac{1}{\varepsilon}\right).$ (57)
Then, substituting (57) into (56), the unique $\varepsilon$ can be expressed as
$\varepsilon^{n+1}=\frac{1}{T_{\mathrm{tot}}}\sum_{i=1}^{T_{\mathrm{tot}}}{\left(\left|y_{i}-\hat{z}_{i}\right|^{2}+\mu_{i}^{z}\right)},$ (58)
where $\left\{\hat{z}_{i}\right\}_{i=1}^{T_{\mathrm{tot}}}$ and $\left\{\mu_{i}^{z}\right\}_{i=1}^{T_{\mathrm{tot}}}$ are computed from GAMP outputs via (III-B).
III-C5 EM Initialization
Since EM approach iteratively converges to a the local maximum of the LLF in (36), we initialize the parameters $\boldsymbol{\alpha}$ as set in [30]
$\lambda^{0}=\frac{\Sigma_{k}}{Q_{\mathrm{tot}}},\mu^{0}=0,\mathsf{SNR}^{0}=10\mathrm{dB},$ (59a)
$\varepsilon^{0}=\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}}{\left(\mathsf{SNR}^{0}+1\right)T_{\mathrm{tot}}},\nu^{0}=\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}-T_{\mathrm{tot}}\varepsilon^{0}}{\mathrm{tr}\left(\mathbf{\Psi}^{\mathsf{T}}\mathbf{\Psi}\right)\lambda^{0}}.$ (59b)
III-D CBS-GAMP Based Full CSI Estimation
Employing CBS dictionaries, the beam-split affected channel is easily formulated as sparse recovery problem at each SC $m\in\mathcal{M}$. In this case, EM and GAMP algorithms are iteratively implemented to alternately update MMSE estimates $\left\{x_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ and prior parameters $\boldsymbol{\alpha}$ at all SCs. By substituting MMSE estimates $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$ into (18), the full CSI is available
$\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m}.$ (60)
IV Low-Complexity Channel Estimation by BSAD
The CBS-GAMP algorithm obtained from Section III-B ensures the accuracy of channel estimation. However, this algorithm may introduce additional computational complexity due to its execution at each SC. It is obvious that the frequency-dependent spatial directions affected by beam split effect share a common physical support. Inspired by this observation, we propose a novel BSAD scheme to reduce computational complexity. In our proposed algorithm, we only employ the CBS-GAMP to acquire the full CSI within a small subset of SCs, namely the first and the $M/2$-th SCs [^1] The selection of SC indices is demonstrated in Section IV-B. Then, employing the obtained full CSI of the subset of SCs, we design EnM algorithm and DS-MUSIC algorithm to determine the common support at BS and RIS, respectively. Finally, employing the obtained common support and the frequency-dependent nature introduced by beam split effect, we transform the channel estimation problem at the remaining SCs into a simple linear regression problem, which can be easily solved using LS method.
IV-A EnM Based DoDs Estimation at BS
In this subsection, we derive the DoDs at BS from the full CSI $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[\bar{m}],\bar{m}\in\left\{1,M/2\right\}$ estimated by the CBS-GAMP algorithm. Recalling (60), the estimated full CSI is consisted of two parts, i.e., the total CBS dictionary $\mathbf{\Pi}_{\bar{m}}$ and the angle-excluded vector $\mathbf{x}_{k,\bar{m}}$, which are known separately. Based on the fact that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\mathbf{x}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$, the common support can be easily determined. However, the inevitable noise and estimation errors will cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, leading to more than $\Sigma_{k}$ non-zero entries. Therefore, we propose a novel EnM method to determine the common support $\zeta^{\mathrm{BS}}$ and then calculate the corresponding physical DoDs at BS, as summarized in Algorithm 2.
Algorithm 2 EnM Based DoDs Estimation at BS
0:  Angle-excluded matrix $\mathbf{x}_{k,\bar{m}},\bar{m}\in\left\{1,M/2\right\}$, the path number $L$ of BS-RIS channel and the quantization level of CBS dictionaries $Q_{\mathrm{R}}$ and $Q_{\mathrm{T}}$.
1:  $\boldsymbol{\epsilon}=\left[\epsilon_{1},\cdots,\epsilon_{Q_{\mathrm{T}}}\right]$ and $\bar{\boldsymbol{\varphi}}=-1+\frac{1+2\boldsymbol{q}}{Q_{\mathrm{T}}},\boldsymbol{q}=1,\dots,Q_{\mathrm{T}}$.
2:  $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}=\mathrm{normalize}\left(\left[\left|x_{k,\bar{m},1}\right|,\cdots,\left|x_{k,\bar{m},Q_{\mathrm{tot}}}\right|\right]\right)$.
3:  for $q\in\left\{1,2,\dots,Q_{\mathrm{T}}\right\}$ do
4:     $\epsilon_{q}=\sum_{\bar{m}\in\left\{1,\frac{M}{2}\right\}}{\left\|\tilde{\boldsymbol{\epsilon}}_{\bar{m}}\left(\left(\left(q-1\right)Q_{\mathrm{B}}+1\right):qQ_{\mathrm{B}}\right)\right\|_{2}}$.
5:  end for
6:  $\zeta^{\mathrm{BS}}=\left\{q|\mathop{\mathrm{maxk}}\limits_{q}\left(\boldsymbol{\epsilon},L\right)\right\}$
7:  $\hat{\boldsymbol{\varphi}}=\bar{\boldsymbol{\varphi}}\left(\zeta^{\mathrm{BS}}\right)$.
7:  $\zeta^{\mathrm{BS}}$ and $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$.
In the proposed EnM method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, in which each entry is the normalized results of the corresponding element in $\mathbf{x}_{k,\bar{m}}$ (Step $2$). As mentioned in Section III-A, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, with each block contains $Q_{\mathrm{B}}$ columns. Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way. By accumulating the $\ell_{2}$-norm for each block of the two normalized energy basis vectors, the merged-energy vector $\boldsymbol{\epsilon}$ can establish a one-to-one correspondence with the physical DoDs at BS in the entire beam space (Step $3-5$). Subsequently, the indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ can constitute the common support $\zeta^{\mathrm{BS}}$ (Step $6$). Finally, with knowledge of $\zeta^{\mathrm{BS}}$, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ is easily obtained according to the sampling rule of the CBS dictionary (Step $7$). Fig. 3 depicts an example of the estimated energy spectrum employing the proposed EnM method. Assuming that the cascaded channel contains $L=3$ BS-RIS paths with the common support $\zeta^{\mathrm{BS}}=\left\{3,4,24\right\}$ and SNR $=20$ dB. It can be observed from the spectrum that the energy is distributed surrounding the estimated DoDs. Furthermore, the energy is concentrated around the actual direction and can be easily distinguished, which validates the effectiveness of the proposed EnM method.
Figure 3: Normalized energy spectrum via EnM algorithm.
IV-B DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
The proposed EnM method ensures the acquisition of the DoDs at BS. However, due to the coupling between the DoAs and DoDs at RIS, the aforementioned method can be challenging to obtain angular information at RIS. Therefore, in this subsection we propose a novel DS-MUSIC algorithm to acquire the coupled DoAs/DoDs at RIS.
Firstly, with knowledge of the estimated DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ from Section IV-A, the estimated cascaded channel $\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}]$ can be projected onto the subspace of ARM of BS using the pseudo-inverse, which is given by
$\begin{split}\mathbf{U}_{k,\bar{m}}&\triangleq\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}](\hat{\mathbf{A}}_{N_{\mathrm{T}},\bar{m}}^{\mathsf{H}}\left(\boldsymbol{\varphi}\right))^{\dagger}\\ &=\mathrm{Diag}\left(\mathbf{h}_{k}[\bar{m}]\right)\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\psi}\right)\mathbf{\Upsilon}_{\bar{m}},\end{split}$ (61)
where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$ and $\left(\mathbf{A}_{N}^{\mathsf{H}}\right)^{\dagger}=(\mathbf{A}_{N}^{\dagger})^{\mathsf{H}}$. Besides, the rest-CSI matrix $\mathbf{U}_{k,\bar{m}}\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ can be block-partitioned column-wise as $\mathbf{U}_{k,\bar{m}}=\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]$. By substituting (7) into (61), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
$\begin{split}\mathbf{u}_{k,\bar{m},l}&=\!\upsilon_{\bar{m},l}\left(\mathbf{A}_{\mathrm{R},\bar{m}}^{*}\left(\boldsymbol{\vartheta}_{k}\right)\boldsymbol{\beta}_{k,\bar{m}}\right)\bullet\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\right)\\ &=\!\upsilon_{\bar{m},l}\!\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\!-\!\vartheta_{k,1}\!\right)\!,\!\cdots\!,\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\!-\!\vartheta_{k,J_{k}}\!\right)\right]\!\boldsymbol{\beta}_{k,\bar{m}},\end{split}$ (62)
where $\upsilon_{l}$ is the $\left(l,l\right)$-th entry in $\mathbf{\Upsilon}_{m}$. Define $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM where $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right],l=1,\dots,L,j=1,\dots,J_{k}$ and $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\upsilon_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ as the amplitude vector, (62) is rewritten as
$\mathbf{u}_{k,\bar{m},l}=\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\boldsymbol{\chi}_{k,\bar{m},l}.$ (63)
To find the coupled physical directions $\left\{{\hat{\gamma}}_{k,l,j}\right\}_{j=1}^{J_{k}}$ in (63), this issue can be regarded as a frequency estimation finding problem. Inspired by this, we propose a novel DS-MUSIC algorithm as described in Algorithm 3, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
Algorithm 3 DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
[…]
1:  $\bar{\boldsymbol{\gamma}}=-2\left(1+\boldsymbol{q}/Q_{\mathrm{R}}\right),\boldsymbol{q}=1,\dots,Q_{\mathrm{B}}$.
2:  for $l\in\left\{1,2,\dots,L\right\}$ do
3:     for ${\bar{m}}\in\left\{1,\frac{M}{2}\right\}$ do
4:        $\mathbf{R}_{k,\bar{m},l}^{u}=\frac{1}{N_{\mathrm{a}}}\tilde{\mathbf{U}}_{k,\bar{m},l}\tilde{\mathbf{U}}_{k,\bar{m},l}^{\mathsf{H}}$.
5:        $\left[\tilde{\mathbf{\Omega}}_{k,\bar{m},l},\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l}\right]=\mathrm{eig}\left(\mathbf{R}_{k,\bar{m},l}^{u}\right)$.
6:        $\mathcal{S}_{l}=\left\{n_{\lambda}|\underset{n_{\lambda}}{\mathrm{mink}}\left(\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l},N_{\mathrm{R}}-J_{k}\right)\right\}$.
7:        $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}=\tilde{\mathbf{\Omega}}_{k,\bar{m},l}\left(\mathcal{S}_{l}\right)$.
8:        for $q\in\left\{1,2,\dots,Q_{\mathrm{B}}\right\}$ do
9:           $\tilde{p}_{q}^{\mathrm{MUSIC}}=\frac{1}{\left\|\left(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\right)^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}}$.
10:        end for
11:        $\mathbf{p}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\!=\!\log\!\left[\!\min\!\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\right)\!\!,\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\!\!,\min\!\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\right)\!\right]$.
[…]
13:     end for
14:     $\mathcal{Q}_{l}=\left\{\frac{q_{1}+q_{\frac{M}{2}}^{\prime}}{2}\left|\underset{q_{1},q_{\frac{M}{2}}^{\prime}}{\min\mathrm{k}}\left(d\left(q_{1},q_{\frac{M}{2}}^{\prime}\right),J_{k}\right)\right.\right\}$ where $q_{1}\in\tilde{\mathcal{Q}}_{1,l}$ and $q_{\frac{M}{2}}^{\prime}\in\tilde{\mathcal{Q}}_{\frac{M}{2},l}$.
15:  end for
16:  $\zeta^{\mathrm{RIS}}=\mathcal{Q}_{1}\cup\cdots\cup\mathcal{Q}_{L}$.
17:  $\hat{\boldsymbol{\gamma}}_{k}=\bar{\boldsymbol{\gamma}}\left(\zeta^{\mathrm{RIS}}\right)$.
17:  $\zeta^{\mathrm{RIS}}$ and $\hat{\boldsymbol{\gamma}}_{k}$.
Due to the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC, some energy from physical directions will leak into the noise space, making it challenging to distinguish physical directions. To address this issue, we preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray contains $N_{\mathrm{sub}}$ reflecting elements. There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}},N_{\mathrm{sub}}>J_{k}$. The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
$\mathbf{u}_{k,\bar{m},l,n}=\mathbf{A}_{\mathrm{R},\bar{m},n}\left(\boldsymbol{\gamma}_{k,l}\right)\boldsymbol{\chi}_{k,\bar{m},l},$ (64)
where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(n:n+N_{\mathrm{sub}}-1\right)$. The covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ is given by
$\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}=\mathbf{u}_{k,\bar{m},l,n}\mathbf{u}_{k,\bar{m},l,n}^{\mathsf{H}}.$ (65)
The covariance matrix $\mathbf{R}_{k,\bar{m},l}^{u}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ of $\mathbf{u}_{k,\bar{m},l}$ is calculated by averaging all $N_{\mathrm{a}}$ $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}$, which is given by
$\mathbf{R}_{k,\bar{m},l}^{u}=\frac{1}{N_{\mathrm{a}}}\sum_{n=1}^{N_{\mathrm{a}}}{\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}}=\frac{1}{N_{\mathrm{a}}}\tilde{\mathbf{U}}_{k,\bar{m},l}\tilde{\mathbf{U}}_{k,\bar{m},l}^{\mathsf{H}},$ (66)
where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ in (64). The eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ yields
$\mathbf{R}_{k,\bar{m},l}^{u}=\tilde{\mathbf{\Omega}}_{k,\bar{m},l}\mathrm{Diag}(\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l})\tilde{\mathbf{\Omega}}_{k,\bar{m},l}^{\mathsf{H}},$ (67)
where $\tilde{\boldsymbol{\Omega}}_{k,\bar{m},l}\triangleq\left[\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,1},\cdots,\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,N_{\mathrm{sub}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ is the eigenbasis matrix which contains $N_{\mathrm{sub}}$ eigenvectors and $\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l}\in\mathbb{C}^{N_{\mathrm{sub}}}$ is the eigenvalue vector. It is obvious that $\mathbf{R}_{k,\bar{m},l}^{u}$ is a Hermitian matrix of all its $N_{\mathrm{sub}}$ eigenvectors are orthogonal to each other. Arrange the eigenvalues of $\mathbf{R}_{k,\bar{m},l}^{u}$ in descending order and organize the corresponding eigenvectors accordingly, it can be rewritten as
$\mathbf{R}_{k,\bar{m},l}^{u}=\mathbf{\Omega}_{k,\bar{m},l}\mathrm{Diag}\left(\boldsymbol{\lambda}_{k,\bar{m},l}\right)\mathbf{\Omega}_{k,\bar{m},l}^{\mathsf{H}},$ (68)
where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ are the sorted eigenbasis matrix and eigenvalue vector, respectively. Subsequently, $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts with $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$. Wherein, $J_{k}$ largest eigenvalues corresponding to the directions of largest variability span the signal subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$, while the remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors span the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which is orthogonal to the signal subspace. Define the sub-ARV $\mathbf{a}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}(\gamma)\in\mathbb{C}^{N_{\mathrm{sub}}}$ as the first $N_{\mathrm{sub}}$ row of $\mathbf{a}_{\mathrm{R},\bar{m}}(\gamma)$. Similar to the definition of sub-ARV, we define $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the sub-CBS dictionary which consists of the first $N_{\mathrm{sub}}$ rows. By employing the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}$, one can scan the entire beam space and calculate the MUSIC-pseudo-spectrum (PS) for the beams aligned with any sampled direction. Define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}\triangleq[\tilde{p}_{k,\bar{m},l,1}^{\mathrm{MUSIC}},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}^{\mathrm{MUSIC}}]^{\mathsf{T}}$ as the MUSIC-PS, wherein $\tilde{p}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}$ is the spectral value, which follows a $\ell_{2}$-norm form as
$\tilde{p}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}},$ (69)
where $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\!\triangleq\!e^{\!-j2\pi\eta_{m}\left(\!\frac{q}{Q_{\mathrm{R}}}\!-1\right)\mathbf{n}_{\mathrm{sub}}}$ with $\mathbf{n}_{\mathrm{sub}}\!=\!\left[1,\cdots,N_{\mathrm{sub}}\right]$. The MUSIC-PS is obtained by mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$. Based on the fact that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ that resides in the signal subspace is orthogonal to the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the $\ell_{2}$-norm in (69) associated with the actual directions should approximate zero. The MUSIC algorithm is typically dependent on peak detection to find actual physical directions, resulting in undetectable marginal peaks. To address this issue, we design a novel extended-log-MUSIC-PS (EL-MUSIC-PS), where we extend the spectrum and map it to the logarithmic domain to enhance the accuracy of estimation. The EL-MUSIC-PS can be express as
$\mathbf{p}_{k,\bar{m},l}^{\mathrm{MUSIC}}=\log\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\right),\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\right)\right].$ (70)
In this form, the actual directions can be determined by employing local-maximum algorithms. Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{q}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{q}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (71)
where $K_{q}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$. It is observed that when $q^{\prime}=q+K_{q}Q_{\mathrm{R}}/\eta_{m}$, $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ will generate the identical beam. Moreover, the maximum number of identical beams $K_{q}^{\max}$ can be obtained by solving the integer equation $K_{q}^{\max}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$, which can be derived as
$K_{q}^{\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (72)
Therefore, more than $J_{k}$ peaks would be detected in the EL-MUSIC-PS in (70), making it challenging to distinguish the true and fake physical directions in one single SC. Thankfully, it is observed from (72) that the spatial directions associated with true physical directions keep frequency-independent. Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction. In this approach, we select two SCs, calculate and compare their EL-MUSIC-PSs on the same plot. The positions where peaks overlap indicate the true physical directions. However, due to the unavoidable noise and estimation errors in full CSI, the EL-MUSIC-PSs of different SCs may exist offsets. Therefore, in practical applications, we select the $J_{k}$ closest peaks to represent true physical directions, and we introduce the peak distance $d$ to measure the distance between two peaks, which is given by
$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=\left|q_{m}-q_{m^{\prime}}^{\prime}\right|,$ (73)
where $q_{m}$ and $q_{m^{\prime}}^{\prime}$ denote the indices of the peaks at two selected SCs, respectively. It can be observed from (72) that the selection of SC $m$ affects the number of identical beams. To avoid additional algorithm complexity, the selection of SCs should satisfy the constraint $K_{q,\max}\leq 2$. Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into it, we can obtain
$m\leq\frac{M}{2}+\frac{Mf_{c}Q_{\mathrm{R}}}{B\left(Q_{\mathrm{B}}\right)}+\frac{1}{2}.$ (74)
Substituting (5) into (73), the peak distance $d$ can be derived as
$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{q}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (75)
To distinguish true physical directions effectively, the peak distance should be sufficiently large. According to (75), it is evident that the intervals between selected SCs should be large enough. Therefore, under the constraint in (74), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation. Fig. 4 depicts an example of estimated EL-MUSIC-PS with the proposed DS approach. Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for a arbitrary UE with SNR $=20$ dB. It can be observed that the peaks overlap in actual physical directions, while that associated with fake spatial directions are totally separated. The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
Figure 4: Normalized EL-MUSIC-PS via DS-MUSIC algorithm.
IV-C LS Estimation for Angle-Excluded Coefficients
With the estimated DoDs at BS and coupled DoAs/DoDs at RIS, (10) can be rewritten as
$\mathbf{h}_{k}^{\mathrm{cas}}[m]=\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\otimes\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\tilde{\mathbf{x}}_{k,m},$ (76)
where $\tilde{\mathbf{x}}_{k,m}\!\triangleq\!\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Upsilon}_{m}\!\in\mathbb{C}^{L^{2}\!J_{k}}$ denotes the angle-excluded vector, $\hat{\boldsymbol{\varphi}}\triangleq[\hat{\varphi}_{1},\dots,\hat{\varphi}_{L}]\in\mathbb{C}^{L}$ denotes the estimated DoDs at BS of all $L$ paths, and $\hat{\boldsymbol{\gamma}}_{k}\triangleq[\hat{\boldsymbol{\gamma}}_{k,1}^{\mathsf{T}},\dots,\hat{\boldsymbol{\gamma}}_{k,L}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{LJ_{k}}$ denotes the estimated coupled directions at RIS. Substituting (76) into (19), the received signal $\mathbf{y}_{k,t}\left[m\right]$ can be rewritten as
$\mathbf{y}_{k,t}\left[m\right]=\tilde{\mathbf{\Psi}}_{m}\tilde{\mathbf{x}}_{k,m},$ (77)
where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\overline{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$. In this case, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained employing LS solution
$\hat{\tilde{\mathbf{x}}}_{k,m}=\left(\tilde{\mathbf{\Psi}}_{m}^{\mathsf{H}}\tilde{\mathbf{\Psi}}_{m}\right)^{-\mathsf{1}}\tilde{\mathbf{\Psi}}_{m}^{\mathsf{H}}\mathbf{y}_{k,t}\left[m\right].$ (78)
Then, the full CSI can be calculated by
$\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\otimes\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\hat{\tilde{\mathbf{x}}}_{k,m}.$ (79)
IV-D Complexity Analysis
In this subsection, we analyze the computational complexity of the proposed channel estimation scheme. Specifically, the complexity of the proposed scheme mainly stems from four major steps: 1) Full CSI estimation via CBS-GAMP approach $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$, 2) DoDs acquisition at BS via EnM method $\mathcal{O}\left(Q_{\mathrm{T}}Q_{\mathrm{B}}\right)$, 3) Coupled directions acquisition at RIS via DS-MUSIC algorithm $\mathcal{O}\left(2N_{r}^{3}\right)$, 4) Angle-excluded coefficients estimation via LS solution $\mathcal{O}\left(\left(M-2\right)L^{6}J_{k}^{3}\right)$. The overall complexity of the proposed scheme is given by
$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right)$ (80)
For comparison, we also discuss the complexity of the conventional OMP scheme [33], BSA-OMP scheme [14] and CBS-GAMP scheme. Since the above algorithms need to be executed at each SC, the total computational complexity thus is multiplied by $M$ compared to the original. The conventional OMP algorithm has the same complexity $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$ as BSA-OMP approach. The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$. Due to the sparse characteristics of THz channel, the number of propagation path satisfies $L,J_{k}\ll Q_{\mathrm{T}},Q_{\mathrm{R}}$. The number of SCs $M$ is also quite large in wideband systems. Therefore, it is clear that the computational complexity of the proposed scheme is much lower than other three approaches.
V Numerical Results and Discussion
In this section, we present numerical results to validate the effectiveness of the proposed scheme. Without loss of generality, we consider a wideband RIS-assisted THz scenario with $f_{c}=100$ GHz, $B=15$ GHz and $M=128$. A BS employing a $N_{\mathrm{T}}$ = 16 ULA transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements. For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths and the AoAs and AoDs are uniformly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$. The quantization level of CBS dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$. As a metric for CSI estimation performance, we use the normalized mean square error (NMSE), which is given by
$\mathrm{NMSE}=\mathbb{E}\left\{\frac{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]-\mathbf{\hat{h}}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}\right\}$ (81)
Besides, we employ the correct probability as a metric for direction estimation performance, which is given by
$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\}$ (82)
where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ is the threshold which is set as half the sample interval. For comparison, we also consider the conventional OMP scheme [33], BSA-OMP scheme [14], CBS-GAMP scheme, and oracle-LS scheme as benchmark schemes in the simulation. It is notice that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
Figure 5: Correct probability of DoDs at BS and coupled DoAs/DoDs at RIS against SNR.
Fig. 5 depicts the correct probability performance of the DoDs at BS and coupled DoAs/DoDs at RIS versus SNR with $T=50$ and $P=30$. It is shown that both the two correct probabilities approach one as SNR increases. This indicates that the proposed scheme can ensure a certain level of reliability even with low SNR conditions and has the ability to obtain accurate physical directions with sufficiently well channel conditions. Besides, the correct probability of the coupled directions at RIS is lower than that of the DoDs at BS in all SNR regions. This is because the estimation of the former relies on the latter. Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves.
Figure 6: NMSE performance comparison against SNR.
Fig. 6 illustrates the NMSE performance versus SNR with $T=50$ and $P=30$. Overall, we can observe that the proposed scheme outperforms conventional-OMP and BSA-OMP algorithm in all SNR regions. This is because the proposed scheme has the ability to exploit unknown channel information, and the proposed dictionaries can effectively cover the beam space under beam split effect. Besides, it can be observed that the performance of the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB. This is because estimation errors generated by partial SCs propagate more significantly to the remaining SCs under poor channel conditions. However, the proposed scheme outperforms CBS-GAMP algorithm in high SNR regions, i.e., SNR is larger than $10$ dB. This is because angle decoupling provides more accurate angular estimation information. This implies that there exists a performance-complexity trade-off between these two algorithms. Moreover, it is obvious that the proposed scheme can approach NMSE performance of the oracle-LS scheme, verifying the effectiveness of the proposed scheme.
Figure 7: NMSE performance comparison against the number of subframes $T$.
Fig. 7 shows the NMSE performance versus the number of subframes $T$, with SNR $=20$ dB and $P=15$. As expected, with the increase in the number of subframes $T$, NMSE performance of all the algorithms improves. We can observe that the proposed scheme can maintain well performance even in the case with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm. Moreover, It can be observed that compared to the CBS-GAMP method, the proposed algorithm can reach similar performance with approximately $50\%$ subframes reduction. This indicates that the proposed scheme can efficiently reduce the pilot overhead for channel estimation.
Figure 8: NMSE performance comparison against the number of time slots $P$.
Fig. 8 provides the NMSE performance versus the number of time slots $P$, with SNR $=20$ dB and $T=30$. As expected, with the increase in the number of slots $P$, NMSE performance of all the algorithms improves. We can observe that the proposed scheme outperforms all other three algorithms and can achieve near-optimal NMSE performance of the ideal oracle-LS upper bound in all considered $P$. The performance gap between the proposed scheme and other existing schemes is quite great, demonstrating that the proposed scheme can effectively reduce the system complexity for channel estimation.
Figure 9: NMSE performance comparison against the number of paths $L,J_{k}$.
Fig. 9 illustrates the NMSE performance versus the number of paths $L$ and $J_{k}$, with $T=50$ and $P=30$. We can observe that as the the number of paths increases, the NMSE performance of both the two schemes degrade. This is because of the increased estimation parameters and challenges in distinguishing directions between more multipaths. Furthermore, as SNR increases, the NMSE of CBS-GAMP algorithm converges to around $5\times 10^{-4}$ regardless of the number of paths, while the NMSE of the proposed scheme remain decreasing in all SNR regions. This indicates that the proposed scheme exhibits a superior performance.
Figure 10: NMSE performance comparison against the number of reflecting elements of RIS $N_{\mathrm{R}}$.
Fig. 10 shows the NMSE performance versus the number of reflecting elements of RIS $N_{\mathrm{R}}$, with $T=50$ and $P=30$. We can observe that as $N_{\mathrm{R}}$ decreases, the NMSE performance of the BSA-OMP algorithm degrade greatly. This is because the small-scale RIS cannot provide enough measurement data diversity to support OMP based algorithms. In contrast, the performance of the proposed and CBS-GAMP schemes slightly improve as $N_{\mathrm{R}}$ decreases. This indicates that these two schemes exhibit robust framework, which are less susceptible to the lack of data diversity.
VI Conclusion
In this paper, we investigated channel estimation in RIS-assisted wideband systems with consideration of beam split effect. We proposed an efficient and low-complexity channel estimation scheme. Specifically, we first proposed a novel CBS-GAMP approach to estimate full CSI within a small subset of SCs. Then, to reduce computational complexity, we proposed a novel BSAD scheme which consists of two direction estimation algorithm, i.e., EnM and DS-MUSIC to acquire angular information at BS and RIS, respectively. With knowledge of angular information, we transformed channel estimation into a simple linear-regression problem and recovered CSI at the remaining SCs using a simple LS algorithm. We demonstrated from numerical results that the proposed scheme can achieve superior NMSE performance and significantly reduce pilot overhead.

### v2

@@
-Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split
+Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split
 Abstract
-To support extremely high data rates, reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is considered to be a promising technology for future sixth-generation networks.
-However, due to the typical employment of hybrid beamforming architecture in THz systems, as well as the passive nature of RIS which lacks the capability to process pilot signals, obtaining channel state information (CSI) is facing significant challenges.
-To accurately estimate the cascaded channel, we propose a novel low-complexity channel estimation scheme, which includes three steps.
-Specifically, we first estimate full CSI within a small subset of subcarriers (SCs).
-Then, we acquire angular information at base station and RIS based on the full CSI.
-Finally, we derive spatial directions and recover full-CSI for the remaining SCs.
-Theoretical analysis and simulation results demonstrate that the proposed scheme can achieve superior performance in terms of normalized mean-square-error and exhibit a lower computational complexity compared with the existing algorithms.
+Reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is emerging as a key technology to support ultra-high data rates in future sixth-generation networks.
+However, the acquisition of accurate channel state information (CSI) in such systems is challenging due to the passive nature of RIS and the hybrid beamforming architecture typically employed in THz systems.
+To address these challenges, we propose a novel low-complexity two-phase channel estimation scheme for RIS-assisted THz systems with beam split effect.
+In the proposed scheme, we first estimate the full CSI over a small subset of subcarriers, then extract angular information at both the base station and RIS.
+Subsequently, we recover the full CSI across remaining subcarriers by determining the corresponding spatial directions and angle-excluded coefficients.
+Theoretical analysis and simulation results demonstrate that the proposed method achieves superior performance in terms of normalized mean-square error while significantly reducing computational complexity compared to existing algorithms.
 Index Terms:
@@
 I Introduction
-Envisioned as a cornerstone technology to meet the requirements of future 6G network, terahertz (THz) communication holds immense potential for a myriad of applications within the field of Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1].
-However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [2].
+Envisioned as a cornerstone technology to meet the requirements of future 6G networks, terahertz (THz) communication holds immense potential for a myriad of applications within the field of the Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1, 2, 3].
+However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [4, 5].
 One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [6].
@@
 By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems.
-However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), of which acquisition can be challenging.
+However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), the acquisition of which can be challenging.
 Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
 I-A Prior Work
-Over the past few years, numerous research works have been dedicated to explore channel estimation in RIS-assisted high frequency band systems [6, 7, 8, 9, 10, 11].
-Specifically, a sparse representation of RIS-assisted millimeter (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [6].
+Over the past few years, numerous research works have been dedicated to exploring channel estimation in RIS-assisted high frequency band systems [9, 10, 11, 12, 13, 14].
+Specifically, a sparse representation of RIS-assisted millimeter wave (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [9].
 In [10], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel.
@@
 In [14], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
-The above solutions have effectively addressed the channel estimation problem of mmWave and narrowband THz systems with utilization of the angular domain sparsity.
+The above solutions have effectively addressed the channel estimation problem of narrowband systems with utilization of the angular domain sparsity.
 However, THz communication typically adopts a hybrid beamforming architecture to reduce hardware costs[15].
-This architecture introduces the beam split effect, where path components split into distinct spatial directions at different SC frequencies due to the wider bandwidth of THz signals and the enormous number of RIS elements [13].
-This phenomenon leads to a considerable degradation in achievable data rates, counteracting the benefits of RIS-assisted THz systems.
-To address this issue, beam split effect for RIS-assisted THz systems have been investigated in [14, 15, 16].
-In particular, [14] introduced a beam-split-aware-orthogonal-matching-pursuit (BSA-OMP) approach to automatically mitigate beam split.
+When considering this architecture, the beam split effect becomes inevitable.
+This effect occurs due to the wider bandwidth of THz signals and the enormous number of RIS elements, which causes path components to split into distinct spatial directions at different subcarrier frequencies [16].
+This phenomenon leads to a considerable degradation in achievable data rates, which counteracts the benefits of RIS-assisted THz systems.
+To address this issue, beam split effect in RIS-assisted THz systems has been investigated in [17, 18, 19].
+In particular, [17] introduced a beam-split-aware orthogonal matching pursuit (BSA-OMP) approach to automatically mitigate beam split.
 In [18], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator.
-The authors in [16] proposed a robust regularized sensing-beam-split-OMP based scheme to estimate beam split affected RIS-assisted cascaded channel.
-The above works have achieved accurate channel estimation under beam split effect with the identical core idea of mitigating beam split effect by a well-designed dictionary covering the entire beamspace.
-However, these algorithms will need to be executed at each SC to ensure estimation accuracy, which may introduce additional unaffordable computations.
-To further reduce the complexity, literatures [17, 18, 19] have investigated methods by prioritizing the solution of physical directions.
+The authors in [19] proposed a robust regularized sensing-beam-split-OMP-based scheme to estimate beam split affected RIS-assisted cascaded channel.
+These works primarily focus on mitigating beam split effects within the context of cascaded channels using well-designed dictionaries covering the entire beamspace.
+However, an alternative approach is to focus on the physical directions of the channel components to reduce computational complexity [20, 21, 22].
 Specifically, the authors in [20] regarded channel estimation problem as a block-sparse recovery problem.
-[18] proposed a beam split pattern detection based wideband channel estimation scheme where the physcical DoAs are prioritized for determination.
-In [19], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals(AoAs) and angle-of-departure(AoDs) are determined and refined to reconstructed the wideband channel with beam split effect.
-However, in RIS-assisted THz systems, due to the additional angular information and the coupling of different physical directions introduced by RIS, the aforementioned solutions might be challenging to maintain their low-complexity advantages.
+[21] proposed a beam split pattern detection based wideband channel estimation scheme where the physical DoAs are prioritized for determination.
+In [22], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals (AoAs) and angle-of-departure (AoDs) are determined and refined to reconstruct the wideband channel with beam split effect.
+Although these algorithms achieve accurate channel estimation, they require execution at each SC or simultaneous processing of all SCs, which may introduce additional unaffordable computations.
+Moreover, the passive nature of RIS complicates the precise acquisition of angles at the RIS, which often results in the availability of only coupled angles.
 Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
 I-B Main Contributions
-To address the issue discussed above, we presented some preliminary results to estimate the RIS-assisted cascaded channel [20].
-In this paper, we significantly expand the previous work and propose a low-complexity channel estimation scheme for RIS-assisted wideband THz systems with beam split.
+To address the issue discussed above, we presented some preliminary results on estimating the RIS-assisted cascaded channel [23].
+In this paper, we significantly expand on the previous work and propose a two-phase efficient channel estimation scheme for RIS-assisted wideband THz systems with beam split.
 Our main contributions are summarized as follows:
-  • We focus on the channel estimation in a general RIS-assisted wideband THz communication scenario with consideration of the beam split effect.
-In this scenario, we propose a novel low-complexity channel estimation scheme with three steps.
-To recover the full CSI within a small subset of SCs, we propose a cyclic-beam-split generalized-approximate-message-passing (CBS-GAMP) approach in Step $1$.
-To reduce computational complexity, we propose a novel beam split angle decoupling (BSAD) scheme where we propose energy-maximum (EnM) and double-sensing-multiple-signal-classification (DS-MUSIC) algorithm to acquire angular information in Step $2$ and we utilize least-squares (LS) solution to obtain the rest angle-excluded information in Step $3$.
-  • For the CBS-GAMP approach, we design simplified CBS dictionaries to provide comprehensive coverage of the entire beam space and construct the sparse representation of cascaded channel.
-We employ this algorithm with expectation-maximum (EM) extensions to acquire full CSI within a small subset of SCs, without the need of prior knowledge of channel distribution and signal-to-noise-ratio (SNR).
-  • For the BSAD scheme, we extend the full CSI acquired by CBS-GAMP approach within a small subset of SCs to all the SCs.
-Specifically, we propose a novel EnM algorithm to estimate the directions-of-departure (DoDs) at BS.
-Additionally, we design an extended-log multiple-signal-classification pseudo-spectrum (EL-MUSIC-PS) and propose a novel DS-MUSIC algorithm to estimate the coupled directions at RIS.
-Subsequently, we incorporate the spatial directions derived from the estimated angular information to the unknown CSI and reconstruct the RIS-assisted cascaded channel by a simple LS solution.
+  • We propose a novel low-complexity channel estimation protocol for general RIS-assisted wideband THz communication scenarios with beam split effect, as depicted in Fig.
+1.
+The channel estimation scheme is divided into two phases: full CSI estimation for two SCs in Phase I and remaining SCs estimation in Phase II.
+By employing information from only two SCs, this protocol can significantly reduce the computational complexity while maintaining complete and accurate channel estimation.
+  • In Phase I, we propose a two-step method to acquire full CSI for the selected two SCs.
+Specifically, in Step 1, we propose a simplified cyclic beam split (CBS) dictionary to reduce the size from the scale of $Q^{2}$ to a linear scale of $Q$.
+Using this dictionary, we reformulate the problem as a generalized sparse recovery problem.
+Moreover, we discuss the selection of different sparse recovery algorithms.
+In Step 2, we first design a heuristic DoD estimation method at BS according to the estimated cascaded channel obtained in Step 1.
+We then reformulate the coupled angles at RIS as a typical DoA estimation problem and design a double sensing-multiple signal classification (DS-MUSIC) algorithm by integrating the angular information from the two selected SCs.
+We also provide insights into the selection of two SCs.
+In Phase II, given angular information from Phase I, we employ a simple least squares (LS) algorithm to estimate the channels for the remaining SCs.
+  • Through the analysis of algorithm complexity, our proposed protocol significantly reduces the computational complexity compared to the existing sparse recovery solutions.
+Additionally, numerical results demonstrate that the NMSE performance of our proposed channel estimation scheme closely approaches that of oracle-LS method.
+By focusing on a two-step estimation process for two SCs and applying a simplified LS algorithm for the remaining SCs, the proposed method can achieve both reduction of computational complexity and acquisition of accurate CSI.
+Figure 1: Two-phase channel estimation protocol and frame structure.
 The remainder of this paper is organized as follows.
-In Section II, we explain the RIS-assisted wideband THz system model with beam split.
-In Section III, we propose a novel CBS-GAMP based full CSI estimation approach.
-In Section IV, we propose a further simplified BSAD scheme.
-In Section V, we discuss the numerical simulations and then draw conclusions in Section VI.
+In Section II, we first describe the RIS-assisted wideband THz communication setup and the beam split effect.
+In Section III, we describe our proposed channel estimation protocol.
+Section V provides an analysis of the computational complexity of our proposed scheme, with a comparison to existing methods.
+Finally, we conclude the paper in Section VI.
 Notations: Boldface lower-case and capital letters represent column vectors and matrices, respectively.
@@
 $\mathbf{I}_{N}$ is $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace.
-$\otimes$ denote Kronecker product, *** and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
-II RIS-assisted wideband THz system model
-In this section, we first present system model and downlink channel estimation protocol for the considered RIS-assisted wideband THz system with beam split effect.
-Then, we formulate the RIS-assisted THz channel model and further elaborate beam split effect.
-II-A RIS-Assisted Wideband THz System Model
-Figure 1: RIS-assisted THz communication system with beam split.
-Figure 2: Downlink channel estimation protocol and frame structure.
-We consider a RIS-assisted wideband THz MIMO system with hybrid analog/digital beamforming architecture over $M$ subcarriers (SCs), as depicted in Fig.
-1.
-Assuming that the BS with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}(N_{\mathrm{RF}}<N_{\mathrm{T}})$ ratio-chains are deployed to serve $K$ single-antenna UEs.
+$\otimes$ denote Kronecker product, * and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
+II System and Channel Model
+II-A Signal and Channel Model
+Figure 2: RIS-assisted THz communication system with beam split effect.
+We consider a RIS-assisted wideband THz MIMO system employing hybrid beamforming architecture that integrates both analog and baseband processing, as depicted in Fig.
+2.
+Assuming that the BS equipped with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}$ (where $N_{\mathrm{RF}}<N_{\mathrm{T}}$) RF chains is deployed to serve $K$ single-antenna UEs.
 The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS.
-Besides, the OFDM system is considered with the central frequency $f_{c}$ and bandwidth $B$.
-The overall downlink channel estimation protocol is shown in Fig.
-2.
-We assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [9].
-For each coherence block, channel estimation is firstly performed during $T$ subframes, and the downlink data transmission is then carried out during the remaining $D-T$ subframes.
+Besides, the OFDM system is considered with $M$ subcarriers, central frequency $f_{c}$ and bandwidth $B$.
+In this paper, we assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [12].
+For each coherence block, channel estimation is firstly performed during $T$ subframes, and downlink data transmission is then carried out during the remaining $D-T$ subframes.
 Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots.
-Assuming that the direct BS-UE links are obstructed by obstacles [17], the downlink channel $\mathbf{H}_{k}^{\mathrm{RIS}}\left[m\right]\in\mathbb{C}^{1\times N_{\mathrm{T}}}$ from BS to UE-$k$ at the $m$-th SC can be expressed as
-$\begin{split}\mathbf{H}_{k}^{\mathrm{RIS}}\left[m\right]&=\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}\right)\mathbf{G}\left[m\right]\\ &=\boldsymbol{\theta}\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right],\end{split}$ (1)
-where $\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right]\triangleq\mathbf{G}\left[m\right]\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ is the RIS-assisted cascaded channel and $\boldsymbol{\theta}=\left[e^{j\theta_{1}},\cdots,e^{j\theta_{N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS.
+Additionally, we assume that the direct BS-UE links are blocked by obstacles [20].
+Therefore, we mainly focus on the downlink channel estimation of the cascaded BS-RIS-UE channel.
 As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[15].
-As mentioned above, we consider that $P$ frequency-dependent baseband beamformers $\mathbf{W}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}\triangleq\mathrm{Diag}\left\{s_{1},\cdots,s_{P}\right\}\in\mathbb{C}^{P\times P}$ per time slot.
-Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain augment in each user direction, wherein each entry is subject to the unit-modulus constraint, i.e., $|\left[\mathbf{W}^{\mathrm{RF}}\right]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ as $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$.
-Define $\mathbf{W}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}[m]=\left[\mathbf{w}_{1}[m],...,\mathbf{w}_{P}[m]\right]$.
-At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by UE-$k$ at the $m$-th SC is denoted as
-$\begin{split}y_{k,t,p}[m]&=\mathbf{\theta}_{t}\mathbf{H}_{k}^{\mathrm{cas}}[m]\mathbf{w}_{t,p}[m]s_{t,p}[m]+n_{k,t,p}[m],\end{split}$ (2)
+In our considered system, we assume that $P$ frequency-dependent baseband beamformers $\mathbf{W}_{t}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{t,1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}_{t}\triangleq\mathrm{Dia{\tiny}g}\left\{s_{t,1},\cdots,s_{t,P}\right\}\in\mathbb{C}^{P\times P}$ per time slot.
+Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain improvement in the direction of RIS with each entry is subject to the unit-modulus constraint, i.e., $|[\mathbf{W}^{\mathrm{RF}}]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ for $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$.
+We assume that the analog beamformer remains unchanged over multiple channel coherence blocks due to the relatively static nature of the RIS physical location compared to the dynamic channel coefficients.
+Therefore, we define $\mathbf{W}_{t}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}_{t}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}_{t}[m]=\left[\mathbf{w}_{t,1}[m],...,\mathbf{w}_{t,P}[m]\right]$.
+Let $\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ and $\mathbf{h}_{k}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}}$ denote the channel from BS to RIS and RIS to $k$-th UE at the $m$-th SC, respectively.
+At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by $k$-th UE at the $m$-th SC can be expressed as
+$\begin{split}y_{k,t,p}[m]\!&=\!\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}_{t}\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m]\\ &=\!\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m],\end{split}$ (1)
 where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN).
+Besides, $\boldsymbol{\theta}_{t}=\left[e^{j\theta_{t,1}},\cdots,e^{j\theta_{t,N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS at subframe $t$ with a constant reflection amplitude constraint [24].
 Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t,p}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
-$\tilde{\mathbf{y}}_{k,t}[m]=\boldsymbol{\theta}_{t}\mathbf{H}_{k}^{\mathrm{cas}}[m]\mathbf{R}_{t}[m]+\tilde{\mathbf{n}}_{k,t}[m].$ (3)
-We adopt the widely used Saleh-Valenzuela channel model [21] for BS-RIS channel $\mathbf{G}[m]$ and RIS-UE channel $\mathbf{h}_{k}[m]$.
-The BS-RIS channel $\mathbf{G}[m]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ at the $m$-th SC is expressed as
+$\tilde{\mathbf{y}}_{k,t}[m]=\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{R}_{t}[m]+\tilde{\mathbf{n}}_{k,t}[m],$ (2)
+where $\tilde{\mathbf{n}}_{k,t}[m]\triangleq\left[n_{k,t,1}[m],...,n_{k,t,P}[m]\right]\in\mathbb{C}^{1\times P}$ denotes the combined AWGN. Therefore, the cascaded channels can be expressed as [25]
+$\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right]=\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}.$ (3)
+We adopt the widely used Saleh-Valenzuela channel model for both BS-RIS channel and RIS-UE channel [26]. The BS-RIS channel $\mathbf{G}[m]$ at the $m$-th SC is expressed as
 $\mathbf{G}[m]=\sum_{l=1}^{L}{\alpha_{l}\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{l})\mathbf{a}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\varphi}_{l}\right)e^{-j2\pi\tau_{l}f_{m}}},$ (4)
@@
 $f_{m}=f_{c}+\frac{B}{M}\left(m-1-\frac{M-1}{2}\right),\forall m=1,2,\dots,M,$ (5)
-where $c$ is the speed of light.
+where $f_{c}$ is the central carrier frequency.
 Besides, $\mathbf{a}_{\mathrm{T},m}\in\mathbb{C}^{N_{\mathrm{T}}}$ and $\mathbf{a}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}}$ are the normalized array response vectors (ARVs) associated with BS and RIS, respectively, given by
-$\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi})=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{\varphi}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
-$\mathbf{a}_{\mathrm{R},m}(\tilde{\psi})=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{\psi}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
+$\mathbf{a}_{\mathrm{T},m}(\tilde{u})\triangleq\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{u}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
+$\mathbf{a}_{\mathrm{R},m}(\tilde{v})\triangleq\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{v}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
 where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th suncarrier.
 $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively.
-Besides, $\mathbf{n}_{t}=\left[0,1,\dots,N_{\mathrm{T}}-1\right]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively.
-By defining the angle vectors $\tilde{\boldsymbol{\psi}}=\left[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}\right]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as
-$\mathbf{G}[m]=\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\boldsymbol{\varphi}}\right),$ (7)
-where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=\left[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})\right]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively.
-$\mathbf{\Upsilon}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix.
+Besides, $\mathbf{n}_{t}=[0,1,\dots,N_{\mathrm{T}}-1]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively.
+By defining the angle vectors $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as
+$\mathbf{G}[m]=\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\boldsymbol{\varphi}}\right),$ (7)
+where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively.
+$\mathbf{\Sigma}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix.
 Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and UE-$k$ at the $m$-th SC can be formulated as
@@
 where $J_{k}$ represents the number of propagation paths between RIS and UE-$k$, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the AoD at RIS for the $j$-th path, respectively.
-By defining $\tilde{\boldsymbol{\vartheta}}_{k}=\left[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}\right]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form, given by
+By defining $\tilde{\boldsymbol{\vartheta}}_{k}=[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form as
 $\mathbf{h}_{k}[m]=\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}),$ (9)
@@
 $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the AoDs at RIS.
-Plugging (7) and (9) into (1), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be written as
-$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Upsilon}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
+Substituting (7) and (9) into (3), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be rewritten as
+$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
 II-B Beam Split Effect
-In RIS-assisted wideband THz systems, beam split effect is essentially caused by the analog beamformer and RIS, which contain the frequency-independent phase shifters [22].
-The antenna spacing $d_{\mathrm{BS}}$ and $d_{\mathrm{RIS}}$ are set according to the central frequency $f_{c}$, with half-wavelength spacing, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$ as a typical configuration, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [23].
+According to (4) and (8), THz channel varies across different SCs due to the beam split effect [^1] The beam split effect can been seen as a severe version of beam squint effect in mmWave band [27]..
+This phenomenon is caused by the ultra-wide bandwidth and extensive antenna arrays and reflecting elements in RIS-assisted wideband THz systems.
+Beams generated by the frequency-independent analog beamforming vector and then reflected by the frequency-independent RIS phase shift matrix would split into totally distinct spatial directions.
+Typically, antenna spacings are set to half the wavelength at the central frequency $f_{c}$, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [28].
+For simplicity, we define the variables $u\triangleq\sin\tilde{u}\in[-1,1]$ and $v\triangleq\sin\tilde{v}\in[-1,1]$ to represent the physical directions.
 Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
-$\mathbf{a}_{\mathrm{T},m}(\varphi)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}\varphi\mathbf{n}_{t}},$ (11a)
-$\mathbf{a}_{\mathrm{R},m}(\psi)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}\psi\mathbf{n}_{r}},$ (11b)
-where $\eta_{m}=f_{m}/f_{c}$ represents the relative frequency.
-Without loss of generality, we define $\varphi=\sin\tilde{\varphi}\in[-1,1]$ and $\psi=\sin\tilde{\psi}\in[-1,1]$ as the physical direction for expression simplicity.
-In THz band, the physical directions $\varphi$ and $\psi$ split into totally separated spatial directions $\eta_{m}\varphi$ and $\eta_{m}\psi$, which is the so-called beam split effect.
-Due to the large bandwidth and enormous number of RIS elements, beam split effect occurs at both the BS and RIS, making channel estimation algorithms suffer from severe performance degradation [24].
-III Full CSI Estimation by CBS-GAMP
-In this section, we propose a novel CBS-GAMP based channel estimation approach where simplified CBS dictionaries are firstly presented to formulate angular domain channel.
-By employing CBS dictionaries, CBS-GAMP algorithm with EM extensions is implemented alternatively for each angular sub-channel to achieve accurate channel estimation.
-III-A Angular Domain Sparse Representation
-In this subsection, we transformed the THz channel to the angular domain sparse representation by utilizing the overcomplete CBS dictionaries.
-Specifically, the CBS dictionaries of the BS and RIS directions are formulated as
+$\mathbf{a}_{\mathrm{T},m}(u)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}u\mathbf{n}_{t}},$ (11a)
+$\mathbf{a}_{\mathrm{R},m}(v)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}v\mathbf{n}_{r}},$ (11b)
+where $\eta_{m}\triangleq f_{m}/f_{c}$ represents the relative frequency. As shown in (11), the spatial directions can be represented by
+$u_{m}=\eta_{m}u,$ (12a)
+$v_{m}=\eta_{m}v.$ (12b)
+In the lower frequency bands such as mmWave, the subcarrier frequencies are relatively closer, which causes slight beam squint with $\eta_{m}\approx 1$ which means that the ARVs at all SC frequencies are almost identical[28, 20].
+However, due to the ultra large bandwidth in THz band, $\eta_{m}$ deviates significantly from one.
+Therefore, the difference between physical directions and spatial directions becomes significant and cannot be ignored.
+In wideband RIS-assisted THz systems, beam split effect occurs at both BS and RIS, which leads to severe performance degradation in channel estimation algorithms [27].
+Therefore, it is crucial to develop an effective channel estimation scheme to acquire accurate CSI under beam split effect.
+III Channel Estimation
+III-A Channel Estimation Protocol
+Traditional channel estimation algorithms require significant redesign of the codebook, as conventional dictionaries fail to cover additional angle information introduced by beam split effect [17].
+This additional angle information will complicates the design of dictionaries.
+Furthermore, the beam split effect disperses the channel for each subcarrier into different spatial directions, which requires well-designed channel estimation algorithms to operate on each subcarrier individually, which can significantly increase computational complexity.
+To address these challenges, we propose a novel low-complexity channel estimation protocol, which is divided into two main phases, as depicted in Fig.
+1.
+The first phase, which is the most challenging aspect of the protocol, mainly focuses on estimating the full CSI for two selected subcarriers, which includes estimation of cascaded channel, the AoDs at BS, and the coupled angles at RIS.
+Due to its complexity, we further divide this phase into two steps.
+In Step 1, we employ sparse recovery algorithms to estimate the cascaded channel for the two subcarriers.
+In Step 2, we focus on heuristic algorithms for common angles estimation.
+Given the full CSI estimation from the first phase, we then employ a simple LS algorithm to estimate the channels for the remaining subcarriers.
+III-B Cascaded Channel Estimation
+In this subsection, we focus on the estimation of cascaded channel for one specific subcarrier.
+In the following steps, we first present the angular domain sparse representation in RIS-assisted THz channels.
+Then, we introduce the simplified dictionary, i.e., CBS dictionary.
+Based on this simplified dictionary, we reformulate the channel estimation problem as a sparse recovery problem.
+Finally, we provide a detailed analysis of existing sparse recovery algorithms.
+III-B1 Angular Domain Sparse Representation
+The sparse angular domain representation of the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ is given by
+$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
+where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS-RIS channel and the angle-excluded vector of RIS-UE channel, respectively.
+Additionally, the overcomplete dictionaries of the BS and RIS directions can be expressed as
 $\mathbf{C}_{\mathrm{T},m}=\left[\mathbf{a}_{\mathrm{T},m}\left(\bar{\vartheta}_{1}\right),\cdots,\mathbf{a}_{\mathrm{T},m}\left(\bar{\vartheta}_{Q_{\mathrm{T}}}\right)\right],$ (14a)
@@
 where $Q_{\mathrm{T}}\gg N_{\mathrm{T}}$ and $Q_{\mathrm{R}}\gg N_{\mathrm{R}}$ are the quantization level of the BS and RIS directions, respectively.
-Besides, $\bar{\vartheta}_{q}=\delta q-\frac{Q+1}{Q}$ with $Q\in\{Q_{\mathrm{T}},Q_{\mathrm{R}}\}$ denotes the spatial direction sample, $\delta=2/Q$ is the interval between two adjacent samples.
-Substituting the ARMs in (10) by the CBS dictionaries and extending the angle-excluded matrix/vector, the sparse angular domain representation of the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ is given by
-$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Upsilon}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
-where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Upsilon}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS and the angle-excluded vector of RIS, respectively.
-According to the operation of transposed Khatri-Rao product and the uniform sampling interval of the CBS dictionary, the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains a considerable amount of duplicated columns.
-For narrowband channel model, the coupled dictionary contains only $Q_{\mathrm{R}}$ distinct columns, which are exactly the first $Q_{\mathrm{R}}$ columns [6].
-However, for wideband channel model with beam split effect, due to the frequency-dependent nature of ARVs, this simplification is no longer suitable.
-Note that sequence $\bar{\vartheta}_{q_{1}}=\delta q_{1}-\frac{Q+1}{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ corresponds one-to-one with $Q_{\mathrm{R}}$ ARVs in dictionary $\mathbf{C}_{\mathrm{R},m}$, the sequence $-\bar{\vartheta}_{q_{2}},\forall q_{2}=1,\dots,Q_{\mathrm{R}}$ can also establish such correspondence with $\mathbf{C}_{\mathrm{R},m}^{*}$.
-Therefore, it can be observed that the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ have the form of $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}}),\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
-The $j$-th column of the coupled dictionary $\tilde{{\xi}}_{\mathrm{R},m,j}$ can be expressed as
-$\tilde{\xi}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (14)
-where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
-Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of distinct columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}\triangleq 2Q_{\mathrm{R}}-1$.
-By employing the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, one can easily construct the coupled CBS dictionary as
+Moreover, $\bar{\vartheta}_{q}=\delta\cdot q-(Q+1)/{Q}$ with $Q\in\{Q_{\mathrm{T}},Q_{\mathrm{R}}\}$ denotes the spatial direction sample, $\delta=2/Q$ is the interval between two adjacent samples.
+III-B2 Simplified Dictionary
+Note that the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ in (13) is derived from the transposed Khatri-Rao product, which contains a significant number of redundant columns due to the fixed sampling intervals.
+For narrowband channel model, this will result in only $Q_{\mathrm{R}}$ unique columns which are exactly the first $Q_{\mathrm{R}}$ columns [9].
+However, for wideband THz channel with beam split effect, this simplification is no longer suitable due to the frequency-dependent nature of the ARVs.
+Proposition 1.
+The number of unique columns of dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{B}=2Q_{R}-1$, and the coupled dictionary can be simplified by including only the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, which can be given as
 $\mathbf{\Xi}_{\mathrm{R},m}\triangleq[\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(Q_{\mathrm{R}}^{2}-Q_{\mathrm{R}}+1:Q_{\mathrm{R}}^{2}\right);\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(2:Q_{\mathrm{R}}\right)],$ (15)
 where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary.
+Figure 3: Coupled dictionary simplification diagram.
+Proof.
+We first define the parameter sequence $\bar{\vartheta}_{q_{1}}=\delta\cdot q_{1}-(Q+1)/{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ which corresponds one-to-one with $Q_{\mathrm{R}}$ ARV columns in the dictionary $\mathbf{C}_{\mathrm{R},m}$.
+Similarly, we define the sequence $\delta_{q_{2}}=-\delta\cdot q_{2}+(Q+1)/Q$ to correspond to $Q_{R}$ ARV columns in the conjugate dictionary $\mathbf{C}_{\mathrm{R},m}^{*}$.
+Thus, the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ can be expressed as $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}})$ for $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
+The $j$-th column of the coupled dictionary $\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}$ can be expressed as
+$\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (16)
+where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
+Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of unique columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$.
+Then, we simplify the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ step by step from (a) to (d) as shown in Fig.
+3.
+As depicted in step (a), we reshape the overcomplete dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ of length $Q_{\mathrm{R}}^{2}$ ARV columns into a three-dimensional matrix where each column represents an ARV.
+Given that ARV vectors can be uniquely determined by known angles in (16), we use the sequences $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$ to index the corresponding ARV columns.
+In step (b), we form the index matrix $\mathbf{Q}\in\mathbb{N}^{Q_{\mathrm{R}}\times Q_{\mathrm{R}}}$ with elements $[\mathbf{Q}]_{i,j}=q_{1}(i)-q_{2}(j)$ for $i,j=1,\dots,Q_{\mathrm{R}}$.
+It is noticed that due to the cyclic shift property of this dictionary, the dictionary base can be represented entirely by combining just the first and last rows.
+This means that we can reconstruct the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ with a sliding window of length $Q_{\mathrm{R}}$.
+Therefore, the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ can be given as in (15).
 Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (16), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
-$\xi_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (16)
-By substituting the simplified $\mathbf{\Xi}_{\mathrm{R},m}$ into (13), the RIS-assisted cascaded channel can be rewritten as
+$\boldsymbol{\xi}_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (17)
+III-B3 Sparse Recovery Reformulation
+By substituting the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ into (13), the RIS-assisted cascaded channel can be rewritten as
 $\mathbf{H}_{k}^{\mathrm{cas}}[m]=\mathbf{\Xi}_{\mathrm{R},m}\mathbf{X}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (18)
-where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Upsilon}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$.
-$\mathcal{P}_{i}$ is the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ which are identical to the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$.
-Therefore, there are maximum $\Sigma_{k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$.
-Employing the property of the Kronecker product, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is given by
+where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$.
+$\mathcal{P}_{i}$ denotes the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ that match the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$.
+Therefore, there are at most ${L}_{\mathrm{tot},k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$.
+Employing the Kronecker product property, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ can be given by
 $\mathbf{h}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\mathbf{x}_{k,m},$ (19)
@@
 where $\mathbf{y}_{k,t}\left[m\right]\in\mathbb{C}^{P}$ is the vector form of $\tilde{\mathbf{y}}_{k,t}[m]$.
-Stacking the measurement vectors collected at $T$ successive subframes $\mathbf{y}_{k}\left[m\right]=\left[\mathbf{y}_{k,1}^{\mathsf{T}}\left[m\right],\cdots,\mathbf{y}_{k,T}^{\mathsf{T}}\left[m\right]\right]^{\mathsf{T}}\in\mathbb{C}^{PT}$, the received measurement signal can be expressed as
+Stacking the measurement vectors collected at $T$ successive subframes $\mathbf{y}_{k}\left[m\right]=[\mathbf{y}_{k,1}^{\mathsf{T}}\left[m\right],\cdots,\mathbf{y}_{k,T}^{\mathsf{T}}\left[m\right]]^{\mathsf{T}}\in\mathbb{C}^{PT}$, the received signal can be expressed as
 $\mathbf{y}_{k}[m]=\mathbf{\Psi}_{m}\mathbf{x}_{k,m}+\mathbf{n}_{k}\left[m\right],$ (21)
-where $\mathbf{\Psi}_{m}=\overline{\mathbf{R}}_{m}\mathbf{\Pi}_{m}$ and
-$\overline{\mathbf{R}}_{m}\triangleq\left[\begin{array}[]{c}\mathbf{R}_{1}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{1}\\ \vdots\\ \mathbf{R}_{T}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{T}\\ \end{array}\right].$ (21)
-It is obvious that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix.
-So far, by employing the CBS dictionaries, the cascaded channel estimation problem has been transformed into the sparse signal recovery problem.
-Our primary task is to reconstruct the total angle-excluded vectors $\mathbf{x}_{k,m}$ with knowledge of the measurement vectors $\mathbf{y}_{k}[m]$ and the observation matrices $\mathbf{\Psi}_{m}$.
-III-B Accurate Sparse Recovery and Channel Reconstruction
-The complex sparse recovery problem has been formulated in Subsection III-A.
+where $\mathbf{\Psi}_{m}=\bar{\mathbf{R}}_{m}\mathbf{\Pi}_{m}$ and
+$\bar{\mathbf{R}}_{m}\triangleq\left[\begin{array}[]{c}\mathbf{R}_{1}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{1}\\ \vdots\\ \mathbf{R}_{T}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{T}\\ \end{array}\right].$ (22)
+Note that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix.
 To map this problem to the real number domain for resolution, we then transform the complex form in (21) to the equivalent real form
 [$\left[\begin{array}[]{c}\!\!\!\Re\{\mathbf{y}_{k}\left[m\right]\}\!\!\!\\ \!\!\!\Im\{\mathbf{y}_{k}\left[m\right]\}\!\!\!\\ \end{array}\right]\!\!\!=\!\!\!\left[\begin{matrix}\Re\{\mathbf{\Psi}_{m}\}&\!\!\!\!-\Im\{\mathbf{\Psi}_{m}\}\\ \Im\{\mathbf{\Psi}_{m}\}&\!\!\!\!\Re\{\mathbf{\Psi}_{m}\}\\ \end{matrix}\right]\!\!\left[\begin{array}[]{c}\!\!\!\!\Re\{\mathbf{x}_{k,m}\}\!\!\!\!\\ \!\!\!\!\Im\{\mathbf{x}_{k,m}\}\!\!\!\!\\ \end{array}\right]\!+\!\left[\begin{array}[]{c}\!\!\!\Re\{\mathbf{n}_{k}\left[m\right]\}\!\!\!\!\!\\ \!\!\!\Im\{\mathbf{n}_{k}\left[m\right]\}\!\!\!\!\!\\ \end{array}\right]\!\!,$ (23)
-By systematically assigning symbols to each matrix in (22), the sparse recovery problem is formulated as
+By systematically assigning symbols to each matrix in (23), the sparse recovery problem can be reformulated as
 $\boldsymbol{y}=\mathbf{\Psi}\boldsymbol{x}+\boldsymbol{n}.$ (24)
-For simplicity, we employ a simplified subscript notation by omitting the index for the matrix and vectors.
-By defining $T_{\mathrm{tot}}\triangleq PT$ and $Q_{\mathrm{tot}}\triangleq Q_{\mathrm{B}}Q_{\mathrm{T}}$, $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$ is the measurement vector, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ denotes the observation matrix and $\boldsymbol{x}\boldsymbol{\in}\mathbb{C}^{Q_{\mathrm{tot}}}$ with sparsity $\Sigma_{\mathrm{tot}}$ denotes the sparse representation vector to be estimated.
-Note that unlike conventional THz channel models where $\Sigma_{\mathrm{tot}}=\Sigma_{k},\forall k=1,\dots,K$, the sparsity of RIS-assisted THz channel satisfies $\Sigma_{\mathrm{tot}}\leq\Sigma_{k}$ and remains uncertain.
-This is because the introduction of simplified dictionary and RIS phase-shift matrix, resulting in coupling between AoAs and AoDs, which can be challenging to be separated.
-In this case, we adopt the Bayesian framework, i.e.
-GAMP approach, to recover the angle-excluded coefficients in $\boldsymbol{x}$, which do not require the sparsity level $\Sigma_{\mathrm{tot}}$ as the prior input.
-Bayesian approaches require only the prior distribution assumption of $\boldsymbol{x}$, where Gaussian distribution model is widely used in rich-scattering environments [25].
-However, this assumption becomes inapplicable for THz band due to the existence of angular domain sparsity makes $\boldsymbol{x}$ a sparse vector with $Q_{\mathrm{tot}}\gg\Sigma_{\mathrm{tot}}$.
-Since the sparse vector $\mathbf{x}\in\mathbb{C}^{Q^{\rm tot}}$ is supported on a subspace with dimension smaller than $Q^{\rm tot}$, a mixture stochastic distribution can be employed to capture its sparsity [26, 27]:
-$p_{X}(x)=\lambda f_{X}(x)+(1-\lambda)\delta(x),$ (24)
-where $\delta(\cdot)$ is the Dirac delta, $f_{X}(x)$ is the active-coefficient probability density function (PDF), and $\lambda$ is the sparsity rate that parameterizes the signal sparsity.
-By the weak law of large numbers [28], $\left\|\boldsymbol{x}\right\|_{0}/Q_{\mathrm{tot}}\xrightarrow{\mathbb{P}}\lambda$, where the $\ell_{0}$-norm $\left\|\cdot\right\|_{0}$ denotes the number of nonzero entries of a vector.
-Then, we adopt the Gaussian distribution as an active-coefficient PDF in (24) which is widely employed as a priori distribution [29, 25], i.e., $f_{X}(x)=\mathcal{N}(x,u,v)$, the mixture stochastic distribution in (24) can be rewritten as
-$p_{X}(x)=\lambda\mathcal{N}(x,u,v)+(1-\lambda)\delta(x),$ (25)
-where $u$ and $\nu$ are the mean and the variance, respectively, which correspond to the large-scale fading coefficient that accounts for the combined effects of shadowing and path loss.
-Inspired by (25), we treat the sparse vector as Bernoulli-Gaussian (BG) signal with unknown priori $\boldsymbol{q}\triangleq\left[\lambda,u,v\right]$ and variance $\varepsilon$ of measurement noise.
-In the Bayesian estimation framework, we employ the widely adopted minimum mean squared error (MMSE) estimation for each $x_{j}$ separately by approximating the marginal posterior distributions $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y}\right)$ [30]
-$\hat{x}_{j}^{\mathrm{MMSE}}=\mathbb{E}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y}\right\}.$ (26)
-We then employ the GAMP based framework to find the true marginal posterior distributions $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y}\right)$, which is summarized in Algorithm 1.
-Algorithm 1 GAMP for MMSE Estimation
-0:  Observation matrix $\mathbf{\Psi}\in\mathbb{R}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$, measurement vector $\boldsymbol{y}\in\mathbb{R}^{T_{\mathrm{tot}}}$, noise variance $\varepsilon$, prior distribution $p_{X}(x;\boldsymbol{q})$, precision tolerance $\tau_{\mathrm{max}}$ and maximum number of iterations $N_{\mathrm{iter}}$.
-1:  $n=1$.
-2:  $\forall j:\hat{x}_{j}(n)=\int_{x}{xp_{X}(x;\boldsymbol{q})}$.
-3:  $\forall j:\mu_{j}^{x}(n)=\int_{x}{\left|x-\hat{x}_{j}(n)\right|^{2}p_{X}(x;\boldsymbol{q})}$.
-4:  $\forall i:\widehat{s_{i}}(n-1)=0$.
-5:  while $n\leq N_{\mathrm{iter}}$ do
-6:     $\forall i:\mu_{i}^{p}(n)=\sum_{j=1}^{Q_{\mathrm{tot}}}{\left|\mathbf{\Psi}_{i,j}\right|^{2}}\mu_{j}^{x}(n)$.
-7:     $\forall i:\hat{p}_{i}(n)=\sum_{j=1}^{Q_{\mathrm{tot}}}{\mathbf{\Psi}_{i,j}}\hat{x}_{j}(n)-\mu_{i}^{p}(n)\hat{s}_{i}(n-1)$.
-8:     $\forall i:\mu_{i}^{z}(n)=\mathrm{Var}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i}(n),\mu_{i}^{p}(n),\boldsymbol{q}\right\}$.
-9:     $\forall i:\hat{z}_{i}(n)=\mathbb{E}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i}(n),\mu_{i}^{p}(n),\boldsymbol{q}\right\}$.
-10:     $\forall i:\mu_{i}^{s}(n)=\frac{1}{\mu_{i}^{p}(n)}\left[1-\frac{\mu_{i}^{z}(n)}{\mu_{i}^{p}(n)}\right]$.
-11:     $\forall i:\hat{s}_{i}(n)=\frac{1}{\mu_{i}^{p}(n)}\left[\hat{z}_{i}(n)-\hat{p}_{i}(n)\right]$.
-12:     $\forall j:\mu_{j}^{r}(n)=\left(\sum_{i=1}^{T_{\mathrm{tot}}}{\left|\mathbf{\Psi}_{i,j}\right|^{2}}\mu_{i}^{s}(n)\right)^{-1}$.
-13:     $\forall j:\hat{r}_{j}(n)=\hat{x}_{j}(n)+\mu_{j}^{r}(n)\sum_{i=1}^{T_{\mathrm{tot}}}{\mathbf{\Psi}_{i,j}^{*}}\hat{s}_{i}(n)$.
-14:     $\forall j:\hat{x}_{j}(n+1)=\mathbb{E}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j}(n),\mu_{j}^{r}(n),\boldsymbol{q}\right\}$.
-15:     $\forall j:\mu_{j}^{x}(n+1)=\mathrm{Var}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j}(n),\mu_{j}^{r}(n),\boldsymbol{q}\right\}$.
-16:     $n\leftarrow n+1$.
-17:     if $\frac{\sum_{j=1}^{Q_{\mathrm{tot}}}{\left|\hat{x}_{j}(n+1)-\hat{x}_{j}(n)\right|^{2}}}{\sum_{j=1}^{Q_{\mathrm{tot}}}{\left|\hat{x}_{j}(n)\right|^{2}}}<\tau_{\mathrm{max}}$ then
-18:        break.
-19:     end if
-20:  end while
-20:  MMSE estimation $\left\{\hat{x}_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$.
-The relationship between the $i$-th measured output $y_{i}$ and the corresponding noiseless output $z_{i}\triangleq\boldsymbol{\psi}_{i}^{\mathsf{T}}\boldsymbol{x}$ in GAMP algorithm can be modeled by the conditional PDF $p_{Y|Z}\left(y_{i}|z_{i};\boldsymbol{q}\right)$, where $\boldsymbol{\psi}_{i}^{\mathsf{T}}$ is the $i$-th row of $\mathbf{\Psi}$.
-The message from the factor node $p\left(z_{i}|\boldsymbol{\psi}_{i}^{\mathsf{T}}\boldsymbol{x}\right)$ to the variable node $z_{i}$ is calculated by integrating all the variable nodes $\left\{x_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ in $\boldsymbol{x}$.
-As $Q_{\mathrm{tot}}$ is quite large and $x_{j}$ is assumed independent and identically distributed (i.i.d.), according to the central limit theorem (CLT), the conditional distribution $z_{i}$ given $x_{j}$ can be approximated by a Gaussian distribution $\mathcal{N}\left(z_{i};\hat{p}_{i},\mu_{i}^{p}\right)$ with mean $\hat{p}_{i}$ and variance $\mu_{i}^{p}$.
-Therefore, the true marginal posterior $p_{Z|\boldsymbol{Y}}\left(z_{i}|\boldsymbol{y};\boldsymbol{q}\right)$ can be expressed as
-$p_{Z|\boldsymbol{Y}}\left(z_{i}|\boldsymbol{y};\hat{p}_{i},\mu_{i}^{p},\boldsymbol{q}\right)\triangleq\frac{p_{Y|Z}\left(y_{i}|z_{i};\boldsymbol{q}\right)\mathcal{N}\left(z_{i};\hat{p}_{i},\mu_{i}^{p}\right)}{\int_{z}{p_{Y|Z}}\left(y_{i}|z;\boldsymbol{q}\right)\mathcal{N}\left(z;\hat{p}_{i},\mu_{i}^{p}\right)},$ (27)
-where the quantities $\hat{p}_{i}$ and $\mu_{i}^{p}$ are updated with iteration $n$ in lines $7$ and $6$ of Algorithm 1, respectively.
-Since the noise $\boldsymbol{n}$ is treated as AWGN, the conditional probability $p_{Y|Z}\left(y_{i}|z_{i};\boldsymbol{q}\right)$ follows Gaussian distribution $\mathcal{N}\left(y_{i};z_{i},\varepsilon\right)$.
-Therefore, the PDF in (27) has moments
-$\hat{z}_{i}=\mathbb{E}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i},\mu_{i}^{p},\boldsymbol{q}\right\}=\frac{\mu_{i}^{p}y_{i}+\varepsilon\hat{p}_{i}}{\mu_{i}^{p}+\varepsilon},$ (28a)
-$\mu_{i}^{z}=\mathrm{Var}_{Z|\boldsymbol{Y}}\left\{z_{i}|\boldsymbol{y};\hat{p}_{i},\mu_{i}^{p},\boldsymbol{q}\right\}=\frac{\mu_{i}^{p}\varepsilon}{\mu_{i}^{p}+\varepsilon},$ (28b)
-where the quantities $\hat{z}_{i}$ and $\mu_{i}^{z}$ are updated with iteration $n$ in lines $9$ and $8$ of Algorithm 1, respectively.
-Employing CLT again, the conditional distribution $x_{j}$ given $z_{i}$ can be approximated by a Gaussian distribution $\mathcal{N}\left(x_{j};\hat{r}_{j},\mu_{j}^{r}\right)$ with mean $\hat{r}_{j}$ and variance $\mu_{j}^{r}$.
-Therefore, the true marginal posterior $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{q}\right)$ can be expressed as
-$p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right)\triangleq\frac{p_{X}\left(x_{j};\boldsymbol{q}\right)\mathcal{N}\left(x_{j};\hat{r}_{j},\mu_{j}^{r}\right)}{\int_{x}{p_{X}}(x;\boldsymbol{q})\mathcal{N}\left(x;\hat{r}_{j},\mu_{j}^{r}\right)},$ (29)
-where the quantities $\hat{r}_{j}$ and $\mu_{j}^{r}$ are updated with iteration $n$ in lines $13$ and $12$ of Algorithm 1, respectively.
-By substituting the BG distribution $p_{X}\left(x_{j};\boldsymbol{q}\right)$ into (29) and after simple algebraic operations, we define $\zeta_{j}$ as the integration at the denominator in (29), which is expressed as
-$\begin{split}&\zeta_{j}\triangleq\int_{x}{p_{X}}(x;\boldsymbol{q})\mathcal{N}\left(x;\hat{r}_{j},\mu_{j}^{r}\right)\\ &=(1-\lambda)\mathcal{N}\left(0;\hat{r}_{j},\mu_{j}^{r}\right)+\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right),\end{split}$ (30)
-Similarly, the numerator in (29) can be derived as
-$\begin{split}&p_{X}\left(x_{j};\boldsymbol{q}\right)\mathcal{N}\left(x_{j};\hat{r}_{j},\mu_{i}^{r}\right)=(1-\lambda)\mathcal{N}\left(0;\hat{r}_{j},\mu_{j}^{r}\right)\delta\left(x_{j}\right)\\ &+\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right)\mathcal{N}\left(x_{j};\gamma_{j},\nu_{j}\right),\end{split}$ (31)
-where $\gamma_{j}\triangleq(\hat{r}_{j}\nu+\mu\mu_{j}^{r})/(\mu_{j}^{r}+\nu)$ and $\nu_{j}\triangleq\mu_{j}^{r}\nu/(\mu_{j}^{r}+\nu)$ are the $\left(\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right)$-dependent quantities and the simplification in (30) and (31) holds by the Gaussian-PDF multiplication rule where
-$\mathcal{N}\!(x;a,\!A)\mathcal{N}\!(x;b,\!B)\!=\!\mathcal{N}\!\!\left(\!x;\!\frac{aB\!\!+\!\!bA}{A\!+\!B}\!,\!\frac{AB}{A\!\!+\!\!B}\!\right)\!\mathcal{N}\!(0;\!a-b,\!A+B).$ (32)
-By defining
-$\pi_{j}\triangleq\frac{\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right)}{\lambda\mathcal{N}\left(0;\hat{r}_{j}-\mu,\mu_{j}^{r}+\nu\right)+(1-\lambda)\mathcal{N}\left(0;\hat{r}_{j},\mu_{j}^{r}\right)},$ (33)
-we can further simplify (29) as
-$p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right)=\left(1-\pi_{j}\right)\delta\left(x_{j}\right)+\pi_{j}\mathcal{N}\left(x_{j};\gamma_{j},\nu_{j}\right).$ (34)
-So far, the posterior $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y}\right)$ is rewritten into the form of BG distribution with quantities $\left(\pi_{j},\gamma_{j},\nu_{j}\right)$.
-Wherein $\pi_{j}$ is the sparsity rate corresponing to the posterior support probability $\operatorname{Pr}\left\{x_{j}\neq 0|\boldsymbol{y};\boldsymbol{q}\right\}$, $\gamma_{j}$ and $\nu_{j}$ are the corresponding mean and variance, respectively.
-Given the PDF $p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{q}\right)$ in (34), the moments can be computed in closed form
-$\hat{x}_{j}=\mathbb{E}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j},\mu,\boldsymbol{q}\right\}=\pi_{j}\gamma_{j},$ (35a)
-$\mu_{j}^{x}=\mathrm{Var}_{X|\boldsymbol{Y}}\left\{x_{j}|\boldsymbol{y};\hat{r}_{j},\mu_{j}^{r},\boldsymbol{q}\right\}=\pi_{j}^{2}\nu_{j},$ (35b)
-where the quantities $\hat{x}_{j}$ and $\mu_{j}^{x}$ are updated with iteration $n$ in lines $14$ and $15$ of Algorithm 1, respectively.
-Essentially, the quantities $\left(\pi_{j},\gamma_{j},\nu_{j}\right)$ all depend on the prior parameters $\boldsymbol{q}$ in (25), which are unknown beforehand.
-Therefore, in next subsection, we employ the EM based approach to learn $\boldsymbol{q}$ from the received measurements $\boldsymbol{y}$.
-III-C EM-Based Prior Parameters Learning
-In this subsection, we aim to determine the maximum-likelihood estimate $\boldsymbol{\alpha}\triangleq\left[\boldsymbol{q},\varepsilon\right]$ of the prior signal parameters and noise variance.
-The prior parameters $\boldsymbol{\alpha}$ can be learned and updated by EM approach.
-Specifically, with knowledge of the posterior probabilities acquired by GAMP algorithm in Section III-B, EM algorithm iteratively increases the log-likelihood function (LLF) to guarantee the convergence to a local maximum of the likelihood $p_{\boldsymbol{Y}}(\boldsymbol{y};\boldsymbol{\alpha})$ [31].
-The LLF can be approximated by GAMP posteriors, which yields EM update as
-$\boldsymbol{\alpha}^{n+1}=\mathop{\mathrm{arg}\max}\limits_{\boldsymbol{\alpha}}\mathbb{E}_{\boldsymbol{X|Y}}\left\{\ln p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\alpha})\right\},$ (36)
-where $p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y})$ is the joint PDF parameterized by $\boldsymbol{\alpha}$, which is given by
-$p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\alpha})=p_{\boldsymbol{Y|X}}\left(\boldsymbol{y}|\boldsymbol{x};\varepsilon\right)p_{\boldsymbol{X}}(\boldsymbol{x};\boldsymbol{q}).$ (37)
-Substituting (37) into (36), the expectation can be divided into two parts
-$\begin{split}&\mathbb{E}_{\boldsymbol{X|Y}}\left\{\ln p_{\boldsymbol{XY}}(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\alpha})\right\}\\ &=\!\!\int_{\boldsymbol{x}}\!\!{\ln p_{\boldsymbol{Y\!|X}}\!\!\left(\boldsymbol{y}|\boldsymbol{x};\!\varepsilon\right)\!p_{\boldsymbol{X\!|Y}}\!\!\left(\boldsymbol{x}|\boldsymbol{y};\!\boldsymbol{\alpha}\right)}\!+\!\!\int_{\boldsymbol{x}}\!\!{\ln\!p_{\boldsymbol{X}}\!(\boldsymbol{x};\!\boldsymbol{q})p_{\boldsymbol{X\!|Y}}\!\!\left(\boldsymbol{x}|\boldsymbol{y};\!\boldsymbol{\alpha}\right)}.\end{split}$ (38)
-As observed in (38), the expectation of LLF in (36) is decoupled in terms of the signal parameters $\boldsymbol{q}$ and noise variance $\varepsilon$, which can be updated separately.
-Following the well-established incremental updating schedule proposed in [32], we update only one parameter in $\boldsymbol{\alpha}$ at one time while keeping other parameters unchanged.
-III-C1 EM update for $\lambda$
-Since $\lambda$ is independent of noise variance $\varepsilon$, the update of $\lambda$ is only relevant to the latter part in (38), which is given by
-$\lambda^{n+1}\!=\mathop{\mathrm{arg}\max}\limits_{\lambda\in\left(0,1\right)}\!\sum_{j=1}^{Q_{\mathrm{tot}}}\!{\int_{x_{j}}\!{\ln p_{X}(x_{j};\lambda,\mu^{n},\nu^{n})p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}},$ (39)
-Clearly, $\lambda^{n+1}$ is the value of $\lambda$ where the expectation in (39) is at the point of zero derivative
-$\sum_{j=1}^{Q_{\mathrm{tot}}}{\int_{x_{j}}{p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\lambda}\ln p_{X}(x_{j};\lambda,\mu^{n},\nu^{n})}}=0.$ (40)
-Substituting (25) into (40), one can derive that
-$\frac{d}{d\lambda}\ln p_{X}\left(x_{j};\lambda,\mu^{n},\nu^{n}\right)=\begin{cases}\frac{1}{\lambda-1}&x_{j}=0\\ \frac{1}{\lambda}&x_{j}\neq 0\end{cases},$ (41)
-Given the calculation in (34) and (41), it is previous that the neighborhood around $x_{j}=0$ which we define as $\mathcal{U}_{e}=U(0,e)$ is nonconsecutive from the remainder $\bar{\mathcal{U}}_{e}=\mathbb{R}\backslash\mathcal{U}_{e}$ in (40).
-Therefore, we calculate the integration in (40) within the neighborhood $\mathcal{U}_{e}$ and the remainder $\bar{\mathcal{U}}_{e}$ separately
-$\lambda\!=\!\underset{e\rightarrow 0}{\lim}\sum_{j=1}^{Q_{\mathrm{tot}}}\!{\frac{\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}{\int_{x_{j}\in\bar{\mathcal{U}}_{e}}\!{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\!+\!\int_{x_{j}\in\mathcal{U}_{e}}\!{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (42)
-Substituting the above observation and (34) into (42), the update of $\lambda$ can be easily obtained
-$\lambda^{n+1}=\underset{e\rightarrow 0}{\lim}\sum_{j=1}^{Q^{\mathrm{tot}}}{\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{p_{X|\boldsymbol{Y}}}\left(x_{j}|\boldsymbol{y};\boldsymbol{q}^{n}\right)}=\frac{1}{Q^{\mathrm{tot}}}\sum_{j=1}^{Q^{\mathrm{tot}}}{\pi_{j}},$ (43)
-where $\left\{\pi_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ are computed from GAMP outputs via (33).
-III-C2 EM update for $\mu$
-Similar to (39), the update for $\mu$ is given by
-$\mu^{n+1}\!=\!\underset{\mu\in\mathbb{R}}{\mathrm{arg}\max}\sum_{j=1}^{Q_{\mathrm{tot}}}\!\!{\int_{x_{j}}{\ln p_{X}\left(x_{j};\lambda^{n},\mu,\nu^{n}\right)p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (44)
-Similarly, $\mu^{n+1}$ corresponding to the maximum expectation in (44) is again at the point of zero derivative
-$\sum_{j=1}^{Q_{\mathrm{tot}}}{\int_{x_{j}}{p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\mu}\ln p_{X}\left(x_{j};\lambda^{n},\mu,\nu^{n}\right)}=0}.$ (45)
-Substituting (25) into (45), one can derive that
-$\frac{d}{d\mu}\ln p_{X}\left(x_{j};\lambda^{n},\mu,\nu^{n}\right)=\begin{cases}0&x_{j}=0\\ \frac{x_{j}-\mu}{\nu^{n}}&x_{j}\neq 0\\ \end{cases},$ (46)
-Excluding the neighborhood $\mathcal{U}_{e}$, (45) can be rewritten as
-$\sum_{j=1}^{Q_{\mathrm{tot}}}{\lim_{e\rightarrow 0}\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{\left(x_{j}-\mu\right)}p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}=0.$ (47)
-Then, the unique $\mu$ can be derived from (47) as
-$\mu^{n+1}=\frac{1}{\lambda^{n+1}Q_{\mathrm{tot}}}\sum_{j=1}^{Q_{\mathrm{tot}}}{\pi_{j}}\gamma_{j},$ (48)
-where $\left\{\pi_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ and $\left\{\gamma_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ are computed from GAMP outputs via (34).
-III-C3 EM update for $\nu$
-Similar to (39), the update for $\nu$ is given by
-$\nu^{n+1}\!=\!\underset{\nu>0}{\mathrm{arg}\max}\sum_{j=1}^{Q_{\mathrm{tot}}}\!\!\!{\int_{x_{j}}{\ln p_{X}\left(x_{j};\lambda^{n},\mu^{n},\nu\right)p_{X|\boldsymbol{Y}}\!\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (49)
-Similarly, $\nu^{n+1}$ corresponding to the maximum expectation in (49) is again at the point of zero derivative
-$\sum_{j=1}^{Q_{\mathrm{tot}}}{\int_{x_{j}}{p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\nu}\ln p_{X}\left(x_{j};\lambda^{n},\mu^{n},\nu\right)}=0}.$ (50)
-Substituting (25) into (50), one can derive that
-$\nu^{n+1}=\begin{cases}0&x_{j}=0\\ \frac{1}{2}\left(\frac{\left|x_{j}-\mu^{n}\right|^{2}}{\left(\nu\right)^{2}}-\frac{1}{\nu}\right)&x_{j}\neq 0\\ \end{cases}.$ (51)
-Excluding the neighborhood $\mathcal{U}_{e}$, (50) can be rewritten as
-$\sum_{j=1}^{Q_{\mathrm{tot}}}{\lim_{e\rightarrow 0}\int_{x_{j}\in\bar{\mathcal{U}}_{e}}{\left(\left|x_{j}-\mu^{n}\right|^{2}-\nu\right)}p_{X|\boldsymbol{Y}}\left(x_{j}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}=0.$ (52)
-Then, the unique $\nu$ can be derived from (52) as
-$\nu^{n+1}=\frac{1}{\lambda^{n+1}Q_{\mathrm{tot}}}\sum_{j=1}^{Q_{\mathrm{tot}}}{\pi_{j}\left(\left|\mu^{n}-\gamma_{j}\right|^{2}+\nu_{j}\right)},$ (53)
-where $\left\{\pi_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$, $\left\{\gamma_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ and $\nu_{j}$ are computed from GAMP outputs via (33) and (34).
-III-C4 EM update for $\varepsilon$
-Given $\boldsymbol{q}^{n}$ as determined above, we finally update the noise variance $\varepsilon$.
-Under the AWGN assumption, the noise is i.i.d and independent from $\boldsymbol{x}$.
-Therefore, the update of $\varepsilon$ is only relevant to the former part in (38), which is given by
-$\varepsilon^{n+1}=\underset{\varepsilon>0}{\mathrm{arg}\max}\sum_{i=1}^{T_{\mathrm{tot}}}{\int_{y_{i}}{\ln p_{Y|\boldsymbol{X}}\left(y_{i}|\boldsymbol{x};\varepsilon\right)p_{\boldsymbol{X}|\boldsymbol{Y}}\left(\boldsymbol{x}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)}}.$ (54)
-Recalling that $z_{i}\triangleq\boldsymbol{\psi}_{i}^{\mathsf{T}}\boldsymbol{x}$ and we define $w_{i}\triangleq y_{i}-z_{i}$ as the AWGN variable, the posterior $p_{Y|\boldsymbol{X}}\left(y_{i}|\boldsymbol{x};\varepsilon\right)$ can be re-expressed as
-$p_{Y|\boldsymbol{X}}\left(y_{i}|\boldsymbol{x};\varepsilon\right)=p_{W}\left(w_{i};\varepsilon\right)=\mathcal{N}\left(w_{i};0,\varepsilon\right).$ (55)
-Similarly, $\varepsilon^{n+1}$ corresponding to the maximum expectation in (54) is again at the point of zero derivative
-$\sum_{i=1}^{T_{\mathrm{tot}}}{\int_{w_{i}}{p\left(w_{i}|\boldsymbol{y};\boldsymbol{\alpha}^{n}\right)\frac{d}{d\varepsilon}\ln p_{W}\left(w_{i};\varepsilon\right)}}=0.$ (56)
-Substituting (55) into (56), one can derive that
-$\frac{d}{d\varepsilon}\ln p_{W}\left(w_{i};\varepsilon\right)=\frac{1}{2}\left(\frac{\left|w_{i}\right|^{2}}{(\varepsilon)^{2}}-\frac{1}{\varepsilon}\right).$ (57)
-Then, substituting (57) into (56), the unique $\varepsilon$ can be expressed as
-$\varepsilon^{n+1}=\frac{1}{T_{\mathrm{tot}}}\sum_{i=1}^{T_{\mathrm{tot}}}{\left(\left|y_{i}-\hat{z}_{i}\right|^{2}+\mu_{i}^{z}\right)},$ (58)
-where $\left\{\hat{z}_{i}\right\}_{i=1}^{T_{\mathrm{tot}}}$ and $\left\{\mu_{i}^{z}\right\}_{i=1}^{T_{\mathrm{tot}}}$ are computed from GAMP outputs via (III-B).
-III-C5 EM Initialization
-Since EM approach iteratively converges to a the local maximum of the LLF in (36), we initialize the parameters $\boldsymbol{\alpha}$ as set in [30]
-$\lambda^{0}=\frac{\Sigma_{k}}{Q_{\mathrm{tot}}},\mu^{0}=0,\mathsf{SNR}^{0}=10\mathrm{dB},$ (59a)
-$\varepsilon^{0}=\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}}{\left(\mathsf{SNR}^{0}+1\right)T_{\mathrm{tot}}},\nu^{0}=\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}-T_{\mathrm{tot}}\varepsilon^{0}}{\mathrm{tr}\left(\mathbf{\Psi}^{\mathsf{T}}\mathbf{\Psi}\right)\lambda^{0}}.$ (59b)
-III-D CBS-GAMP Based Full CSI Estimation
-Employing CBS dictionaries, the beam-split affected channel is easily formulated as sparse recovery problem at each SC $m\in\mathcal{M}$.
-In this case, EM and GAMP algorithms are iteratively implemented to alternately update MMSE estimates $\left\{x_{j}\right\}_{j=1}^{Q_{\mathrm{tot}}}$ and prior parameters $\boldsymbol{\alpha}$ at all SCs.
-By substituting MMSE estimates $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$ into (18), the full CSI is available
+For simplicity, we use a simplified subscript notation by omitting the index for the matrices and vectors.
+By defining $T_{\mathrm{tot}}\triangleq PT$ and $Q_{\mathrm{tot}}\triangleq Q_{\mathrm{B}}Q_{\mathrm{T}}$, let $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$ denote the measurement vector, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ denote the observation matrix and $\boldsymbol{x}\in\mathbb{C}^{Q_{\mathrm{tot}}}$ denote the sparse representation vector to be estimated with sparsity ${L}_{\mathrm{tot}}$.
+III-B4 Sparse Recovery Algorithm Analysis
+In designing our channel estimation algorithm, the selection of an appropriate sparse recovery algorithm is crucial due to the unique sparsity characteristics of RIS-assisted THz channels.
+We considered LS, OMP, and GAMP for this purpose.
+a) LS: The LS algorithm, while straightforward, solves the sparse recovery problem by minimizing the residual sum of squares.
+However, its computational complexity, which is $\mathcal{O}(Q_{\mathrm{tot}}^{3})$, will become significantly high for large matrix dimensions.
+Consequently, LS is not practical for the initial full CSI estimation phase due to its high computational demands.
+b) OMP: OMP is a widely used greedy algorithm that reconstructs the signal by iteratively selecting the dictionary atoms that best match the current residual.
+The computational complexity of OMP is $\mathcal{O}(L_{\mathrm{tot}}Q_{\mathrm{tot}}^{2})$, which is relatively efficient.
+However, OMP has a nature of “hard estimation”, which means that it makes definitive decisions at each step without retaining detailed probabilistic information about the signal.
+This approach can lead to suboptimal performance, particularly in noisy environments or with complex signal structures.
+c) GAMP: In contrast to OMP, GAMP uses a probabilistic approach within a message-passing framework to perform global updates, which will maintain a distribution of possible values for each signal element.
+The computational complexity of GAMP is $\mathcal{O}(Q_{\mathrm{tot}}^{2})$ which is similar to OMP but involves more sophisticated probabilistic calculations.
+GAMP’s “soft estimation” process retains probabilistic information throughout the iterations, which leads to more accurate estimation.
+This makes GAMP particularly advantageous in environments with high noise levels or complex signals.
+Despite the similar computational complexity of GAMP and OMP, GAMP’s superior accuracy and robustness make it the preferred choice for initial full CSI estimation [^2] Detailed investigation of CBS-based GAMP scheme has been discussed in our previous work [19].
+Additionally, due to the probabilistic model employed by the “soft estimation” process, GAMP can ensure more reliable angular information determination and avoid the errors caused by “hard estimation”.
+Moreover, for the final stage of channel estimation involving the remaining subcarriers, we employ the LS algorithm.
+The angular information obtained from the initial estimation phase significantly reduces the dictionary matrix dimensions, which will make LS computationally feasible and effective.
+This stage can benefit more from the simplicity and accuracy of LS method.
+III-C Common Angle Estimation
+In this subsection, we focus on the estimation of common angles.
+Using the selected GAMP algorithm from Section III-B, we can obtain the sparse signal vector $\boldsymbol{x}$, denoted as $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$.
+By substituting $\hat{\mathbf{x}}_{k,m}$ into (19), the estimated cascaded channel can be expressed as
 $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m}.$ (25)
-IV Low-Complexity Channel Estimation by BSAD
-The CBS-GAMP algorithm obtained from Section III-B ensures the accuracy of channel estimation.
-However, this algorithm may introduce additional computational complexity due to its execution at each SC.
-It is obvious that the frequency-dependent spatial directions affected by beam split effect share a common physical support.
-Inspired by this observation, we propose a novel BSAD scheme to reduce computational complexity.
-In our proposed algorithm, we only employ the CBS-GAMP to acquire the full CSI within a small subset of SCs, namely the first and the $M/2$-th SCs [^1] The selection of SC indices is demonstrated in Section IV-B.
-Then, employing the obtained full CSI of the subset of SCs, we design EnM algorithm and DS-MUSIC algorithm to determine the common support at BS and RIS, respectively.
-Finally, employing the obtained common support and the frequency-dependent nature introduced by beam split effect, we transform the channel estimation problem at the remaining SCs into a simple linear regression problem, which can be easily solved using LS method.
-IV-A EnM Based DoDs Estimation at BS
-In this subsection, we derive the DoDs at BS from the full CSI $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[\bar{m}],\bar{m}\in\left\{1,M/2\right\}$ estimated by the CBS-GAMP algorithm.
-Recalling (60), the estimated full CSI is consisted of two parts, i.e., the total CBS dictionary $\mathbf{\Pi}_{\bar{m}}$ and the angle-excluded vector $\mathbf{x}_{k,\bar{m}}$, which are known separately.
-Based on the fact that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\mathbf{x}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$, the common support can be easily determined.
-However, the inevitable noise and estimation errors will cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, leading to more than $\Sigma_{k}$ non-zero entries.
-Therefore, we propose a novel EnM method to determine the common support $\zeta^{\mathrm{BS}}$ and then calculate the corresponding physical DoDs at BS, as summarized in Algorithm 2.
-Algorithm 2 EnM Based DoDs Estimation at BS
-0:  Angle-excluded matrix $\mathbf{x}_{k,\bar{m}},\bar{m}\in\left\{1,M/2\right\}$, the path number $L$ of BS-RIS channel and the quantization level of CBS dictionaries $Q_{\mathrm{R}}$ and $Q_{\mathrm{T}}$.
+This method implies that we could perform the estimation for each subcarrier independently to obtain the channel estimates for all subcarriers.
+However, this requires implementing the estimation algorithm $M$ times, which would result in an unacceptably high computational complexity.
+While the beam split effect complicates the channel estimation process, it also offers some insights.
+It is noticed in (11) that different subcarriers share common physical directions, with the only different term $\eta_{m}$.
+This indicates that the physical angles remain constant across different subcarriers.
+Inspired by this, we perform the cascaded channel estimation on only a small subset of subcarriers to acquire the common physical angles.
+We select $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]$ for $m\in\{1,M/2\}$.
+Given this estimation, we can further estimate the common directions, i.e., the DoDs at the BS and the coupled directions at the RIS.
+III-C1 Common DoDs Estimation at BS
+Given the estimated cascaded channel $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m},\bar{m}\in\left\{1,M/2\right\}$, it is noticed that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\hat{\mathbf{x}}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$.
+This means that given the accurate estimation of $\hat{\mathbf{x}}$, we only need to determine its non-zero positions corresponding to the columns in the matrix $\mathbf{\Pi}_{\bar{m}}$.
+Since $\mathbf{\Pi}_{\bar{m}}\triangleq\left(\mathbf{C}_{\mathrm{T},\bar{m}}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},\bar{m}}\right)$ is in one-to-one correspondence with the ARVs, this relationship allows us to directly obtain the DoDs at BS.
+However, inevitable noise and estimation errors can cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, which will lead to more than $\Sigma_{k}$ non-zero entries.
+To address this, we propose a heuristic method to determine the accurate non-zero positions, referred to as common physical support, and calculate the corresponding physical DoDs at the BS, as summarized in Algorithm 1.
+Algorithm 1 DoDs Estimation at BS
+0:  Angle-excluded matrix $\mathbf{x}_{k,\bar{m}},\bar{m}\in\left\{1,M/2\right\}$, the path number $L$ of BS-RIS channel and the quantization level of overcomplete dictionaries $Q_{\mathrm{R}}$ and $Q_{\mathrm{T}}$.
 1:  $\boldsymbol{\epsilon}=\left[\epsilon_{1},\cdots,\epsilon_{Q_{\mathrm{T}}}\right]$ and $\bar{\boldsymbol{\varphi}}=-1+\frac{1+2\boldsymbol{q}}{Q_{\mathrm{T}}},\boldsymbol{q}=1,\dots,Q_{\mathrm{T}}$.
@@
 5:  end for
-6:  $\zeta^{\mathrm{BS}}=\left\{q|\mathop{\mathrm{maxk}}\limits_{q}\left(\boldsymbol{\epsilon},L\right)\right\}$
-7:  $\hat{\boldsymbol{\varphi}}=\bar{\boldsymbol{\varphi}}\left(\zeta^{\mathrm{BS}}\right)$.
-7:  $\zeta^{\mathrm{BS}}$ and $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$.
-In the proposed EnM method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, in which each entry is the normalized results of the corresponding element in $\mathbf{x}_{k,\bar{m}}$ (Step $2$).
-As mentioned in Section III-A, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, with each block contains $Q_{\mathrm{B}}$ columns.
+6:  $\mathcal{I}^{\mathrm{BS}}=\left\{q|\mathop{\mathrm{maxk}}\limits_{q}\left(\boldsymbol{\epsilon},L\right)\right\}$
+7:  $\hat{\boldsymbol{\varphi}}=\bar{\boldsymbol{\varphi}}\left(\mathcal{I}^{\mathrm{BS}}\right)$.
+7:  $\mathcal{I}^{\mathrm{BS}}$ and $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$.
+In the proposed method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, where each entry is the normalized value of the corresponding element in $\hat{\mathbf{x}}_{k,\bar{m}}$ (Step $2$).
+As observed in Section III-B3, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, each containing $Q_{\mathrm{B}}$ columns.
 Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way.
-By accumulating the $\ell_{2}$-norm for each block of the two normalized energy basis vectors, the merged-energy vector $\boldsymbol{\epsilon}$ can establish a one-to-one correspondence with the physical DoDs at BS in the entire beam space (Step $3-5$).
-Subsequently, the indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ can constitute the common support $\zeta^{\mathrm{BS}}$ (Step $6$).
-Finally, with knowledge of $\zeta^{\mathrm{BS}}$, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ is easily obtained according to the sampling rule of the CBS dictionary (Step $7$).
+By accumulating the $\ell_{2}$-norm for each block, we form the merged-energy vector $\boldsymbol{\epsilon}$ to establish a one-to-one correspondence with the physical DoDs at BS (Step $3-5$).
+The indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ will form the common physical support $\mathcal{I}^{\mathrm{BS}}$ (Step $6$).
+Finally, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ can be obtained using the knowledge of $\mathcal{I}^{\mathrm{BS}}$ and the sampling rule of the CBS dictionary (Step $7$).
 Fig.
-3 depicts an example of the estimated energy spectrum employing the proposed EnM method.
-Assuming that the cascaded channel contains $L=3$ BS-RIS paths with the common support $\zeta^{\mathrm{BS}}=\left\{3,4,24\right\}$ and SNR $=20$ dB.
-It can be observed from the spectrum that the energy is distributed surrounding the estimated DoDs.
-Furthermore, the energy is concentrated around the actual direction and can be easily distinguished, which validates the effectiveness of the proposed EnM method.
+4 depicts an example of the estimated energy spectrum employing the proposed method.
+The cascaded channel is assumed to contain $L=3$ BS-RIS paths with the common physical support $\mathcal{I}^{\mathrm{BS}}=\left\{3,4,24\right\}$ and SNR $=20$ dB.
+The spectrum reveals that the energy is distributed surrounding the estimated DoDs which closely matches the actual directions and can be easily distinguished.
+This validates the effectiveness of the proposed method.
 Figure 4: Normalized energy spectrum via EnM algorithm.
-IV-B DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
-The proposed EnM method ensures the acquisition of the DoDs at BS.
-However, due to the coupling between the DoAs and DoDs at RIS, the aforementioned method can be challenging to obtain angular information at RIS.
-Therefore, in this subsection we propose a novel DS-MUSIC algorithm to acquire the coupled DoAs/DoDs at RIS.
-Firstly, with knowledge of the estimated DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ from Section IV-A, the estimated cascaded channel $\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}]$ can be projected onto the subspace of ARM of BS using the pseudo-inverse, which is given by
-$\begin{split}\mathbf{U}_{k,\bar{m}}&\triangleq\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}](\hat{\mathbf{A}}_{N_{\mathrm{T}},\bar{m}}^{\mathsf{H}}\left(\boldsymbol{\varphi}\right))^{\dagger}\\ &=\mathrm{Diag}\left(\mathbf{h}_{k}[\bar{m}]\right)\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\psi}\right)\mathbf{\Upsilon}_{\bar{m}},\end{split}$ (61)
-where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$ and $\left(\mathbf{A}_{N}^{\mathsf{H}}\right)^{\dagger}=(\mathbf{A}_{N}^{\dagger})^{\mathsf{H}}$.
-Besides, the rest-CSI matrix $\mathbf{U}_{k,\bar{m}}\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ can be block-partitioned column-wise as $\mathbf{U}_{k,\bar{m}}=\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]$.
+III-C2 Common Coupled Angle Estimation at RIS
+With the estimated DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$, the estimated cascaded channel $\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}]$ can be projected onto the subspace of ARM of BS using the pseudo-inverse, which can be given by
+$\begin{split}\mathbf{U}_{k,\bar{m}}&\triangleq\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}](\hat{\mathbf{A}}_{N_{\mathrm{T}},\bar{m}}^{\mathsf{H}}\left(\boldsymbol{\varphi}\right))^{\dagger}\\ &=\mathrm{Diag}\left(\mathbf{h}_{k}[\bar{m}]\right)\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\psi}\right)\mathbf{\Sigma}_{\bar{m}},\end{split}$ (26)
+where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$, and $\mathbf{U}_{k,\bar{m}}\triangleq\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ denotes the rest-CSI matrix, which is only related to the AoAs and AoDs at RIS.
 By substituting (7) into (26), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
-$\begin{split}\mathbf{u}_{k,\bar{m},l}&=\!\upsilon_{\bar{m},l}\left(\mathbf{A}_{\mathrm{R},\bar{m}}^{*}\left(\boldsymbol{\vartheta}_{k}\right)\boldsymbol{\beta}_{k,\bar{m}}\right)\bullet\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\right)\\ &=\!\upsilon_{\bar{m},l}\!\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\!-\!\vartheta_{k,1}\!\right)\!,\!\cdots\!,\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\!-\!\vartheta_{k,J_{k}}\!\right)\right]\!\boldsymbol{\beta}_{k,\bar{m}},\end{split}$ (62)
-where $\upsilon_{l}$ is the $\left(l,l\right)$-th entry in $\mathbf{\Upsilon}_{m}$.
-Define $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM where $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right],l=1,\dots,L,j=1,\dots,J_{k}$ and $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\upsilon_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ as the amplitude vector, (62) is rewritten as
-$\mathbf{u}_{k,\bar{m},l}=\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\boldsymbol{\chi}_{k,\bar{m},l}.$ (63)
-To find the coupled physical directions $\left\{{\hat{\gamma}}_{k,l,j}\right\}_{j=1}^{J_{k}}$ in (63), this issue can be regarded as a frequency estimation finding problem.
-Inspired by this, we propose a novel DS-MUSIC algorithm as described in Algorithm 3, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
+$\begin{split}\mathbf{u}_{k,\bar{m},l}&=\Sigma_{\bar{m},l}\left(\mathbf{A}_{\mathrm{R},\bar{m}}^{*}\left(\boldsymbol{\vartheta}_{k}\right)\boldsymbol{\beta}_{k,\bar{m}}\right)\bullet\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\right)\\ &=\Sigma_{\bar{m},l}\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,1}\!\right),\dots,\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,J_{k}}\!\right)\right]\!\boldsymbol{\beta}_{k,\bar{m}},\end{split}$ (27)
+where $\Sigma_{\bar{m},l}$ denotes the $\left(l,l\right)$-th element in $\Sigma_{\bar{m}}$, and $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right]$ represents the coupled angle for $l=1,\dots,L$ and $j=1,\dots,J_{k}$.
+By defining $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM, (27) can be rewritten as
+$\mathbf{u}_{k,\bar{m},l}=\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l,j}\right)\boldsymbol{\chi}_{k,\bar{m},l},$ (28)
+where $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with the path gains and phase shifts for each coupled angle.
+The expression in (28) represents a typical DoA estimation problem, which can be effectively addressed using the MUSIC algorithm due to its high resolution and robustness against noise [29].
+However, there are two key challenges in directly applying the MUSIC algorithm.
+Firstly, the effective angle range for $\boldsymbol{\gamma}_{k,l,j}$, which is the difference between two angles $\psi_{l}$ and $\theta_{k,j}$, is doubled compared to the classical angle range.
+Moreover, the relative frequency parameter $\eta_{m}$ introduced by beam split effect further expands this range.
+This results in an increased number of potential estimated angles, making it difficult to distinguish the true physical directions.
+Secondly, the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC can cause energy leakage into the noise space, which further complicates the identification of physical directions.
+To address these challenges, we combine the information from two selected subcarriers and apply spatial smoothing preprocessing to the projected rest-CSI matrix $\mathbf{U}_{k,\bar{m}}$.
+Inspired by the smooth-MUSIC algorithm, we propose a novel DS-MUSIC algorithm to accurately estimate the coupled angles, as described in Algorithm 2, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
+We first preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray containing $N_{\mathrm{sub}}$ reflecting elements.
+There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total, which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}}$, $N_{\mathrm{sub}}>J_{k}$.
+The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
+$\mathbf{u}_{k,\bar{m},l,n}=\mathbf{A}_{\mathrm{R},\bar{m},n}\left(\boldsymbol{\gamma}_{k,l}\right)\boldsymbol{\chi}_{k,\bar{m},l},$ (29)
+where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(:,n:n+N_{\mathrm{sub}}-1\right)$.
+Therefore, the covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ can be written as
+$\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}=\mathbf{u}_{k,\bar{m},l,n}\mathbf{u}_{k,\bar{m},l,n}^{\mathsf{H}}.$ (30)
 Algorithm 2 DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
@@
 10:        end for
-11:        $\mathbf{p}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\!=\!\log\!\left[\!\min\!\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\right)\!\!,\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\!\!,\min\!\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\right)\!\right]$.
+11:        $\mathbf{p}_{k,\bar{m},l}=\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right),\tilde{\mathbf{p}}_{k,\bar{m},l},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right)\right]$.
 […]
@@
 15:  end for
-16:  $\zeta^{\mathrm{RIS}}=\mathcal{Q}_{1}\cup\cdots\cup\mathcal{Q}_{L}$.
-17:  $\hat{\boldsymbol{\gamma}}_{k}=\bar{\boldsymbol{\gamma}}\left(\zeta^{\mathrm{RIS}}\right)$.
-17:  $\zeta^{\mathrm{RIS}}$ and $\hat{\boldsymbol{\gamma}}_{k}$.
-Due to the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC, some energy from physical directions will leak into the noise space, making it challenging to distinguish physical directions.
-To address this issue, we preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray contains $N_{\mathrm{sub}}$ reflecting elements.
-There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}},N_{\mathrm{sub}}>J_{k}$.
-The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
-$\mathbf{u}_{k,\bar{m},l,n}=\mathbf{A}_{\mathrm{R},\bar{m},n}\left(\boldsymbol{\gamma}_{k,l}\right)\boldsymbol{\chi}_{k,\bar{m},l},$ (64)
-where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(n:n+N_{\mathrm{sub}}-1\right)$.
-The covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ is given by
-$\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}=\mathbf{u}_{k,\bar{m},l,n}\mathbf{u}_{k,\bar{m},l,n}^{\mathsf{H}}.$ (65)
-The covariance matrix $\mathbf{R}_{k,\bar{m},l}^{u}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ of $\mathbf{u}_{k,\bar{m},l}$ is calculated by averaging all $N_{\mathrm{a}}$ $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}$, which is given by
+16:  $\mathcal{I}^{\mathrm{RIS}}=\mathcal{Q}_{1}\cup\cdots\cup\mathcal{Q}_{L}$.
+17:  $\hat{\boldsymbol{\gamma}}_{k}=\bar{\boldsymbol{\gamma}}\left(\mathcal{I}^{\mathrm{RIS}}\right)$.
+17:  $\mathcal{I}^{\mathrm{RIS}}$ and $\hat{\boldsymbol{\gamma}}_{k}$.
+To implement the spatial smoothing, the covariance matrix $\mathbf{R}_{k,\bar{m},l}^{u}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ of $\mathbf{u}_{k,\bar{m},l}$ is calculated by averaging all $N_{\mathrm{a}}$ subarray covariance matrices $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}$, which can be expressed by
 $\mathbf{R}_{k,\bar{m},l}^{u}=\frac{1}{N_{\mathrm{a}}}\sum_{n=1}^{N_{\mathrm{a}}}{\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}}=\frac{1}{N_{\mathrm{a}}}\tilde{\mathbf{U}}_{k,\bar{m},l}\tilde{\mathbf{U}}_{k,\bar{m},l}^{\mathsf{H}},$ (31)
-where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ in (64).
-The eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ yields
+where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace-supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ as defined in (29).
+Then, the eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ can be expressed as
 $\mathbf{R}_{k,\bar{m},l}^{u}=\tilde{\mathbf{\Omega}}_{k,\bar{m},l}\mathrm{Diag}(\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l})\tilde{\mathbf{\Omega}}_{k,\bar{m},l}^{\mathsf{H}},$ (32)
-where $\tilde{\boldsymbol{\Omega}}_{k,\bar{m},l}\triangleq\left[\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,1},\cdots,\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,N_{\mathrm{sub}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ is the eigenbasis matrix which contains $N_{\mathrm{sub}}$ eigenvectors and $\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l}\in\mathbb{C}^{N_{\mathrm{sub}}}$ is the eigenvalue vector.
-It is obvious that $\mathbf{R}_{k,\bar{m},l}^{u}$ is a Hermitian matrix of all its $N_{\mathrm{sub}}$ eigenvectors are orthogonal to each other.
-Arrange the eigenvalues of $\mathbf{R}_{k,\bar{m},l}^{u}$ in descending order and organize the corresponding eigenvectors accordingly, it can be rewritten as
+where $\tilde{\boldsymbol{\Omega}}_{k,\bar{m},l}\triangleq\left[\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,1},\cdots,\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,N_{\mathrm{sub}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ denotes the eigenbasis matrix which contains $N_{\mathrm{sub}}$ eigenvectors and $\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l}\in\mathbb{C}^{N_{\mathrm{sub}}}$ is the eigenvalue vector.
+Note that $\mathbf{R}_{k,\bar{m},l}^{u}$ is a Hermitian matrix, which means that all its $N_{\mathrm{sub}}$ eigenvectors are orthogonal to each other.
+By arranging the eigenvalues of $\mathbf{R}_{k,\bar{m},l}^{u}$ in descending order and organizing the corresponding eigenvectors accordingly, it can be rewritten as
 $\mathbf{R}_{k,\bar{m},l}^{u}=\mathbf{\Omega}_{k,\bar{m},l}\mathrm{Diag}\left(\boldsymbol{\lambda}_{k,\bar{m},l}\right)\mathbf{\Omega}_{k,\bar{m},l}^{\mathsf{H}},$ (33)
-where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ are the sorted eigenbasis matrix and eigenvalue vector, respectively.
-Subsequently, $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts with $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$.
-Wherein, $J_{k}$ largest eigenvalues corresponding to the directions of largest variability span the signal subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$, while the remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors span the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which is orthogonal to the signal subspace.
-Define the sub-ARV $\mathbf{a}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}(\gamma)\in\mathbb{C}^{N_{\mathrm{sub}}}$ as the first $N_{\mathrm{sub}}$ row of $\mathbf{a}_{\mathrm{R},\bar{m}}(\gamma)$.
-Similar to the definition of sub-ARV, we define $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the sub-CBS dictionary which consists of the first $N_{\mathrm{sub}}$ rows.
-By employing the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}$, one can scan the entire beam space and calculate the MUSIC-pseudo-spectrum (PS) for the beams aligned with any sampled direction.
-Define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}\triangleq[\tilde{p}_{k,\bar{m},l,1}^{\mathrm{MUSIC}},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}^{\mathrm{MUSIC}}]^{\mathsf{T}}$ as the MUSIC-PS, wherein $\tilde{p}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}$ is the spectral value, which follows a $\ell_{2}$-norm form as
-$\tilde{p}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}},$ (69)
-where $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\!\triangleq\!e^{\!-j2\pi\eta_{m}\left(\!\frac{q}{Q_{\mathrm{R}}}\!-1\right)\mathbf{n}_{\mathrm{sub}}}$ with $\mathbf{n}_{\mathrm{sub}}\!=\!\left[1,\cdots,N_{\mathrm{sub}}\right]$.
-The MUSIC-PS is obtained by mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$.
-Based on the fact that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ that resides in the signal subspace is orthogonal to the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the $\ell_{2}$-norm in (69) associated with the actual directions should approximate zero.
-The MUSIC algorithm is typically dependent on peak detection to find actual physical directions, resulting in undetectable marginal peaks.
-To address this issue, we design a novel extended-log-MUSIC-PS (EL-MUSIC-PS), where we extend the spectrum and map it to the logarithmic domain to enhance the accuracy of estimation.
-The EL-MUSIC-PS can be express as
-$\mathbf{p}_{k,\bar{m},l}^{\mathrm{MUSIC}}=\log\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\right),\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\right)\right].$ (70)
-In this form, the actual directions can be determined by employing local-maximum algorithms. Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
-$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{q}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{q}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (71)
-where $K_{q}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$.
-It is observed that when $q^{\prime}=q+K_{q}Q_{\mathrm{R}}/\eta_{m}$, $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ will generate the identical beam.
-Moreover, the maximum number of identical beams $K_{q}^{\max}$ can be obtained by solving the integer equation $K_{q}^{\max}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$, which can be derived as
-$K_{q}^{\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (72)
-Therefore, more than $J_{k}$ peaks would be detected in the EL-MUSIC-PS in (70), making it challenging to distinguish the true and fake physical directions in one single SC.
+where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ denote the sorted eigenbasis matrix and eigenvalue vector, respectively.
+Therefore, the matrix $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts: $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$.
+The part $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ consists of the $J_{k}$ largest eigenvalues corresponding to the directions of largest variability, which spans the signal subspace.
+The remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors form $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, spans the noise subspace and is orthogonal to the signal subspace.
+To calculate the MUSIC pseudo-spectrum, the practical implementation requires sampling the estimated angles.
+We define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}\triangleq[\tilde{p}_{k,\bar{m},l,1},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}]^{\mathsf{T}}$ as the MUSIC pseudo-spectrum, where $\tilde{p}_{k,\bar{m},l,q}$ is the sampling spectral value for $q=1,\dots,Q_{\mathrm{B}}$.
+To align these sampled values with our proposed algorithm, we adjust the angle sampling within the range of the CBS dictionary.
+Therefore, we define the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the first $N_{\mathrm{sub}}$ rows of the original CBS dictionary in (15).According to (17), each element in the sub-CBS dictionary $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ can be derived as
+$\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\triangleq e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}},q=1,\dots,Q_{\mathrm{B}},$ (34)
+where $\mathbf{n}_{\mathrm{sub}}=\left[1,\cdots,N_{\mathrm{sub}}\right]$.
+By mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the MUSIC algorithm constructs the pseudo-spectrum by evaluating the inverse of the projection’s norm at each potential direction, which can be given as
+$\tilde{p}_{k,\bar{m},l,q}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}}.$ (35)
+Given that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ in the signal subspace should ideally have minimal projection onto the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which means that the $\ell_{2}$-norm in (35) associated with the actual directions should approximate zero.
+The MUSIC algorithm typically relies on peak detection to find actual physical directions.
+To avoid undetectable marginal peaks, we extend the pseudo-spectrum by inserting the minimum value at its edges.
+The extended pseudo-spectrum can be express as
+$\mathbf{p}_{k,\bar{m},l}=\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right),\tilde{\mathbf{p}}_{k,\bar{m},l},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right)\right].$ (36)
+The potential directions can then be determined by employing local-maximum algorithms.
+However, due to the extended range of coupled angles, the estimated number of directions will be $K_{\text{cp}}$ times the original.
+Proposition 2.
+The maximum number of identical beams can be given as
+$K_{\rm{cp},\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (37)
+Proof.
+Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
+$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{\text{cp}}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{\text{cp}}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (38)
+where $K_{\text{cp}}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$.
+It is observed that when $q^{\prime}=q+K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, the element in sub-CBS dictionary will generate the identical beam, i.e., $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}=\xi_{\mathrm{R},m,q^{\prime}}^{\mathrm{sub}}$.
+Therefore, the maximum number of identical beams $K_{\text{cp},\max}$ can be obtained by solving the integer equation $K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$.
+By rearranging the terms and employing the relationship $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$, we can obtain
+$K_{\text{cp}}=\frac{\left(2Q_{\mathrm{R}}-1\right)\eta_{m}}{Q_{\mathrm{R}}}.$ (39)
+Therefore, the maximum value of $K_{\text{cp}}$ is the integer part of this expression, which can be given as
+$K_{\rm{cp},\max}=\left\lfloor\eta_{m}\left(2-\frac{1}{Q_{\mathrm{R}}}\right)\right\rfloor+1.$ (40)
+According to Proposition 2, instead of the expected $J_{k}$ peaks in the MUSIC pseudo-spectrum, $K_{\text{cp},\max}J_{k}$ peaks will be detected.
+This makes the algorithm infeasible for distinguishing the true physical directions from fake ones within a single SC.
 Thankfully, it is observed from (37) that the spatial directions associated with true physical directions keep frequency-independent.
 Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction.
-In this approach, we select two SCs, calculate and compare their EL-MUSIC-PSs on the same plot.
+In this approach, we select two SCs to calculate and compare their extended MUSIC pseudo-spectrum.
 The positions where peaks overlap indicate the true physical directions.
-However, due to the unavoidable noise and estimation errors in full CSI, the EL-MUSIC-PSs of different SCs may exist offsets.
-Therefore, in practical applications, we select the $J_{k}$ closest peaks to represent true physical directions, and we introduce the peak distance $d$ to measure the distance between two peaks, which is given by
+However, due to the unavoidable noise and estimation errors, the pseudo-spectrum of different SCs may exist offsets.
+Therefore, in practical implement, we select the $J_{k}$ closest peaks to represent true physical directions.
+We introduce the peak distance $d$ to measure the distance between two peaks, which is given by
 $d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=\left|q_{m}-q_{m^{\prime}}^{\prime}\right|,$ (41)
 where $q_{m}$ and $q_{m^{\prime}}^{\prime}$ denote the indices of the peaks at two selected SCs, respectively.
-It can be observed from (72) that the selection of SC $m$ affects the number of identical beams.
-To avoid additional algorithm complexity, the selection of SCs should satisfy the constraint $K_{q,\max}\leq 2$.
-Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into it, we can obtain
+So far, the detailed steps of the proposed DS-MUSIC algorithm are summarized in Algorithm 2.
+To ensure the algorithm is both effective and efficient, it is crucial to select the appropriate SCs.
+Since $K_{\text{cp},\max}$ in (37) is affected by the choice of SCs, we should aim to satisfy $K_{\text{cp},\max}\leq 2$, which means each angle has at most two identical peaks.
+Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into (37), we can obtain
 $m\leq\frac{M}{2}+\frac{Mf_{c}Q_{\mathrm{R}}}{B\left(Q_{\mathrm{B}}\right)}+\frac{1}{2}.$ (42)
-Substituting (5) into (73), the peak distance $d$ can be derived as
-$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{q}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (75)
-To distinguish true physical directions effectively, the peak distance should be sufficiently large.
+Besides, substituting (5) into (41), the peak distance $d$ can be derived as
+$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{\text{cp}}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (43)
+To distinguish true physical directions effectively, the peak distance should also be sufficiently large.
 According to (43), it is evident that the intervals between selected SCs should be large enough.
@@
 Fig.
-4 depicts an example of estimated EL-MUSIC-PS with the proposed DS approach.
-Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for a arbitrary UE with SNR $=20$ dB.
-It can be observed that the peaks overlap in actual physical directions, while that associated with fake spatial directions are totally separated.
+5 depicts an example of estimated extended MUSIC pseudo-spectrum with the proposed DS approach.
+Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for an arbitrary UE with SNR $=20$ dB.
+It can be observed that the peaks overlap at actual physical directions, while those associated with fake spatial directions are completely separated.
 The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
-Figure 4: Normalized EL-MUSIC-PS via DS-MUSIC algorithm.
-IV-C LS Estimation for Angle-Excluded Coefficients
-With the estimated DoDs at BS and coupled DoAs/DoDs at RIS, (10) can be rewritten as
+Figure 5: Normalized extended MUSIC pseudo-spectrum via DS-MUSIC algorithm.
+III-D Channel Estimation for Other Subcarriers
+Through the proposed two-step full CSI estimation, we obtained the cascaded channels for two subcarriers and the corresponding common physical angle.
+Based on this knowledge, the cascaded channel in (10) can be rewritten as
 $\mathbf{h}_{k}^{\mathrm{cas}}[m]=\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\otimes\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\tilde{\mathbf{x}}_{k,m},$ (44)
-where $\tilde{\mathbf{x}}_{k,m}\!\triangleq\!\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Upsilon}_{m}\!\in\mathbb{C}^{L^{2}\!J_{k}}$ denotes the angle-excluded vector, $\hat{\boldsymbol{\varphi}}\triangleq[\hat{\varphi}_{1},\dots,\hat{\varphi}_{L}]\in\mathbb{C}^{L}$ denotes the estimated DoDs at BS of all $L$ paths, and $\hat{\boldsymbol{\gamma}}_{k}\triangleq[\hat{\boldsymbol{\gamma}}_{k,1}^{\mathsf{T}},\dots,\hat{\boldsymbol{\gamma}}_{k,L}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{LJ_{k}}$ denotes the estimated coupled directions at RIS.
+where $\tilde{\mathbf{x}}_{k,m}\!\triangleq\!\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m}\!\in\mathbb{C}^{L^{2}\!J_{k}}$ represents the angle-excluded vector, $\hat{\boldsymbol{\varphi}}\triangleq[\hat{\varphi}_{1},\dots,\hat{\varphi}_{L}]\in\mathbb{C}^{L}$ is the estimated DoDs at BS of all $L$ paths, and $\hat{\boldsymbol{\gamma}}_{k}\triangleq[\hat{\boldsymbol{\gamma}}_{k,1}^{\mathsf{T}},\dots,\hat{\boldsymbol{\gamma}}_{k,L}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{LJ_{k}}$ denotes the estimated coupled directions at RIS.
 Substituting (44) into (20), the received signal $\mathbf{y}_{k,t}\left[m\right]$ can be rewritten as
 $\mathbf{y}_{k,t}\left[m\right]=\tilde{\mathbf{\Psi}}_{m}\tilde{\mathbf{x}}_{k,m},$ (45)
-where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\overline{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$.
-In this case, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained employing LS solution
+where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\bar{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$.
+The column size of the $\tilde{\mathbf{\Psi}}_{m}$ matrix is reduce from $Q_{\mathrm{B}}Q_{\mathrm{T}}$ of overcomplete dictionary to a multipath-related size $L^{2}J_{k}$.
+This significant reduction in the column size enhances both the efficiency and feasibility of the LS solution.
+Therefore, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained as
 $\hat{\tilde{\mathbf{x}}}_{k,m}=\left(\tilde{\mathbf{\Psi}}_{m}^{\mathsf{H}}\tilde{\mathbf{\Psi}}_{m}\right)^{-\mathsf{1}}\tilde{\mathbf{\Psi}}_{m}^{\mathsf{H}}\mathbf{y}_{k,t}\left[m\right].$ (46)
-Then, the full CSI can be calculated by
+Substituting (46) into (44), the full CSI for other SCs $m\in\mathcal{M}\backslash\{1,M/2\}$ can be expressed as
 $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\otimes\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\hat{\tilde{\mathbf{x}}}_{k,m}.$ (47)
-IV-D Complexity Analysis
-In this subsection, we analyze the computational complexity of the proposed channel estimation scheme.
-Specifically, the complexity of the proposed scheme mainly stems from four major steps: 1) Full CSI estimation via CBS-GAMP approach $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$, 2) DoDs acquisition at BS via EnM method $\mathcal{O}\left(Q_{\mathrm{T}}Q_{\mathrm{B}}\right)$, 3) Coupled directions acquisition at RIS via DS-MUSIC algorithm $\mathcal{O}\left(2N_{r}^{3}\right)$, 4) Angle-excluded coefficients estimation via LS solution $\mathcal{O}\left(\left(M-2\right)L^{6}J_{k}^{3}\right)$.
-The overall complexity of the proposed scheme is given by
-$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right)$ (80)
-For comparison, we also discuss the complexity of the conventional OMP scheme [33], BSA-OMP scheme [14] and CBS-GAMP scheme.
-Since the above algorithms need to be executed at each SC, the total computational complexity thus is multiplied by $M$ compared to the original.
-The conventional OMP algorithm has the same complexity $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$ as BSA-OMP approach.
+IV Analysis of Computational Complexity
+In this section, we analyze the computational complexity of the proposed channel estimation protocol.
+We also compare our results with the other existing algorithms.
+Specifically, the complexity of the proposed scheme mainly stems from four major components:
+  1. 1.
+    Cascaded channel estimation for two selected SCs by CBS-based GAMP approach: $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
+  2. 2.
+    DoDs acquisition at BS: $\mathcal{O}\left(Q_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
+  3. 3.
+    Coupled directions acquisition at the RIS using DS-MUSIC algorithm: $\mathcal{O}\left(2N_{r}^{3}\right)$.
+  4. 4.
+    Channel estimation for other SCs: $\mathcal{O}\left(\left(M-2\right)L^{6}J_{k}^{3}\right)$.
+The overall complexity of the proposed scheme can be expressed as
+$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right).$ (48)
+For comparison, we also discuss the complexity of the conventional OMP scheme [30], BSA-OMP scheme [17] and CBS-GAMP scheme [23].
+These algorithms perform under traditional channel estimation frameworks which means they need to be executed at each SC or simultaneously process all SCs.
+The computational complexity of conventional OMP algorithm is $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$.
+The BSA-OMP approach has the same complexity as the conventional OMP, $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$, since it similarly processes each subcarrier independently.
 The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
-Due to the sparse characteristics of THz channel, the number of propagation path satisfies $L,J_{k}\ll Q_{\mathrm{T}},Q_{\mathrm{R}}$.
-The number of SCs $M$ is also quite large in wideband systems.
-Therefore, it is clear that the computational complexity of the proposed scheme is much lower than other three approaches.
+Additionally, the number of SCs $M$ is typically large in wideband systems.
+Therefore, it is evident that the computational complexity of the proposed scheme is significantly lower than the other approaches mentioned.
 V Numerical Results and Discussion
 In this section, we present numerical results to validate the effectiveness of the proposed scheme.
-Without loss of generality, we consider a wideband RIS-assisted THz scenario with $f_{c}=100$ GHz, $B=15$ GHz and $M=128$.
-A BS employing a $N_{\mathrm{T}}$ = 16 ULA transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements.
-For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths and the AoAs and AoDs are uniformly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$.
-The quantization level of CBS dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$.
-As a metric for CSI estimation performance, we use the normalized mean square error (NMSE), which is given by
-$\mathrm{NMSE}=\mathbb{E}\left\{\frac{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]-\mathbf{\hat{h}}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}\right\}$ (81)
-Besides, we employ the correct probability as a metric for direction estimation performance, which is given by
-$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\}$ (82)
+Without loss of generality, we consider a wideband RIS-assisted THz scenario with central carrier frequency $f_{c}=100$ GHz, bandwidth $B=15$ GHz and $M=128$ SCs.
+A BS employing a ULA with $N_{\mathrm{T}}$ = 16 antennas transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements.
+For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths, and the AoAs and AoDs are randomly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$.
+The quantization level of overcomplete dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$.
+We use the normalized mean square error (NMSE) as a metric for CSI estimation performance, which is given by
+$\mathrm{NMSE}=\mathbb{E}\left\{\frac{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]-\mathbf{\hat{h}}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}\right\}.$ (49)
+Additionally, we employ the correct probability as a metric for direction estimation performance, which is given by
+$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\},$ (50)
 where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ is the threshold which is set as half the sample interval.
-For comparison, we also consider the conventional OMP scheme [33], BSA-OMP scheme [14], CBS-GAMP scheme, and oracle-LS scheme as benchmark schemes in the simulation.
-It is notice that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
+For comparison, we also consider the conventional OMP scheme [30], BSA-OMP scheme [17], CBS-GAMP scheme [23], and oracle-LS scheme [31] as benchmark schemes in the simulation.
+It is noted that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
 Figure 6: Correct probability of DoDs at BS and coupled DoAs/DoDs at RIS against SNR.
@@
 6 depicts the correct probability performance of the DoDs at BS and coupled DoAs/DoDs at RIS versus SNR with $T=50$ and $P=30$.
-It is shown that both the two correct probabilities approach one as SNR increases.
-This indicates that the proposed scheme can ensure a certain level of reliability even with low SNR conditions and has the ability to obtain accurate physical directions with sufficiently well channel conditions.
-Besides, the correct probability of the coupled directions at RIS is lower than that of the DoDs at BS in all SNR regions.
-This is because the estimation of the former relies on the latter.
+It is shown that both the correct probabilities approach one as SNR increases.
+This indicates that the proposed scheme can ensure a certain level of reliability even under low SNR conditions and has the ability to obtain accurate physical directions with sufficient channel conditions.
+Additionally, the correct probability of the coupled directions at RIS is consistently lower than that of the DoDs at BS across all SNR values.
+This is because the estimation of the coupled directions at RIS depends on the accuracy of the DoDs estimation at BS.
 Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves.
@@
 7 illustrates the NMSE performance versus SNR with $T=50$ and $P=30$.
-Overall, we can observe that the proposed scheme outperforms conventional-OMP and BSA-OMP algorithm in all SNR regions.
+Overall, it can be observed that the proposed scheme outperforms conventional OMP and BSA-OMP algorithm across all SNR regions.
 This is because the proposed scheme has the ability to exploit unknown channel information, and the proposed dictionaries can effectively cover the beam space under beam split effect.
-Besides, it can be observed that the performance of the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB.
+Additionally, the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB.
 This is because estimation errors generated by partial SCs propagate more significantly to the remaining SCs under poor channel conditions.
@@
 This is because angle decoupling provides more accurate angular estimation information.
-This implies that there exists a performance-complexity trade-off between these two algorithms.
-Moreover, it is obvious that the proposed scheme can approach NMSE performance of the oracle-LS scheme, verifying the effectiveness of the proposed scheme.
+This suggests that there exists a performance-complexity trade-off between these two algorithms.
+Moreover, it is obvious that the proposed scheme can approach the NMSE performance of oracle-LS scheme, which verifies the effectiveness of the proposed scheme.
 Figure 8: NMSE performance comparison against the number of subframes $T$.
@@
 8 shows the NMSE performance versus the number of subframes $T$, with SNR $=20$ dB and $P=15$.
-As expected, with the increase in the number of subframes $T$, NMSE performance of all the algorithms improves.
-We can observe that the proposed scheme can maintain well performance even in the case with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm.
-Moreover, It can be observed that compared to the CBS-GAMP method, the proposed algorithm can reach similar performance with approximately $50\%$ subframes reduction.
+As expected, the NMSE performance of all the algorithms improves with the increase in the number of subframes $T$.
+It can be observed that the proposed scheme maintains good performance even with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm.
+Moreover, compared to the CBS-GAMP method, the proposed algorithm achieves similar performance with approximately a $50\%$ subframes reduction.
 This indicates that the proposed scheme can efficiently reduce the pilot overhead for channel estimation.
@@
 9 provides the NMSE performance versus the number of time slots $P$, with SNR $=20$ dB and $T=30$.
-As expected, with the increase in the number of slots $P$, NMSE performance of all the algorithms improves.
-We can observe that the proposed scheme outperforms all other three algorithms and can achieve near-optimal NMSE performance of the ideal oracle-LS upper bound in all considered $P$.
-The performance gap between the proposed scheme and other existing schemes is quite great, demonstrating that the proposed scheme can effectively reduce the system complexity for channel estimation.
+As expected, the NMSE performance of all the algorithms improves with the increase in the number of slots $P$.
+It can be observed that the proposed scheme outperforms all other three algorithms and can achieve near-optimal NMSE performance of the ideal oracle-LS upper bound in all considered values of $P$.
+The performance gap between the proposed scheme and other existing schemes is quite large.
+This indicates that the proposed scheme can effectively reduce the system complexity for channel estimation.
 Figure 10: NMSE performance comparison against the number of paths $L,J_{k}$.
@@
 10 illustrates the NMSE performance versus the number of paths $L$ and $J_{k}$, with $T=50$ and $P=30$.
-We can observe that as the the number of paths increases, the NMSE performance of both the two schemes degrade.
-This is because of the increased estimation parameters and challenges in distinguishing directions between more multipaths.
-Furthermore, as SNR increases, the NMSE of CBS-GAMP algorithm converges to around $5\times 10^{-4}$ regardless of the number of paths, while the NMSE of the proposed scheme remain decreasing in all SNR regions.
-This indicates that the proposed scheme exhibits a superior performance.
+It can be observed that as the the number of paths increases, the NMSE performance of both the two schemes degrade.
+This is because of the increased estimation parameters and challenges in distinguishing directions among more multipath components.
+Furthermore, as SNR increases, the NMSE of CBS-GAMP algorithm converges to around $5\times 10^{-4}$ regardless of the number of paths.
+In contrast, the NMSE of the proposed scheme remain decreasing across all SNR regions.
+This demonstrates that the proposed scheme maintains better estimation accuracy and is more robust against variations in the number of paths compared to the CBS-GAMP algorithm.
 Figure 11: NMSE performance comparison against the number of reflecting elements of RIS $N_{\mathrm{R}}$.
@@
 11 shows the NMSE performance versus the number of reflecting elements of RIS $N_{\mathrm{R}}$, with $T=50$ and $P=30$.
-We can observe that as $N_{\mathrm{R}}$ decreases, the NMSE performance of the BSA-OMP algorithm degrade greatly.
-This is because the small-scale RIS cannot provide enough measurement data diversity to support OMP based algorithms.
+It can be observed that as $N_{\mathrm{R}}$ decreases, the NMSE performance of the BSA-OMP algorithm degrades significantly.
+This is because a small-scale RIS cannot provide enough measurement data diversity to support OMP based algorithms.
 In contrast, the performance of the proposed and CBS-GAMP schemes slightly improve as $N_{\mathrm{R}}$ decreases.
@@
 VI Conclusion
-In this paper, we investigated channel estimation in RIS-assisted wideband systems with consideration of beam split effect.
-We proposed an efficient and low-complexity channel estimation scheme.
-Specifically, we first proposed a novel CBS-GAMP approach to estimate full CSI within a small subset of SCs.
-Then, to reduce computational complexity, we proposed a novel BSAD scheme which consists of two direction estimation algorithm, i.e., EnM and DS-MUSIC to acquire angular information at BS and RIS, respectively.
-With knowledge of angular information, we transformed channel estimation into a simple linear-regression problem and recovered CSI at the remaining SCs using a simple LS algorithm.
-We demonstrated from numerical results that the proposed scheme can achieve superior NMSE performance and significantly reduce pilot overhead.
+In this paper, we investigated channel estimation in RIS-assisted wideband THz systems with consideration of beam split effect.
+We proposed a two-phase channel estimation scheme to achieve accurate CSI acquisition with reduced computational complexity.
+Specifically, in Phase I, we introduced a two-step method to achieve full CSI estimation over two SCs.
+In Phase II, we employed angular information obtained from Phase I to estimate the channels for the remaining SCs using a simple LS algorithm.
+Theoretical analysis and numerical results demonstrated that our scheme achieved near-oracle performance in terms of NMSE.
+Moreover, compared to existing solutions, the proposed two-phase strategy significantly reduced computational complexity while maintaining a high level of accuracy in CSI acquisition.

### v3

@@
 Abstract
-Reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is emerging as a key technology to support ultra-high data rates in future sixth-generation networks.
+Reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is emerging as a key technology to support the ultra-high data rates in future sixth-generation networks.
 However, the acquisition of accurate channel state information (CSI) in such systems is challenging due to the passive nature of RIS and the hybrid beamforming architecture typically employed in THz systems.
 To address these challenges, we propose a novel low-complexity two-phase channel estimation scheme for RIS-assisted THz systems with beam split effect.
-In the proposed scheme, we first estimate the full CSI over a small subset of subcarriers, then extract angular information at both the base station and RIS.
-Subsequently, we recover the full CSI across remaining subcarriers by determining the corresponding spatial directions and angle-excluded coefficients.
+In the proposed scheme, we first estimate the full CSI over a small subset of subcarriers (SCs), then extract angular information at both the base station and RIS.
+Subsequently, we recover the full CSI across remaining SCs by determining the corresponding spatial directions and angle-excluded coefficients.
 Theoretical analysis and simulation results demonstrate that the proposed method achieves superior performance in terms of normalized mean-square error while significantly reducing computational complexity compared to existing algorithms.
@@
 One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [6].
-To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide LoS paths by intelligently manipulating the propagation environment [7].
+To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide virtual LoS path by intelligently manipulating the propagation environment [7].
 RIS can dynamically adjust reflection properties of a large number of passive reflecting elements to optimize signal transmission and reception [8].
 By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems.
-However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), the acquisition of which can be challenging.
+Despite these benefits, the RIS elements remain passive and lack inherent signal processing capabilities.
+Accordingly, the performance enhancement enabled by RIS relies significantly on the accurate channel state information (CSI), which is challenging to acquire.
 Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
-I-A Prior Work
-Over the past few years, numerous research works have been dedicated to exploring channel estimation in RIS-assisted high frequency band systems [9, 10, 11, 12, 13, 14].
-Specifically, a sparse representation of RIS-assisted millimeter wave (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [9].
+I-A Related Works
+Over the past few years, numerous research works have been dedicated to exploring channel estimation in RIS-assisted high frequency systems [9, 10, 11, 12, 13, 14].
+Specifically, a sparse representation of RIS-assisted millimeter wave (mmWave) channel has been established, enabling the application of compressed sensing based techniques for CSI acquisition [9].
 In [10], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel.
-To reduce computational complexity, [11] proposed a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation.
-The authors in [12] proposed a two-phase channel estimation strategy to reduce pilot overhead.
-[13] investigated cooperative-beam-training based channel estimation procedure.
+To reduce computational complexity, [11] proposes a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation.
+The authors in [12] propose a two-phase channel estimation strategy to reduce pilot overhead.
+[13] investigates cooperative-beam-training based channel estimation procedure.
 In [14], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
 The above solutions have effectively addressed the channel estimation problem of narrowband systems with utilization of the angular domain sparsity.
-However, THz communication typically adopts a hybrid beamforming architecture to reduce hardware costs[15].
-When considering this architecture, the beam split effect becomes inevitable.
-This effect occurs due to the wider bandwidth of THz signals and the enormous number of RIS elements, which causes path components to split into distinct spatial directions at different subcarrier frequencies [16].
-This phenomenon leads to a considerable degradation in achievable data rates, which counteracts the benefits of RIS-assisted THz systems.
-To address this issue, beam split effect in RIS-assisted THz systems has been investigated in [17, 18, 19].
-In particular, [17] introduced a beam-split-aware orthogonal matching pursuit (BSA-OMP) approach to automatically mitigate beam split.
+However, mmWave/THz communications typically adopt a hybrid beamforming architecture to reduce hardware costs[15], making beam squint/beam split effect inevitable.
+These effects occur due to the wider bandwidth of mmWave/THz signals and the enormous number of RIS elements, which causes path components to split into distinct spatial directions at different subcarrier (SC) frequencies [16].
+Such phenomenon significantly degrades achievable data rates, thereby counteracting the benefits of RIS-assisted THz systems.
+To address this issue, beam squint/beam split effect in RIS-assisted mmWave/THz systems has been investigated in [17, 18, 19, 20].
+In particular, [17] introduces a beam-split-aware orthogonal matching pursuit (BSA-OMP) approach to automatically mitigate beam split effect.
 In [18], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator.
-The authors in [19] proposed a robust regularized sensing-beam-split-OMP-based scheme to estimate beam split affected RIS-assisted cascaded channel.
-These works primarily focus on mitigating beam split effects within the context of cascaded channels using well-designed dictionaries covering the entire beamspace.
-However, an alternative approach is to focus on the physical directions of the channel components to reduce computational complexity [20, 21, 22].
-Specifically, the authors in [20] regarded channel estimation problem as a block-sparse recovery problem.
-[21] proposed a beam split pattern detection based wideband channel estimation scheme where the physical DoAs are prioritized for determination.
-In [22], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals (AoAs) and angle-of-departure (AoDs) are determined and refined to reconstruct the wideband channel with beam split effect.
-Although these algorithms achieve accurate channel estimation, they require execution at each SC or simultaneous processing of all SCs, which may introduce additional unaffordable computations.
-Moreover, the passive nature of RIS complicates the precise acquisition of angles at the RIS, which often results in the availability of only coupled angles.
+The authors in [19] propose a robust regularized sensing-beam-split-OMP-based scheme to estimate beam split affected RIS-assisted cascaded channel.
+The authors in [20] regard the wideband channel estimation problem as a block-sparse recovery problem by concatenating signals at different SCs.
+These works primarily focus on mitigating beam split effect within the context of cascaded channels using well-designed dictionaries covering the entire beamspace.
+Although they can achieve accurate channel estimation, it requires execution at each SC or simultaneous processing of all SCs, which may introduce additional unaffordable computations.
+An alternative approach is to focus on the physical directions of the channel components to reduce computational complexity [21, 22, 23, 24, 25].
+In mmWave band, angle-based channel estimation methods are used to simplify the wideband channel estimation process under the beam squint effect [21, 22].
+The authors in [21] first estimate angle-of-arrivals (AoAs) at BS and then employed BSA-based algorithms to acquire full CSI.
+[22] presents a twin-stage OMP method to estimate the path angles of the RIS-assisted cascaded two-hop channel.
+[23] introduces a Newtonized OMP (NOMP) based algorithm to detect the channel parameters.
+These approaches perform angle estimation by leveraging the principle that the frequency-dependent spatial angles change while frequency-independent physical angles remain constant across different SCs.
+This principle can also be applied in THz systems, however, due to the significantly larger beam spacing in THz band compared to mmWave systems, the frequency-dependent spatial angles exhibit greater deviations.
+Moreover, the solutions discussed above fail to leverage the wider bandwidth available in THz band to better differentiate between true physical and false spatial angles.
+Consequently, the angle-based methods used in mmWave systems may become neither efficient nor effective when applied to the THz band.
+To this end, several low-complexity estimation schemes have been developed specifically for wideband THz systems [24, 25].
+For example, [24] proposes a beam-split-pattern detection-based estimation method that prioritizes physical DoA determination.
+In [25], a three-stage wideband channel estimation approach is presented, where AoAs and angles-of-departure (AoDs) are successively refined to reconstruct the channel under beam split effect.
+These methods achieve a favorable trade-off between complexity and performance in traditional architectures.
+However, these methods are not readily applicable to RIS-assisted THz systems due to the lack of mechanisms for angular estimation at RIS.
+Moreover, the passive nature of RIS further complicates the accurate acquisition of angular information, making it difficult to resolve the coupled AoAs/AoDs.
 Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
+Figure 1: Illustration of the two-phase channel estimation protocol and corresponding frame structure.
 I-B Main Contributions
-To address the issue discussed above, we presented some preliminary results on estimating the RIS-assisted cascaded channel [23].
+To address the issues discussed above, we presented some preliminary results on estimating the RIS-assisted cascaded channel [26].
 In this paper, we significantly expand on the previous work and propose a two-phase efficient channel estimation scheme for RIS-assisted wideband THz systems with beam split.
 Our main contributions are summarized as follows:
-  • We propose a novel low-complexity channel estimation protocol for general RIS-assisted wideband THz communication scenarios with beam split effect, as depicted in Fig.
-1.
-The channel estimation scheme is divided into two phases: full CSI estimation for two SCs in Phase I and remaining SCs estimation in Phase II.
-By employing information from only two SCs, this protocol can significantly reduce the computational complexity while maintaining complete and accurate channel estimation.
-  • In Phase I, we propose a two-step method to acquire full CSI for the selected two SCs.
-Specifically, in Step 1, we propose a simplified cyclic beam split (CBS) dictionary to reduce the size from the scale of $Q^{2}$ to a linear scale of $Q$.
-Using this dictionary, we reformulate the problem as a generalized sparse recovery problem.
-Moreover, we discuss the selection of different sparse recovery algorithms.
-In Step 2, we first design a heuristic DoD estimation method at BS according to the estimated cascaded channel obtained in Step 1.
-We then reformulate the coupled angles at RIS as a typical DoA estimation problem and design a double sensing-multiple signal classification (DS-MUSIC) algorithm by integrating the angular information from the two selected SCs.
-We also provide insights into the selection of two SCs.
-In Phase II, given angular information from Phase I, we employ a simple least squares (LS) algorithm to estimate the channels for the remaining SCs.
-  • Through the analysis of algorithm complexity, our proposed protocol significantly reduces the computational complexity compared to the existing sparse recovery solutions.
-Additionally, numerical results demonstrate that the NMSE performance of our proposed channel estimation scheme closely approaches that of oracle-LS method.
-By focusing on a two-step estimation process for two SCs and applying a simplified LS algorithm for the remaining SCs, the proposed method can achieve both reduction of computational complexity and acquisition of accurate CSI.
-Figure 1: Two-phase channel estimation protocol and frame structure.
+  • We propose a novel low-complexity channel estimation scheme tailored for RIS-assisted wideband THz communication systems under the beam split effect.
+Unlike conventional methods that rely on full-band sparse recovery across all subcarriers, the proposed scheme adopts a two-phase protocol to achieve accurate CSI estimation with substantially reduced computations.
+Specifically, as illustrated in Fig.
+1, Phase I estimates the full CSI at two carefully selected SCs, while Phase II efficiently reconstructs the remaining CSI using only a simple least squares (LS) algorithm, thereby avoiding the need of computationally expensive recovery procedures across the full bandwidth.
+  • The proposed low-complexity design is primarily enabled by the two-step approach implemented in Phase I.
+In Step 1, we recover the full CSI at two selected SCs utilizing a simplified cyclic beam split (CBS) dictionary, which is compressed from the original size of $Q^{2}$ to a linear scale of $12Q-$ 1, thereby significantly reducing the complexity of sparse recovery.
+In Step 2, we estimate the physical directions at both the BS and RIS based on the recovered CSI.
+Specifically, we propose an energy maximum (EnM) algorithm for directions-of-departure (DoDs) estimation at BS and a double sensing-MUSIC (DS-MUSIC) algorithm to resolve the coupled angles at RIS.
+Instead of mitigating the severe beam misalignment introduced by beam split effect, the proposed algorithm leverages the resulting spatial separation of RIS angles within the extended angular range, which facilitates accurate estimation of physical directions.
+  • We provide a comprehensive complexity analysis, which demonstrates that our proposed protocol significantly reduces computational complexity compared to existing sparse recovery solutions.
+Furthermore, we conduct extensive numerical evaluations under both mmWave and THz system settings.
+The results demonstrate that: (i) the proposed two-phase scheme efficiently exploits the beam split properties to enhance estimation accuracy compared to mmWave solutions; and (ii) it achieves superior NMSE performance over conventional baselines, approaching that of the oracle-LS benchmark.
+These findings highlight the potential of our method for efficient channel estimation in RIS-assisted THz communication systems.
 The remainder of this paper is organized as follows.
-In Section II, we first describe the RIS-assisted wideband THz communication setup and the beam split effect.
-In Section III, we describe our proposed channel estimation protocol.
-Section V provides an analysis of the computational complexity of our proposed scheme, with a comparison to existing methods.
-Finally, we conclude the paper in Section VI.
-Notations: Boldface lower-case and capital letters represent column vectors and matrices, respectively.
-$\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively.
-$\mathbb{E}$ and $\mathrm{Var}$ represent statistical expectation and variance, respectively.
+Section II presents the system model for RIS-assisted wideband THz communication with beam split effect along with the associated channel estimation protocol.
+Section III, IV and V describe the proposed two-phase channel estimation framework in detail.
+Section VI analyzes the computational complexity, and Section VII provides simulation results and performance comparisons with existing methods.
+Finally, Section VIII concludes the paper.
+Notations: Boldface lower-case and capital letters represent the column vectors and matrices, respectively.
+$\left(\cdot\right)^{*}$, $(\cdot)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote the conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively.
+$\mathbb{E}$ and $\mathrm{Var}$ represent the statistical expectation and variance, respectively.
 $\mathbb{N}$, $\mathbb{C}$ and $\mathbb{R}$ denote the set of integers, complex numbers and real numbers, respectively.
-$\mathbf{I}_{N}$ is $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace.
-$\otimes$ denote Kronecker product, * and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
+$\mathbf{I}_{N}$ is the $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace.
+$\otimes$, $\odot$ and $\bullet$ denote the Kronecker product, Khatri-Rao product and transposed Khatri-Rao product, respectively.
+$[\mathbf{A}]_{i,j}$ represents the $(i,j)$-th element of matrix $\mathbf{A}$, and $[\mathbf{a}]_{i}$ represents the $i$-th element of vector or set $\mathbf{a}$.
+Besides, the operator $\text{Diag}(\mathbf{x})$ constructs a diagonal matrix whose diagonal elements are given by the elements of the vector $\mathbf{x}$.
+The notation $a:b$ denotes the sequence $\{a,a+1,\dots,b\}$ with $a\leq b$.
+The imaginary unit is denoted as $\jmath=\sqrt{-1}$.
 II System and Channel Model
 II-A Signal and Channel Model
-Figure 2: RIS-assisted THz communication system with beam split effect.
-We consider a RIS-assisted wideband THz MIMO system employing hybrid beamforming architecture that integrates both analog and baseband processing, as depicted in Fig.
-2.
+As depicted in Fig.
+2, we consider a RIS-assisted wideband THz MIMO system employing hybrid beamforming architecture that integrates both analog and baseband processing.
 Assuming that the BS equipped with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}$ (where $N_{\mathrm{RF}}<N_{\mathrm{T}}$) RF chains is deployed to serve $K$ single-antenna UEs.
 The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS.
-Besides, the OFDM system is considered with $M$ subcarriers, central frequency $f_{c}$ and bandwidth $B$.
+Besides, the orthogonal frequency division multiplexing (OFDM) system is considered with $M$ SCs, central frequency $f_{c}$ and bandwidth $B$.
+Figure 2: Illustration of RIS-assisted THz communication system with beam split effect.
 In this paper, we assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [12].
-For each coherence block, channel estimation is firstly performed during $T$ subframes, and downlink data transmission is then carried out during the remaining $D-T$ subframes.
+For each coherence block, channel estimation is first performed during $T$ subframes, and downlink data transmission is then carried out during the remaining $D-T$ subframes.
 Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots.
-Additionally, we assume that the direct BS-UE links are blocked by obstacles [20].
+Additionally, we assume that the direct BS-UE links are blocked by obstacles [20] [^1] The proposed scheme can be easily extended to one scenario where direct BS-UE and cascaded BS-RIS-UE links coexist.
+The direct channel can be estimated via conventional channel estimation approaches by turning off RIS..
 Therefore, we mainly focus on the downlink channel estimation of the cascaded BS-RIS-UE channel.
 As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[15].
-In our considered system, we assume that $P$ frequency-dependent baseband beamformers $\mathbf{W}_{t}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{t,1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}_{t}\triangleq\mathrm{Dia{\tiny}g}\left\{s_{t,1},\cdots,s_{t,P}\right\}\in\mathbb{C}^{P\times P}$ per time slot.
-Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain improvement in the direction of RIS with each entry is subject to the unit-modulus constraint, i.e., $|[\mathbf{W}^{\mathrm{RF}}]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ for $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$.
-We assume that the analog beamformer remains unchanged over multiple channel coherence blocks due to the relatively static nature of the RIS physical location compared to the dynamic channel coefficients.
-Therefore, we define $\mathbf{W}_{t}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}_{t}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}_{t}[m]=\left[\mathbf{w}_{t,1}[m],...,\mathbf{w}_{t,P}[m]\right]$.
+During $P$ successive time slots, the frequency-dependent baseband beamformer vectors $\left\{\mathbf{w}_{t,p}^{\mathrm{BB}}[m]\right\}_{p=1}^{P}\in\mathbb{C}^{N_{\mathrm{RF}}}$ are sequentially employed combining with the frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$.
+Each entry of $\mathbf{W}^{\mathrm{RF}}$ is subject to the unit-modulus constraint, i.e., $|[\mathbf{W}^{\mathrm{RF}}]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ for $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$.
+We assume that the analog beamformer remains unchanged over multiple channel coherence blocks due to the relatively static nature of the RIS physical location compared to the dynamic channel coefficients [12] [^2] Note that the angular information remains unknown for channel estimation..
+Therefore, we define $\mathbf{w}_{t,p}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{w}_{t,p}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ as the hybrid beamformer vector.
 Let $\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ and $\mathbf{h}_{k}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}}$ denote the channel from BS to RIS and RIS to $k$-th UE at the $m$-th SC, respectively.
 At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by $k$-th UE at the $m$-th SC can be expressed as
-$\begin{split}y_{k,t,p}[m]\!&=\!\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}_{t}\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m]\\ &=\!\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m],\end{split}$ (1)
+$\ y_{k,t,p}[m]=\!\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t}[m]\!+\!n_{k,t,p}[m],$ (1)
 where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN).
-Besides, $\boldsymbol{\theta}_{t}=\left[e^{j\theta_{t,1}},\cdots,e^{j\theta_{t,N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS at subframe $t$ with a constant reflection amplitude constraint [24].
-Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t,p}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
+Besides, $\boldsymbol{\theta}_{t}=\left[e^{\jmath\theta_{t,1}},\cdots,e^{\jmath\theta_{t,N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS at subframe $t$ with a constant reflection amplitude constraint [27].
+Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
 $\tilde{\mathbf{y}}_{k,t}[m]=\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{R}_{t}[m]+\tilde{\mathbf{n}}_{k,t}[m],$ (2)
@@
 We adopt the widely used Saleh-Valenzuela channel model for both BS-RIS channel and RIS-UE channel [29]. The BS-RIS channel $\mathbf{G}[m]$ at the $m$-th SC is expressed as
-$\mathbf{G}[m]=\sum_{l=1}^{L}{\alpha_{l}\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{l})\mathbf{a}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\varphi}_{l}\right)e^{-j2\pi\tau_{l}f_{m}}},$ (4)
-where $L$ represents the number of propagation paths between BS and RIS, $\alpha_{l}$, $\tau_{l}$, $\tilde{\psi}_{l}$ and $\tilde{\varphi}_{l}$ represent the complex gain, the time delay, the physical angle-of-departure (AoD) at BS and angle-of-arrival (AoA) at RIS for the $l$-th path, respectively.
+$\mathbf{G}[m]=\sum_{l=1}^{L}{\alpha_{l}\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{l})\mathbf{a}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\varphi}_{l}\right)e^{-\jmath 2\pi\tau_{l}f_{m}}},$ (4)
+where $L$ represents the number of propagation paths between BS and RIS, $\alpha_{l}$, $\tau_{l}$, $\tilde{\psi}_{l}$ and $\tilde{\varphi}_{l}$ represent the complex gain, the time delay, the physical DoA at BS and DoD at RIS for the $l$-th path, respectively.
 Besides, $f_{m}$ denotes the frequency of the $m$-th SC, which is given by
@@
 Besides, $\mathbf{a}_{\mathrm{T},m}\in\mathbb{C}^{N_{\mathrm{T}}}$ and $\mathbf{a}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}}$ are the normalized array response vectors (ARVs) associated with BS and RIS, respectively, given by
-$\mathbf{a}_{\mathrm{T},m}(\tilde{u})\triangleq\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{u}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
-$\mathbf{a}_{\mathrm{R},m}(\tilde{v})\triangleq\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{v}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
-where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th suncarrier.
+$\mathbf{a}_{\mathrm{T},m}(\tilde{u})\triangleq\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-\jmath 2\pi\sin\left(\tilde{u}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
+$\mathbf{a}_{\mathrm{R},m}(\tilde{v})\triangleq\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-\jmath 2\pi\sin\left(\tilde{v}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
+where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th SC.
 $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively.
-Besides, $\mathbf{n}_{t}=[0,1,\dots,N_{\mathrm{T}}-1]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively.
-By defining the angle vectors $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as
+Besides, $\mathbf{n}_{t}=[0,1,\dots,N_{\mathrm{T}}-1]^{\mathsf{T}}$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]^{\mathsf{T}}$, respectively.
+By defining the angle vectors $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be rewritten as
 $\mathbf{G}[m]=\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\boldsymbol{\varphi}}\right),$ (7)
-where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively.
-$\mathbf{\Sigma}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix.
-Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and UE-$k$ at the $m$-th SC can be formulated as
-$\mathbf{h}_{k}[m]=\sum_{j=1}^{J_{k}}{\beta_{k,j}\mathbf{a}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\vartheta}_{k,j})e^{-j2\pi\mu_{k,j}f_{m}}},$ (8)
-where $J_{k}$ represents the number of propagation paths between RIS and UE-$k$, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the AoD at RIS for the $j$-th path, respectively.
-By defining $\tilde{\boldsymbol{\vartheta}}_{k}=[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form as
+where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the DoDs at BS and DoAs at RIS, respectively.
+$\mathbf{\Sigma}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-\jmath 2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-\jmath 2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix.
+Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and $k$-th UE at the $m$-th SC can be formulated as
+$\mathbf{h}_{k}[m]=\sum_{j=1}^{J_{k}}{\beta_{k,j}\mathbf{a}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\vartheta}_{k,j})e^{-\jmath 2\pi\mu_{k,j}f_{m}}},$ (8)
+where $J_{k}$ represents the number of propagation paths between RIS and $k$-th UE, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the DoD at RIS for the $j$-th path, respectively.
+By defining $\tilde{\boldsymbol{\vartheta}}_{k}=[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}]$, $\mathbf{h}_{k}[m]$ can be rewritten as
 $\mathbf{h}_{k}[m]=\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}),$ (9)
-where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-j2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-j2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE.
-$\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the AoDs at RIS.
+where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-\jmath 2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-\jmath 2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE.
+$\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the DoDs at RIS.
 Substituting (7) and (9) into (3), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be rewritten as
-$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
+$\mathbf{H}_{k}^{\mathrm{cas}}[m]=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).$ (10)
 II-B Beam Split Effect
@@
 This phenomenon is caused by the ultra-wide bandwidth and extensive antenna arrays and reflecting elements in RIS-assisted wideband THz systems.
-Beams generated by the frequency-independent analog beamforming vector and then reflected by the frequency-independent RIS phase shift matrix would split into totally distinct spatial directions.
+The beams generated by the frequency-independent analog beamforming vector and then reflected by the frequency-independent RIS phase shift matrix would split into totally distinct spatial directions.
 Typically, antenna spacings are set to half the wavelength at the central frequency $f_{c}$, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [31].
@@
 Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
-$\mathbf{a}_{\mathrm{T},m}(u)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}u\mathbf{n}_{t}},$ (11a)
-$\mathbf{a}_{\mathrm{R},m}(v)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}v\mathbf{n}_{r}},$ (11b)
-where $\eta_{m}\triangleq f_{m}/f_{c}$ represents the relative frequency. As shown in (11), the spatial directions can be represented by
-$u_{m}=\eta_{m}u,$ (12a)
-$v_{m}=\eta_{m}v.$ (12b)
-In the lower frequency bands such as mmWave, the subcarrier frequencies are relatively closer, which causes slight beam squint with $\eta_{m}\approx 1$ which means that the ARVs at all SC frequencies are almost identical[28, 20].
-However, due to the ultra large bandwidth in THz band, $\eta_{m}$ deviates significantly from one.
-Therefore, the difference between physical directions and spatial directions becomes significant and cannot be ignored.
-In wideband RIS-assisted THz systems, beam split effect occurs at both BS and RIS, which leads to severe performance degradation in channel estimation algorithms [27].
+$\mathbf{a}_{\mathrm{T},m}(u)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-\jmath\pi\eta_{m}u\mathbf{n}_{t}},$ (11a)
+$\mathbf{a}_{\mathrm{R},m}(v)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-\jmath\pi\eta_{m}v\mathbf{n}_{r}},$ (11b)
+where $\eta_{m}\triangleq f_{m}/f_{c}$ represents the relative frequency.
+As shown in (11), the spatial directions can be represented by $u_{m}=\eta_{m}u$ and $v_{m}=\eta_{m}v$.
+In the lower frequency bands, such as mmWave, the SC frequencies are relatively closer, resulting in only minor deviations of $\eta_{m}$ from one.
+Consequently, beam squint remains limited, and the ARVs exhibit only slight variations across different SCs [31, 20].
+In contrast, the ultra-large bandwidth in THz band leads to a significantly larger deviation in $\eta_{m}$ compared to mmWave, making the discrepancy between physical and spatial directions become pronounced and non-negligible.
+In wideband RIS-assisted THz systems, beam split effect occurs at both BS and RIS, resulting in severe performance degradation in channel estimation algorithms [30].
 Therefore, it is crucial to develop an effective channel estimation scheme to acquire accurate CSI under beam split effect.
-III Channel Estimation
-III-A Channel Estimation Protocol
-Traditional channel estimation algorithms require significant redesign of the codebook, as conventional dictionaries fail to cover additional angle information introduced by beam split effect [17].
-This additional angle information will complicates the design of dictionaries.
-Furthermore, the beam split effect disperses the channel for each subcarrier into different spatial directions, which requires well-designed channel estimation algorithms to operate on each subcarrier individually, which can significantly increase computational complexity.
-To address these challenges, we propose a novel low-complexity channel estimation protocol, which is divided into two main phases, as depicted in Fig.
+II-C Channel Estimation Protocol
+Traditional channel estimation algorithms require substantial modifications to the dictionaries, as conventional designs fail to capture the additional angular information introduced by the beam split effect [17].
+Furthermore, the beam split effect disperses the channel of each SC into different spatial directions, thereby necessitating well-designed channel estimation algorithms that operate independently on each SC, which significantly increases the computational complexity.
+To address these challenges, we propose the following low-complexity channel estimation protocol, which is divided into two main phases, as depicted in Fig.
 1.
-The first phase, which is the most challenging aspect of the protocol, mainly focuses on estimating the full CSI for two selected subcarriers, which includes estimation of cascaded channel, the AoDs at BS, and the coupled angles at RIS.
-Due to its complexity, we further divide this phase into two steps.
-In Step 1, we employ sparse recovery algorithms to estimate the cascaded channel for the two subcarriers.
-In Step 2, we focus on heuristic algorithms for common angles estimation.
-Given the full CSI estimation from the first phase, we then employ a simple LS algorithm to estimate the channels for the remaining subcarriers.
-III-B Cascaded Channel Estimation
-In this subsection, we focus on the estimation of cascaded channel for one specific subcarrier.
-In the following steps, we first present the angular domain sparse representation in RIS-assisted THz channels.
-Then, we introduce the simplified dictionary, i.e., CBS dictionary.
+Phase I mainly focuses on estimating the full CSI for two selected SCs, which includes estimation of cascaded channels, the DoDs at BS, and the coupled angles at RIS.
+Specifically, we further divide this phase into two steps.
+In Step 1, we employ sparse recovery algorithms to estimate the cascaded channel for the selected two SCs.
+In Step 2, we focus on the heuristic algorithms for common angles estimation.
+Given the full CSI estimation from Phase I, we then employ a simple LS algorithm to estimate the channels for the remaining SCs in Phase II.
+III Cascaded Channel Estimation for two SCs
+In this section, we focus on the estimation of the cascaded channel for two specific SCs.
+We begin by presenting the angular-domain sparse representation of RIS-assisted THz channels.
+Next, we introduce the simplified dictionary, referred to as the CBS dictionary.
 Based on this simplified dictionary, we reformulate the channel estimation problem as a sparse recovery problem.
-Finally, we provide a detailed analysis of existing sparse recovery algorithms.
-III-B1 Angular Domain Sparse Representation
+Finally, we provide a comprehensive analysis of existing sparse recovery algorithms.
+III-A Angular Domain Sparse Representation
 The sparse angular domain representation of the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ is given by
-$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
-where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS-RIS channel and the angle-excluded vector of RIS-UE channel, respectively.
+$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\bar{\mathbf{X}}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (12)
+where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary.
+Moreover, $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}^{2}\times Q_{\mathrm{T}}}$, where $\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ non-zero entries corresponding to the angle-excluded matrix of BS-RIS channel and the angle-excluded vector of RIS-UE channel, respectively.
+Consequently, $\bar{\mathbf{X}}_{k,m}$ is an $L\times J_{k}$-sparse matrix with its non-zero entries corresponding to the set $\{\!\bar{x}_{k\!,m\!,l\!,j}|\bar{x}_{k\!,m\!,l\!,j}\!\triangleq\!\alpha_{l}\beta_{k,\!j}e^{\!-\jmath 2\pi\!\left(\!\tau_{l}\!+\!\mu_{k,j}\!\right)\!f_{m}}\!,\!l\!=\!1,\!\dots\!,\!L,\!j\!\!=\!1,\!\dots\!,\!J_{k}\!\}$.
 Additionally, the overcomplete dictionaries of the BS and RIS directions can be expressed as
@@
 Moreover, $\bar{\vartheta}_{q}=\delta\cdot q-(Q+1)/{Q}$ with $Q\in\{Q_{\mathrm{T}},Q_{\mathrm{R}}\}$ denotes the spatial direction sample, $\delta=2/Q$ is the interval between two adjacent samples.
-III-B2 Simplified Dictionary
-Note that the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ in (13) is derived from the transposed Khatri-Rao product, which contains a significant number of redundant columns due to the fixed sampling intervals.
-For narrowband channel model, this will result in only $Q_{\mathrm{R}}$ unique columns which are exactly the first $Q_{\mathrm{R}}$ columns [9].
+III-B Simplified Dictionary
+Note that the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains only $Q_{\mathrm{R}}$ unique columns which are exactly the first $Q_{\mathrm{R}}$ columns in narrowband systems[9].
 However, for wideband THz channel with beam split effect, this simplification is no longer suitable due to the frequency-dependent nature of the ARVs.
 Proposition 1.
-The number of unique columns of dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{B}=2Q_{R}-1$, and the coupled dictionary can be simplified by including only the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, which can be given as
+The dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains $Q_{B}=2Q_{R}-1$ unique columns, and the coupled dictionary can be simplified by including only the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, which can be expressed as
 $\mathbf{\Xi}_{\mathrm{R},m}\triangleq[\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(Q_{\mathrm{R}}^{2}-Q_{\mathrm{R}}+1:Q_{\mathrm{R}}^{2}\right);\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(2:Q_{\mathrm{R}}\right)],$ (14)
-where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary.
-Figure 3: Coupled dictionary simplification diagram.
+where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary, and the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
+$\boldsymbol{\xi}_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (15)
 Proof.
-We first define the parameter sequence $\bar{\vartheta}_{q_{1}}=\delta\cdot q_{1}-(Q+1)/{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ which corresponds one-to-one with $Q_{\mathrm{R}}$ ARV columns in the dictionary $\mathbf{C}_{\mathrm{R},m}$.
-Similarly, we define the sequence $\delta_{q_{2}}=-\delta\cdot q_{2}+(Q+1)/Q$ to correspond to $Q_{R}$ ARV columns in the conjugate dictionary $\mathbf{C}_{\mathrm{R},m}^{*}$.
-Thus, the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ can be expressed as $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}})$ for $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
-The $j$-th column of the coupled dictionary $\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}$ can be expressed as
-$\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (16)
-where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
-Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of unique columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$.
-Then, we simplify the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ step by step from (a) to (d) as shown in Fig.
-3.
-As depicted in step (a), we reshape the overcomplete dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ of length $Q_{\mathrm{R}}^{2}$ ARV columns into a three-dimensional matrix where each column represents an ARV.
-Given that ARV vectors can be uniquely determined by known angles in (16), we use the sequences $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$ to index the corresponding ARV columns.
-In step (b), we form the index matrix $\mathbf{Q}\in\mathbb{N}^{Q_{\mathrm{R}}\times Q_{\mathrm{R}}}$ with elements $[\mathbf{Q}]_{i,j}=q_{1}(i)-q_{2}(j)$ for $i,j=1,\dots,Q_{\mathrm{R}}$.
-It is noticed that due to the cyclic shift property of this dictionary, the dictionary base can be represented entirely by combining just the first and last rows.
-This means that we can reconstruct the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ with a sliding window of length $Q_{\mathrm{R}}$.
-Therefore, the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ can be given as in (15).
-Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (16), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
-$\boldsymbol{\xi}_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (17)
-III-B3 Sparse Recovery Reformulation
+The proof is provided in Appendix A. ∎
+III-C Sparse Recovery Reformulation
 By substituting the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ into (12), the RIS-assisted cascaded channel can be rewritten as
-$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\mathbf{\Xi}_{\mathrm{R},m}\mathbf{X}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (18)
-where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$.
+$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\mathbf{\Xi}_{\mathrm{R},m}\mathbf{X}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}}.$ (16)
+Recalling (14) that the simplified dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ is constructed by removing duplicate columns from the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, the corresponding angle excluded matrix $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the merged version of $\bar{\mathbf{X}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$.
+Specifically, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$.
 $\mathcal{P}_{i}$ denotes the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ that match the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$.
-Therefore, there are at most ${L}_{\mathrm{tot},k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$.
+In the cascaded channel, the equivalent angles are the combination of DoDs at BS and coupled angles at RIS, which may overlap and result in fewer effective paths.
+Therefore, there are ${L}_{\mathrm{tot}}\leq L\times J_{k}$ non-zero entries in $\mathbf{X}_{k,m}$.
 Employing the Kronecker product property, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ can be given by
@@
 Note that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix.
-To map this problem to the real number domain for resolution, we then transform the complex form in (21) to the equivalent real form
+To facilitate the application of following sparse recovery algorithms and improve numerical stability [32], we transform the complex form in (19) to the equivalent real form
 [$\left[\begin{array}[]{c}\!\!\!\Re\{\mathbf{y}_{k}\left[m\right]\}\!\!\!\\ \!\!\!\Im\{\mathbf{y}_{k}\left[m\right]\}\!\!\!\\ \end{array}\right]\!\!\!=\!\!\!\left[\begin{matrix}\Re\{\mathbf{\Psi}_{m}\}&\!\!\!\!-\Im\{\mathbf{\Psi}_{m}\}\\ \Im\{\mathbf{\Psi}_{m}\}&\!\!\!\!\Re\{\mathbf{\Psi}_{m}\}\\ \end{matrix}\right]\!\!\left[\begin{array}[]{c}\!\!\!\!\Re\{\mathbf{x}_{k,m}\}\!\!\!\!\\ \!\!\!\!\Im\{\mathbf{x}_{k,m}\}\!\!\!\!\\ \end{array}\right]\!+\!\left[\begin{array}[]{c}\!\!\!\Re\{\mathbf{n}_{k}\left[m\right]\}\!\!\!\!\!\\ \!\!\!\Im\{\mathbf{n}_{k}\left[m\right]\}\!\!\!\!\!\\ \end{array}\right]\!\!,$ (21)
+where ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$ represent the real and imaginary parts, respectively.
 By systematically assigning symbols to each matrix in (21), the sparse recovery problem can be reformulated as
 $\boldsymbol{y}=\mathbf{\Psi}\boldsymbol{x}+\boldsymbol{n}.$ (22)
-For simplicity, we use a simplified subscript notation by omitting the index for the matrices and vectors.
-By defining $T_{\mathrm{tot}}\triangleq PT$ and $Q_{\mathrm{tot}}\triangleq Q_{\mathrm{B}}Q_{\mathrm{T}}$, let $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$ denote the measurement vector, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ denote the observation matrix and $\boldsymbol{x}\in\mathbb{C}^{Q_{\mathrm{tot}}}$ denote the sparse representation vector to be estimated with sparsity ${L}_{\mathrm{tot}}$.
-III-B4 Sparse Recovery Algorithm Analysis
-In designing our channel estimation algorithm, the selection of an appropriate sparse recovery algorithm is crucial due to the unique sparsity characteristics of RIS-assisted THz channels.
-We considered LS, OMP, and GAMP for this purpose.
-a) LS: The LS algorithm, while straightforward, solves the sparse recovery problem by minimizing the residual sum of squares.
-However, its computational complexity, which is $\mathcal{O}(Q_{\mathrm{tot}}^{3})$, will become significantly high for large matrix dimensions.
-Consequently, LS is not practical for the initial full CSI estimation phase due to its high computational demands.
-b) OMP: OMP is a widely used greedy algorithm that reconstructs the signal by iteratively selecting the dictionary atoms that best match the current residual.
-The computational complexity of OMP is $\mathcal{O}(L_{\mathrm{tot}}Q_{\mathrm{tot}}^{2})$, which is relatively efficient.
-However, OMP has a nature of “hard estimation”, which means that it makes definitive decisions at each step without retaining detailed probabilistic information about the signal.
-This approach can lead to suboptimal performance, particularly in noisy environments or with complex signal structures.
-c) GAMP: In contrast to OMP, GAMP uses a probabilistic approach within a message-passing framework to perform global updates, which will maintain a distribution of possible values for each signal element.
-The computational complexity of GAMP is $\mathcal{O}(Q_{\mathrm{tot}}^{2})$ which is similar to OMP but involves more sophisticated probabilistic calculations.
-GAMP’s “soft estimation” process retains probabilistic information throughout the iterations, which leads to more accurate estimation.
-This makes GAMP particularly advantageous in environments with high noise levels or complex signals.
-Despite the similar computational complexity of GAMP and OMP, GAMP’s superior accuracy and robustness make it the preferred choice for initial full CSI estimation [^2] Detailed investigation of CBS-based GAMP scheme has been discussed in our previous work [19].
-Additionally, due to the probabilistic model employed by the “soft estimation” process, GAMP can ensure more reliable angular information determination and avoid the errors caused by “hard estimation”.
-Moreover, for the final stage of channel estimation involving the remaining subcarriers, we employ the LS algorithm.
-The angular information obtained from the initial estimation phase significantly reduces the dictionary matrix dimensions, which will make LS computationally feasible and effective.
-This stage can benefit more from the simplicity and accuracy of LS method.
-III-C Common Angle Estimation
-In this subsection, we focus on the estimation of common angles.
-Using the selected GAMP algorithm from Section III-B, we can obtain the sparse signal vector $\boldsymbol{x}$, denoted as $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$.
+For simplicity, we adopt a simplified subscript notation by omitting the indices for the matrices and vectors.
+Let $T_{\mathrm{tot}}\triangleq 2PT$ and $Q_{\mathrm{tot}}\triangleq 2Q_{\mathrm{B}}Q_{\mathrm{T}}$.
+Define $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ and $\boldsymbol{x}\in\mathbb{C}^{Q_{\mathrm{tot}}}$ as the measurement vector, the observation matrix and the sparse representation vector to be estimated with sparsity ${L}_{\mathrm{tot}}$, respectively.
+This sparse recovery formulation forms the basis for the subsequent channel estimation process.
+In what follows, we analyze different sparse recovery algorithms and motivate the selection of GAMP as the preferred approach for the cascaded channel estimation.
+III-D Sparse Recovery Algorithm Analysis
+Selecting an appropriate sparse recovery algorithm is critical for channel estimation in RIS-assisted THz systems due to the unique sparsity characteristics and error propagation effects.
+This section evaluates three widely used algorithms: LS, OMP, and GAMP.
+a) LS: The LS algorithm solves the linear regression problem by minimizing the residual sum of squares.
+However, it lacks sparsity constraints, which prevents it from effectively capturing the sparse structure of the channel.
+As a result, LS is unsuitable for the initial full CSI estimation phase, where a sparse representation is required.
+b) OMP: OMP is a greedy algorithm that reconstructs the signal by iteratively selecting dictionary atoms that best match the current residual.
+Since it follows a “hard estimation” approach, it makes definitive decisions at each iteration without retaining probabilistic information.
+This approach makes OMP highly sensitive to noise and prone to errors in angle estimation.
+Without a probabilistic model, it cannot compensate for these errors in later processing, which reduces the robustness of the overall estimation framework.
+c) GAMP: Unlike OMP, GAMP adopts a probabilistic message-passing framework that maintains a distribution over signal values throughout the estimation process.
+This “soft estimation” mechanism allows GAMP to incorporate uncertainty in the estimation process and mitigate the impact of noise.
+Instead of selecting a single best estimate at each step, GAMP provides a probability distribution for each angle estimate, which improves the reliability of angular information.
+This characteristic is particularly advantageous in THz channel estimation, where accurate CSI acquisition directly affects overall system performance.
+In summary, GAMP provides superior accuracy and robustness, which makes it the preferred choice for initial full CSI estimation [^4] A detailed investigation of the CBS-based GAMP scheme is provided in our previous work [19]..
+Although GAMP can estimate the full channel across all subcarriers, performing this estimation for every subcarrier individually or jointly would lead to substantial computational complexity.
+To balance estimation accuracy and computational efficiency, we apply GAMP to only two selected subcarriers.
+The estimated angular parameters are then used to reconstruct the CSI for the remaining SCs.
+IV Common Angle Estimation
+Utilizing the selected GAMP algorithm from Section III, we can obtain the sparse signal vector $\boldsymbol{x}$, denoted as $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$.
 By substituting $\hat{\mathbf{x}}_{k,m}$ into (17), the estimated cascaded channel can be expressed as
 $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m}.$ (23)
-This method implies that we could perform the estimation for each subcarrier independently to obtain the channel estimates for all subcarriers.
+This method implies that we could perform the estimation for each SC independently to obtain the channel estimates for all SCs.
 However, this requires implementing the estimation algorithm $M$ times, which would result in an unacceptably high computational complexity.
-While the beam split effect complicates the channel estimation process, it also offers some insights.
-It is noticed in (11) that different subcarriers share common physical directions, with the only different term $\eta_{m}$.
-This indicates that the physical angles remain constant across different subcarriers.
-Inspired by this, we perform the cascaded channel estimation on only a small subset of subcarriers to acquire the common physical angles.
-We select $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]$ for $m\in\{1,M/2\}$.
+While the beam split effect complicates the channel estimation process, it is noticed in (11) that different SCs share common physical directions, with the only different term $\eta_{m}$ [20].
+This indicates that the physical angles remain constant across different SCs, which we refer to as the common angles.
+Inspired by this, we perform the cascaded channel estimation on only a small subset of SCs to acquire the common physical angles.
+We select $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]$ for $m\in\{1,M/2\}$, which will be explained in Section IV-B.
 Given this estimation, we can further estimate the common directions, i.e., the DoDs at the BS and the coupled directions at the RIS.
-III-C1 Common DoDs Estimation at BS
-Given the estimated cascaded channel $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m},\bar{m}\in\left\{1,M/2\right\}$, it is noticed that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\hat{\mathbf{x}}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$.
+IV-A Common DoDs Estimation at BS
+Given the estimated cascaded channel $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m},\bar{m}\in\left\{1,M/2\right\}$, we propose an energy maximum (EnM) algorithm to estimate the common DoDs at BS [^5] The two SCs at positions 1 and $M/2$ are selected due to the requirement of coupled angle estimation at RIS in Section IV-B..
+Notice that while the EnM algorithm can operate on any single SC, leveraging CSI from multiple SCs improves estimation accuracy [22].
+Therefore, we accumulate the CSI obtained from cascaded channel estimation in Step 1.
+It is noticed that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\hat{\mathbf{x}}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$.
 This means that given the accurate estimation of $\hat{\mathbf{x}}$, we only need to determine its non-zero positions corresponding to the columns in the matrix $\mathbf{\Pi}_{\bar{m}}$.
 Since $\mathbf{\Pi}_{\bar{m}}\triangleq\left(\mathbf{C}_{\mathrm{T},\bar{m}}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},\bar{m}}\right)$ is in one-to-one correspondence with the ARVs, this relationship allows us to directly obtain the DoDs at BS.
-However, inevitable noise and estimation errors can cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, which will lead to more than $\Sigma_{k}$ non-zero entries.
+However, inevitable noise and estimation errors can cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, which will lead to more than ${L}_{\mathrm{tot}}$ non-zero entries.
 To address this, we propose a heuristic method to determine the accurate non-zero positions, referred to as common physical support, and calculate the corresponding physical DoDs at the BS, as summarized in Algorithm 1.
-Algorithm 1 DoDs Estimation at BS
+Algorithm 1 EnM Based DoDs Estimation at BS
 0:  Angle-excluded matrix $\mathbf{x}_{k,\bar{m}},\bar{m}\in\left\{1,M/2\right\}$, the path number $L$ of BS-RIS channel and the quantization level of overcomplete dictionaries $Q_{\mathrm{R}}$ and $Q_{\mathrm{T}}$.
@@
 In the proposed method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, where each entry is the normalized value of the corresponding element in $\hat{\mathbf{x}}_{k,\bar{m}}$ (Step $2$).
-As observed in Section III-B3, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, each containing $Q_{\mathrm{B}}$ columns.
+As observed in Section III-C, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, each containing $Q_{\mathrm{B}}$ columns.
 Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way.
 By accumulating the $\ell_{2}$-norm for each block, we form the merged-energy vector $\boldsymbol{\epsilon}$ to establish a one-to-one correspondence with the physical DoDs at BS (Step $3-5$).
-The indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ will form the common physical support $\mathcal{I}^{\mathrm{BS}}$ (Step $6$).
-Finally, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ can be obtained using the knowledge of $\mathcal{I}^{\mathrm{BS}}$ and the sampling rule of the CBS dictionary (Step $7$).
+Then, the indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ will form the common physical support $\mathcal{I}^{\mathrm{BS}}$ (Step $6$), where the function $\mathrm{maxk}\left(\boldsymbol{\epsilon},L\right)$ returns the $L$ largest elements in vector $\boldsymbol{\epsilon}$ Finally, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ can be obtained using the knowledge of $\mathcal{I}^{\mathrm{BS}}$ and the sampling rule of the CBS dictionary (Step $7$).
 Fig.
@@
 Figure 3: Normalized energy spectrum via EnM algorithm.
-III-C2 Common Coupled Angle Estimation at RIS
+IV-B Common Coupled Angle Estimation at RIS
 With the estimated DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$, the estimated cascaded channel $\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}]$ can be projected onto the subspace of ARM of BS using the pseudo-inverse, which can be given by
 $\begin{split}\mathbf{U}_{k,\bar{m}}&\triangleq\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}](\hat{\mathbf{A}}_{N_{\mathrm{T}},\bar{m}}^{\mathsf{H}}\left(\boldsymbol{\varphi}\right))^{\dagger}\\ &=\mathrm{Diag}\left(\mathbf{h}_{k}[\bar{m}]\right)\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\psi}\right)\mathbf{\Sigma}_{\bar{m}},\end{split}$ (24)
-where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$, and $\mathbf{U}_{k,\bar{m}}\triangleq\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ denotes the rest-CSI matrix, which is only related to the AoAs and AoDs at RIS.
+where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$, and $\mathbf{U}_{k,\bar{m}}\triangleq\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ denotes the rest-CSI matrix, which is only related to the DoAs and DoDs at RIS.
 By substituting (7) into (24), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
 $\begin{split}\mathbf{u}_{k,\bar{m},l}&=\Sigma_{\bar{m},l}\left(\mathbf{A}_{\mathrm{R},\bar{m}}^{*}\left(\boldsymbol{\vartheta}_{k}\right)\boldsymbol{\beta}_{k,\bar{m}}\right)\bullet\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\right)\\ &=\Sigma_{\bar{m},l}\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,1}\!\right),\dots,\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,J_{k}}\!\right)\right]\!\boldsymbol{\beta}_{k,\bar{m}},\end{split}$ (25)
-where $\Sigma_{\bar{m},l}$ denotes the $\left(l,l\right)$-th element in $\Sigma_{\bar{m}}$, and $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right]$ represents the coupled angle for $l=1,\dots,L$ and $j=1,\dots,J_{k}$.
-By defining $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM, (27) can be rewritten as
+where $\Sigma_{\bar{m},l}$ denotes the $\left(l,l\right)$-th element in $\mathbf{\Sigma}_{\bar{m}}$, and $\gamma_{k,l,j}=\psi_{l}-\vartheta_{k,j}\in\left[-2,2\right]$ represents the coupled angle for $l=1,\dots,L$ and $j=1,\dots,J_{k}$.
+Define $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a_{\mathrm{R},\bar{m}}}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM, (25) can be rewritten as
 $\mathbf{u}_{k,\bar{m},l}=\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l,j}\right)\boldsymbol{\chi}_{k,\bar{m},l},$ (26)
-where $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with the path gains and phase shifts for each coupled angle.
+where $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with path gains and phase shifts for each coupled angle.
 The expression in (26) represents a typical DoA estimation problem, which can be effectively addressed using the MUSIC algorithm due to its high resolution and robustness against noise [33].
@@
 This results in an increased number of potential estimated angles, making it difficult to distinguish the true physical directions.
-Secondly, the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC can cause energy leakage into the noise space, which further complicates the identification of physical directions.
-To address these challenges, we combine the information from two selected subcarriers and apply spatial smoothing preprocessing to the projected rest-CSI matrix $\mathbf{U}_{k,\bar{m}}$.
-Inspired by the smooth-MUSIC algorithm, we propose a novel DS-MUSIC algorithm to accurately estimate the coupled angles, as described in Algorithm 2, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
+Secondly, the approximate orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements per SC can cause energy leakage into the noise space, further complicating the identification of physical directions.
+To address these challenges, we propose a novel double-sensing MUSIC (DS-MUSIC) algorithm.
+Firstly, we combine the information from two selected SCs and apply spatial smoothing preprocessing to the projected rest-CSI matrix $\mathbf{U}_{k,\bar{m}}$.
+Inspired by the spatial smoothing MUSIC algorithm [34], we then introduce the proposed DS-MUSIC algorithm to accurately estimate the coupled angles, as described in Algorithm 2, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
 We first preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray containing $N_{\mathrm{sub}}$ reflecting elements.
-There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total, which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}}$, $N_{\mathrm{sub}}>J_{k}$.
+There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ overlapping subarrays in total.
+To ensure reliable angle estimation, we enforce the condition $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}},N_{\mathrm{sub}}>J_{k}$ [^6] Increasing $N_{\mathrm{a}}$ provides more snapshots for constructing the sample covariance matrix, which improves its estimation accuracy and the stability of the MUSIC algorithm; on the other hand, a larger $N_{\mathrm{sub}}$ improves the angular resolution by increasing the aperture of each subarray.
+Since $N_{\mathrm{a}}+N_{\mathrm{sub}}-1=N_{\mathrm{R}}$ is fixed, these parameters must be jointly chosen to balance spatial resolution and covariance estimation accuracy..
 The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
@@
 where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(:,n:n+N_{\mathrm{sub}}-1\right)$.
+Additionally, $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with the path gains and phase shifts for each coupled angle.
 Therefore, the covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ can be written as
@@
 Algorithm 2 DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
-[…]
-1:  $\bar{\boldsymbol{\gamma}}=-2\left(1+\boldsymbol{q}/Q_{\mathrm{R}}\right),\boldsymbol{q}=1,\dots,Q_{\mathrm{B}}$.
+0:  Projected matrix $\mathbf{U}_{k,\bar{m}}$, the path number $L$ and $J_{k}$ of the BS-RIS and RIS-$k$-th UE channel $\forall k\in\mathcal{K}$, the number of subarrays $N_{\mathrm{a}}$ and quantization level $Q_{\mathrm{R}}$.
+1:  $[\bar{\boldsymbol{\gamma}}]_{q}=-2\left(1+q/Q_{\mathrm{R}}\right),q=1,\dots,Q_{\mathrm{B}}$.
 2:  for $l\in\left\{1,2,\dots,L\right\}$ do
@@
 8:        for $q\in\left\{1,2,\dots,Q_{\mathrm{B}}\right\}$ do
-9:           $\tilde{p}_{q}^{\mathrm{MUSIC}}=\frac{1}{\left\|\left(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\right)^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}}$.
+9:           $\tilde{\mathbf{p}}_{k,\bar{m},l,q}=\frac{1}{\left\|\left(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\right)^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}}$.
 10:        end for
@@
 $\mathbf{R}_{k,\bar{m},l}^{u}=\frac{1}{N_{\mathrm{a}}}\sum_{n=1}^{N_{\mathrm{a}}}{\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}}=\frac{1}{N_{\mathrm{a}}}\tilde{\mathbf{U}}_{k,\bar{m},l}\tilde{\mathbf{U}}_{k,\bar{m},l}^{\mathsf{H}},$ (29)
-where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace-supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ as defined in (29).
-Then, the eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ can be expressed as
+where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ denotes the concatenation of the subspace-supporting vectors $\mathbf{u}_{k,\bar{m},l,n}$ as defined in (27).
+The eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ can be expressed as
 $\mathbf{R}_{k,\bar{m},l}^{u}=\tilde{\mathbf{\Omega}}_{k,\bar{m},l}\mathrm{Diag}(\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l})\tilde{\mathbf{\Omega}}_{k,\bar{m},l}^{\mathsf{H}},$ (30)
@@
 The part $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ consists of the $J_{k}$ largest eigenvalues corresponding to the directions of largest variability, which spans the signal subspace.
-The remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors form $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, spans the noise subspace and is orthogonal to the signal subspace.
+The $N_{\mathrm{sub}}-J_{k}$ remaining eigenvectors, which form $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, span the noise subspace and are orthogonal to the signal subspace.
 To calculate the MUSIC pseudo-spectrum, the practical implementation requires sampling the estimated angles.
-We define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}\triangleq[\tilde{p}_{k,\bar{m},l,1},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}]^{\mathsf{T}}$ as the MUSIC pseudo-spectrum, where $\tilde{p}_{k,\bar{m},l,q}$ is the sampling spectral value for $q=1,\dots,Q_{\mathrm{B}}$.
+We define $\mathbf{\tilde{p}}_{k,\bar{m},l}\triangleq[\tilde{p}_{k,\bar{m},l,1},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}]^{\mathsf{T}}$ as the MUSIC pseudo-spectrum, where $\tilde{p}_{k,\bar{m},l,q}$ is the sampling spectral value for $q=1,\dots,Q_{\mathrm{B}}$.
 To align these sampled values with our proposed algorithm, we adjust the angle sampling within the range of the CBS dictionary.
-Therefore, we define the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the first $N_{\mathrm{sub}}$ rows of the original CBS dictionary in (15).According to (17), each element in the sub-CBS dictionary $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ can be derived as
-$\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\triangleq e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}},q=1,\dots,Q_{\mathrm{B}},$ (34)
+Therefore, we define the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the first $N_{\mathrm{sub}}$ rows of the original CBS dictionary in (14).
+According to (15), each element in the sub-CBS dictionary $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ can be derived as
+$\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\triangleq e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}},q=1,\dots,Q_{\mathrm{B}},$ (32)
 where $\mathbf{n}_{\mathrm{sub}}=\left[1,\cdots,N_{\mathrm{sub}}\right]$.
-By mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the MUSIC algorithm constructs the pseudo-spectrum by evaluating the inverse of the projection’s norm at each potential direction, which can be given as
-$\tilde{p}_{k,\bar{m},l,q}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}}.$ (35)
-Given that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ in the signal subspace should ideally have minimal projection onto the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which means that the $\ell_{2}$-norm in (35) associated with the actual directions should approximate zero.
+By mapping the sampled beams $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the MUSIC algorithm constructs the pseudo-spectrum by evaluating the inverse of the projection’s norm at each potential direction, which can be given as
+$\tilde{p}_{k,\bar{m},l,q}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}}.$ (33)
+Therefore, the MUSIC pseudo-spectrum can be express as
+$\mathbf{p}_{k,\bar{m},l}=\left[\tilde{p}_{k,\bar{m},l,1},\cdots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}\right].$ (34)
+Given that any beam $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ in the signal subspace should ideally have minimal projection onto the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which means that the $\ell_{2}$-norm in (33) associated with the actual directions should approximate zero.
 The MUSIC algorithm typically relies on peak detection to find actual physical directions.
-To avoid undetectable marginal peaks, we extend the pseudo-spectrum by inserting the minimum value at its edges.
-The extended pseudo-spectrum can be express as
-$\mathbf{p}_{k,\bar{m},l}=\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right),\tilde{\mathbf{p}}_{k,\bar{m},l},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right)\right].$ (36)
 The potential directions can then be determined by employing local-maximum algorithms.
-However, due to the extended range of coupled angles, the estimated number of directions will be $K_{\text{cp}}$ times the original.
+However, due to the extended range of coupled angles, the estimated number of directions will be $K_{m,\text{cp}}$ times the original with $K_{\text{cp}}<K_{m,\text{cp,max}}$.
 Proposition 2.
 The maximum number of identical beams can be given as
-$K_{\rm{cp},\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (37)
+$K_{m,\rm{cp},\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1,$ (35)
+[…]
 Proof.
-Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
-$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{\text{cp}}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{\text{cp}}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (38)
-where $K_{\text{cp}}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$.
-It is observed that when $q^{\prime}=q+K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, the element in sub-CBS dictionary will generate the identical beam, i.e., $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}=\xi_{\mathrm{R},m,q^{\prime}}^{\mathrm{sub}}$.
-Therefore, the maximum number of identical beams $K_{\text{cp},\max}$ can be obtained by solving the integer equation $K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$.
+Recalling from (32) that the sampled beams $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}=e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}$.
+Based on the periodicity property of the complex exponential function, which satisfies $e^{-\jmath 2\pi K_{m,\text{cp}}}=1$ for any integer $K_{m,\text{cp}}\in\mathbb{N}$, $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be equivalently expressed as
+$\begin{split}\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-\jmath 2\pi K_{m,\text{cp}}\mathbf{n}_{\mathrm{sub}}}e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{m,\text{cp}}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}}.\end{split}$ (36)
+It is observed that when $q^{\prime}=q+K_{m,\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, the element in sub-CBS dictionary will generate the identical beam, i.e., $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}=\boldsymbol{\xi}_{\mathrm{R},m,q^{\prime}}^{\mathrm{sub}}$.
+Therefore, the maximum number of identical beams $K_{m,\text{cp},\max}$ can be obtained by solving the integer equation $K_{m,\text{cp}}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$.
 By rearranging the terms and employing the relationship $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$, we can obtain
-$K_{\text{cp}}=\frac{\left(2Q_{\mathrm{R}}-1\right)\eta_{m}}{Q_{\mathrm{R}}}.$ (39)
-Therefore, the maximum value of $K_{\text{cp}}$ is the integer part of this expression, which can be given as
-$K_{\rm{cp},\max}=\left\lfloor\eta_{m}\left(2-\frac{1}{Q_{\mathrm{R}}}\right)\right\rfloor+1.$ (40)
+$K_{m,\text{cp}}=\frac{Q_{\mathrm{B}}\eta_{m}}{Q_{\mathrm{R}}}.$ (37)
+Therefore, the maximum value of $K_{m,\text{cp}}$ is the integer part of this expression, which can be given as
+$K_{m,\rm{cp},\max}=\left\lfloor\eta_{m}\left(2-\frac{1}{Q_{\mathrm{R}}}\right)\right\rfloor+1.$ (38)
 According to Proposition 2, instead of the expected $J_{k}$ peaks in the MUSIC pseudo-spectrum, $K_{\text{cp},\max}J_{k}$ peaks will be detected.
 This makes the algorithm infeasible for distinguishing the true physical directions from fake ones within a single SC.
-Thankfully, it is observed from (37) that the spatial directions associated with true physical directions keep frequency-independent.
+Thankfully, it is observed from (36) that the period $K_{m,\text{cp}}/\eta_{m}$ varies with SC frequency.
+Therefore, the periodic false spatial directions are frequency-dependent, while the true physical directions remain frequency-independent [20].
 Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction.
-In this approach, we select two SCs to calculate and compare their extended MUSIC pseudo-spectrum.
-The positions where peaks overlap indicate the true physical directions.
-However, due to the unavoidable noise and estimation errors, the pseudo-spectrum of different SCs may exist offsets.
-Therefore, in practical implement, we select the $J_{k}$ closest peaks to represent true physical directions.
-We introduce the peak distance $d$ to measure the distance between two peaks, which is given by
+In this approach, we select two SCs to calculate and compare their MUSIC pseudo-spectrum, as these two SCs provide the necessary information to uniquely determine the coupled angles at RIS.
+Assuming $\theta_{q_{\mathrm{p}}}$ to be the true physical direction, the corresponding sampled-beam $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q_{\mathrm{p}}}$ can be determined according to (33).
+Then, the unique angular index $q_{\mathrm{p}}$ can be easily decoupled from (32), which is frequency-independent and remains consistent across all the SCs[20].
+In contrast, the index of periodic false angles, given by $q^{\prime}_{\mathrm{p}}=q_{\mathrm{p}}+K_{m,\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, varies among different SCs.
+Therefore, the positions where peaks overlap indicate the true physical directions.
+However, due to unavoidable noise and estimation errors, the pseudo-spectrum across different SCs may exhibit offsets.
+To estimate the true physical directions accurately, we first identify the $J_{k}$ closest peaks, which correspond to the overlapping peaks across different SCs.
+Therefore, we introduce the peak distance $d$ to quantify this deviation, given by:
 $d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=\left|q_{m}-q_{m^{\prime}}^{\prime}\right|,$ (39)
@@
 So far, the detailed steps of the proposed DS-MUSIC algorithm are summarized in Algorithm 2.
-To ensure the algorithm is both effective and efficient, it is crucial to select the appropriate SCs.
-Since $K_{\text{cp},\max}$ in (37) is affected by the choice of SCs, we should aim to satisfy $K_{\text{cp},\max}\leq 2$, which means each angle has at most two identical peaks.
+Selecting the closest peaks ensures consistency in the estimated directions and mitigates the impact of frequency-dependent variations.
+However, while the closest peaks are used for estimating the true directions, it is also essential to ensure that 1) the number of non-overlapping peaks are limited, and 2) these non-overlapping peaks are sufficiently separated.
+Satisfying these conditions allows effective discrimination between true and false physical directions.
+This can be achieved by selecting the appropriate SCs.
+Specifically, since $K_{m,\text{cp},\max}$ in (35) is affected by the choice of SCs, we should aim to satisfy $K_{m,\text{cp},\max}\leq 2$, which means each angle has at most two identical peaks.
 Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into (38), we can obtain
-$m\leq\frac{M}{2}+\frac{Mf_{c}Q_{\mathrm{R}}}{B\left(Q_{\mathrm{B}}\right)}+\frac{1}{2}.$ (42)
+$m<\frac{M}{2}+\frac{Mf_{c}}{BQ_{\mathrm{B}}}+\frac{1}{2}.$ (40)
 Besides, substituting (5) into (39), the peak distance $d$ can be derived as
-$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{\text{cp}}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (43)
-To distinguish true physical directions effectively, the peak distance should also be sufficiently large.
+$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{m,\text{cp}}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (41)
+To distinguish true physical directions effectively, the distance among non-overlapping peaks should also be sufficiently large.
 According to (41), it is evident that the intervals between selected SCs should be large enough.
-Therefore, under the constraint in (42), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation.
+Therefore, under the constraint in (40), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation [^7] Note that while using more subcarriers may offer additional spatial information, our two-subcarrier design leverages the deterministic beam split property for efficient peak alignment.
+Extending the alignment strategy to more than two subcarriers would require complex correspondence matching, which may introduce ambiguity and computational overhead without significant accuracy improvement..
 Fig.
-5 depicts an example of estimated extended MUSIC pseudo-spectrum with the proposed DS approach.
+4 depicts an example of estimated MUSIC pseudo-spectrum utilizing the proposed DS approach.
 Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for an arbitrary UE with SNR $=20$ dB.
@@
 The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
-Figure 5: Normalized extended MUSIC pseudo-spectrum via DS-MUSIC algorithm.
-III-D Channel Estimation for Other Subcarriers
-Through the proposed two-step full CSI estimation, we obtained the cascaded channels for two subcarriers and the corresponding common physical angle.
+Figure 4: Normalized MUSIC pseudo-spectrum via DS-MUSIC algorithm.
+V Channel Estimation for Other SCs
+Through the proposed two-step full CSI estimation, we obtained the cascaded channels for two SCs and the corresponding common physical angle.
 Based on this knowledge, the cascaded channel in (10) can be rewritten as
@@
 where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\bar{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$.
-The column size of the $\tilde{\mathbf{\Psi}}_{m}$ matrix is reduce from $Q_{\mathrm{B}}Q_{\mathrm{T}}$ of overcomplete dictionary to a multipath-related size $L^{2}J_{k}$.
-This significant reduction in the column size enhances both the efficiency and feasibility of the LS solution.
-Therefore, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained as
+The column size of the $\tilde{\mathbf{\Psi}}_{m}$ matrix is reduced from $Q_{\mathrm{B}}Q_{\mathrm{T}}$ of the overcomplete dictionary to a multipath-related size $L^{2}J_{k}$.
+Moreover, the incorporation of angular information transforms the observation matrix $\tilde{\mathbf{\Psi}}_{m}$ transformed into a full-rank matrix.
+Consequently, the sparse representation in (19) is reverted to the non-sparse frequency domain.
+This significant reduction in the column size and reconstruction of signal structure reformulate the channel estimation problem of remaining SCs a low-dimensional linear regression problem, enhancing both the efficiency and feasibility of the LS solution.
+Therefore, by employing LS algorithm, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained as
 $\hat{\tilde{\mathbf{x}}}_{k,m}=\left(\tilde{\mathbf{\Psi}}_{m}^{\mathsf{H}}\tilde{\mathbf{\Psi}}_{m}\right)^{-\mathsf{1}}\tilde{\mathbf{\Psi}}_{m}^{\mathsf{H}}\mathbf{y}_{k,t}\left[m\right].$ (44)
@@
 $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\otimes\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\hat{\tilde{\mathbf{x}}}_{k,m}.$ (45)
-IV Analysis of Computational Complexity
-In this section, we analyze the computational complexity of the proposed channel estimation protocol.
-We also compare our results with the other existing algorithms.
-Specifically, the complexity of the proposed scheme mainly stems from four major components:
+VI Analysis of Computational Complexity
+In this section, we analyze the computational complexity of the proposed channel estimation protocol and compare it with that of existing algorithms.
+Specifically, the overall complexity of the proposed scheme is primarily attributed to the following four components:
   1. 1.
-    Cascaded channel estimation for two selected SCs by CBS-based GAMP approach: $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
+    Cascaded channel estimation for two selected SCs by CBS-based GAMP approach: $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}N_{\mathrm{iter}}\right)$, where $N_{\mathrm{iter}}$ denotes the number of iterations required for GAMP to converge.
   2. 2.
@@
   3. 3.
-    Coupled directions acquisition at the RIS using DS-MUSIC algorithm: $\mathcal{O}\left(2N_{r}^{3}\right)$.
+    Coupled directions acquisition at the RIS using DS-MUSIC algorithm: $\mathcal{O}\left(2N_{\mathrm{R}}^{3}\right)$.
   4. 4.
@@
 The overall complexity of the proposed scheme can be expressed as
-$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right).$ (48)
+$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}N_{\mathrm{iter}}+2N_{\mathrm{R}}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right).$ (46)
 For comparison, we also discuss the complexity of the conventional OMP scheme [35], BSA-OMP scheme [17] and CBS-GAMP scheme [26].
 These algorithms perform under traditional channel estimation frameworks which means they need to be executed at each SC or simultaneously process all SCs.
-The computational complexity of conventional OMP algorithm is $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$.
-The BSA-OMP approach has the same complexity as the conventional OMP, $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$, since it similarly processes each subcarrier independently.
-The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
+The computational complexity of conventional OMP algorithm is $\mathcal{O}\left(MPTQ_{\mathrm{T}}{Q_{\mathrm{R}}^{2}}\right)$.
+The BSA-OMP approach has the same complexity as the conventional OMP, $\mathcal{O}\left(MPTQ_{\mathrm{T}}{Q_{\mathrm{R}}^{2}}\right)$, since it similarly processes each SC independently.
+The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}N_{\mathrm{iter}}\right)$.
 Additionally, the number of SCs $M$ is typically large in wideband systems.
 Therefore, it is evident that the computational complexity of the proposed scheme is significantly lower than the other approaches mentioned.
-V Numerical Results and Discussion
-In this section, we present numerical results to validate the effectiveness of the proposed scheme.
-Without loss of generality, we consider a wideband RIS-assisted THz scenario with central carrier frequency $f_{c}=100$ GHz, bandwidth $B=15$ GHz and $M=128$ SCs.
+VII Numerical Results and Discussion
+In this section, we present numerical results compared with both mmWave and THz benchmarks to validate the effectiveness of the proposed scheme.
+Without loss of generality, we consider wideband RIS-assisted communication scenarios with $M=128$ SCs.
 A BS employing a ULA with $N_{\mathrm{T}}$ = 16 antennas transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements.
-For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths, and the AoAs and AoDs are randomly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$.
-The quantization level of overcomplete dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$.
-We use the normalized mean square error (NMSE) as a metric for CSI estimation performance, which is given by
+Moreover, the hybrid beamformer vector $\mathbf{w}_{t,p}[m]$ is assumed to be randomly distributed [17].
+VII-A Comparison with mmWave Solution
+In this subsection, we present simulation results to illustrate the distinct characteristics and impacts of beam squint and beam split effects, and further demonstrate the unique advantages of the proposed framework in comparison with mmWave solutions.
+VII-A1 Setup
+For the mmWave system, the central carrier frequency is set to $f_{c}=28$ GHz with a bandwidth of $B=600$ MHz, following the configuration in [23].
+For the considered THz system, the central carrier frequency is increased to $f_{c}=100$ GHz with a wider bandwidth of $B=15$ GHz.
+Furthermore, the channel characterization adopts the same parameters as those specified in [24], and the path gains are model as $\alpha_{l}\sim\mathcal{CN}(0,1)$ and the time delays are constrained by $\tau_{l}\leq\tau_{\max}$ with $\tau^{\mathrm{mmWave}}_{\max}=53$ ns and $\tau^{\mathrm{THz}}_{\max}=20$ ns for mmWave and THz systems, respectively.
+Additionally, we assume that the DoDs at BS are available as prior knowledge in this subsection, since the considered benchmark focused only on estimating the angles at RIS.
+Accordingly, the number of propagation paths is set to $L=1$ and $J_{k}=2,\forall k$, respectively.
+Furthermore, the DoAs and DoDs are randomly generated within the range $\left[-\pi/2,\pi/2\right]$.
+To further highlight the effect of beam squint and beam split, we consider the DoDs at RIS to be under large-angle reflection conditions, which significantly deviate from the normal incidence.
+VII-A2 Benchmark
+We adopt the NOMP based mmWave solution proposed in [23] as the benchmrk method. This approach is implemented for both mmWave and THz band, referred to as ‘NOMP-mmWave’ and ‘NOMP-THz’, respectively.
+VII-A3 Metric
+The root mean square error (RMSE) is utilized to evaluate the accuracy of angle estimation, defined as
+$\mathrm{RMSE}=\sqrt{\mathbb{E}\left\{\left|\theta-\hat{\theta}\right|\right\}},$ (47)
+where $\theta$ and $\hat{\theta}$ represent the true angle and the estimated angle, respectively.
+VII-A4 Performance Analysis
+For a fair comparison, identical pilot resources are allocated to both the NOMP-based solution and the proposed approach for estimating cascaded angles at RIS.
+Let $M_{\mathrm{sub}}$ denote the number of SCs used for angle estimation.
+The resource parameters of NOMP algorithm are set to $T=50$ and $M_{\mathrm{sub}}=12$, with the hybrid beamforming functionality disabled, consistent with the original implementation settings in [23].
+Accordingly, the proposed method utilizes $T=25$, $P=12$ and $M_{\mathrm{sub}}=2$ resource elements.
+Fig.
+5 shows the RMSE performance of the coupled DoAs/DoDs estimation at RIS.
+While the NOMP-based mmWave solution achieves satisfactory performance in conventional mmWave scenarios, it fails to adequately address the beam split effect in THz systems.
+This degradation is primarily attributed to the substantially larger SC spacing in THz band, which induces more severe beam misalignment compared to the relatively mild beam squint effect observed at mmWave frequencies.
+As a result, conventional mmWave solutions fail to mitigate this misalignment by merely accumulating information from multiple SCs.
+In contrast, the proposed approach integrates a specifically designed DS-MUSIC algorithm, which distinguishes the true angular components from the frequency-dependent false ones by leveraging the inter-SC misalignment within an extended angular range, thereby achieving significantly enhanced estimation accuracy.
+Moreover, increasing the angular dictionary size $Q_{\mathrm{R}}$ leads to additional performance gains by mitigating quantization errors in the angle domain.
+Figure 5: RMSE performance of coupled DoAs/DoDs at RIS against SNR.
+VII-B Comparison with THz Solutions
+In this subsection, we investigate the impact of different system parameters on performance, and provide a comprehensive comparison with state-of-the-art algorithms in THz band.
+VII-B1 Setup
+We consider the THz system as defined in Section VII-A.
+Without loss of generality, the DoAs and DoDs are randomly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$ in this subsection [18]  [^8] The quantization errors introduced by discrete angular grids assumption are small for large dictionary quantization level $Q_{\mathrm{T}}$ and $Q_{\mathrm{R}}$, which do not affect significantly the normalized array gain and can be neglected..
+Consequently, the quantization level of overcomplete dictionaries can be reduced to $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$ for simplicity, and the 1-st and 64-th SCs are selected for the full CSI estimation in Phase I Step 1.
+Furthermore, we consider a more practical multipath model where the location of both BS and RIS is unknown and there exists $L=J_{k}=3,\forall k$ propagation paths for both BS-RIS and RIS-UEs channels.
+VII-B2 Benchmarks
+For comparison, we also evaluate the following benchmark schemes in the simulation. Unless otherwise specified, they are configured identically to the proposed method to ensure a fair comparison.
+  • Conventional-OMP [35]: Traditional dictionaries are adopted to span the beam space and the CSI is reconstructed by OMP algorithm at each SC.
+  • BSA-OMP [17]: Beam-split-adapted dictionaries are employed to cover the beam space and the CSI is recover by OMP algorithm at each SC.
+  • CBS-GAMP [26]: CBS dictionaries are employed to cover the beam space and the CSI is recover by GAMP algorithm at each SC.
+  • Oracle-LS: The angular information is assumed perfectly known and the angle-excluded information is estimated by LS algorithm.
+This algorithm is regarded as the performance upper bound of the proposed scheme.
+VII-B3 Metrics
+We use the normalized mean square error (NMSE) as a metric for full CSI estimation performance, which is given by
 $\mathrm{NMSE}=\mathbb{E}\left\{\frac{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]-\mathbf{\hat{h}}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}\right\}.$ (48)
 Additionally, we employ the correct probability as a metric for direction estimation performance, which is given by
-$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\},$ (50)
-where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ is the threshold which is set as half the sample interval.
-For comparison, we also consider the conventional OMP scheme [30], BSA-OMP scheme [17], CBS-GAMP scheme [23], and oracle-LS scheme [31] as benchmark schemes in the simulation.
-It is noted that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
+$\mathrm{P}_{\mathrm{c}}=\mathbb{E}\left\{\mathrm{Pr}\left[\left|\theta-\hat{\theta}\right|<\varepsilon_{\mathrm{th}}\right]\right\},$ (49)
+where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ denotes the threshold which is set as half the sampling interval.
+VII-B4 Performance Analysis
 Figure 6: Correct probability of DoDs at BS and coupled DoAs/DoDs at RIS against SNR.
@@
 Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves.
+This discrepancy highlights the propagation of errors from Phase I to Phase II: the angular estimation errors at BS directly impact the estimation accuracy of the coupled directions at RIS, which in turn affects the CSI estimation for other SCs in subsequent steps.
+However, as the SNR increases, the influence of these propagated errors diminishes, which enables Phase II to achieve improved CSI recovery performance.
 Figure 7: NMSE performance comparison against SNR.
@@
 This suggests that there exists a performance-complexity trade-off between these two algorithms.
-Moreover, it is obvious that the proposed scheme can approach the NMSE performance of oracle-LS scheme, which verifies the effectiveness of the proposed scheme.
+Moreover, it is observed that the proposed scheme progressively approaches the NMSE performance of oracle-LS benchmark as SNR increases.
+Under high SNR conditions, the performance gap between the two schemes becomes negligible, further verifying the effectiveness of the proposed scheme.
 Figure 8: NMSE performance comparison against the number of subframes $T$.
@@
 As expected, the NMSE performance of all the algorithms improves with the increase in the number of subframes $T$.
-It can be observed that the proposed scheme maintains good performance even with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm.
+It can be observed that the proposed scheme maintains good performance even with a short length of subframes, e.g., $T=20$, demonstrating the robustness of the proposed algorithm.
 Moreover, compared to the CBS-GAMP method, the proposed algorithm achieves similar performance with approximately a $50\%$ subframes reduction.
@@
 This indicates that these two schemes exhibit robust framework, which are less susceptible to the lack of data diversity.
-VI Conclusion
+VIII Conclusion
 In this paper, we investigated channel estimation in RIS-assisted wideband THz systems with consideration of beam split effect.
@@
 Moreover, compared to existing solutions, the proposed two-phase strategy significantly reduced computational complexity while maintaining a high level of accuracy in CSI acquisition.
+Appendix A Proof of Proposition 1
+We first define the parameter sequence $\bar{\vartheta}_{q_{1}}=\delta\cdot q_{1}-(Q_{\mathrm{R}}+1)/{Q_{\mathrm{R}}},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ which corresponds one-to-one with $Q_{\mathrm{R}}$ ARVs in the dictionary $\mathbf{C}_{\mathrm{R},m}$.
+Similarly, we define the sequence $\delta_{q_{2}}=-\delta\cdot q_{2}+(Q_{\mathrm{R}}+1)/Q_{\mathrm{R}}$ to correspond to $Q_{R}$ ARVs in the conjugate dictionary $\mathbf{C}_{\mathrm{R},m}^{*}$.
+Thus, the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ can be expressed as $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}})$ for $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$.
+The $j$-th column of the coupled dictionary $\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}$ can be expressed as
+$\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (50)
+where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}\in 1,\cdots,Q_{\mathrm{R}}$.
+Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of unique columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$.
+Figure 12: Coupled dictionary simplification diagram.
+Then, we simplify the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ step by step from (a) to (d) as shown in Fig.
+12.
+As depicted in step (a), we reshape the overcomplete dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ of length $Q_{\mathrm{R}}^{2}$ ARVs into a three-dimensional matrix where each column represents an ARV.
+Given that ARV can be uniquely determined by known angles in (50), we use the ordered sequence $\mathcal{L}_{Q}=\left(1,\dots,Q_{\mathrm{R}}\right)$ to index the corresponding ARVs.
+In step (b), we form the index matrix $\mathbf{Q}\in\mathbb{N}^{Q_{\mathrm{R}}\times Q_{\mathrm{R}}}$ with elements $[\mathbf{Q}]_{i,j}=(\mathcal{L}_{Q})_{i}-(\mathcal{L}_{Q})_{j}$ for $i,j=1,\dots,Q_{\mathrm{R}}$.
+It is noticed that due to the cyclic shift property of this dictionary, the dictionary base can be represented entirely by combining just the first and last rows.
+This means that we can reconstruct the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ with a sliding window of length $Q_{\mathrm{R}}$.
+Therefore, the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ can be given as in (14).
+Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (50), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
+$\boldsymbol{\xi}_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (51)